### v2

# Analysis and Simulation of a Fluid-Heat System in a Thin, Rough Layer in Contact With a Solid Bulk Domain

Abstract
We investigate the effective coupling between heat and fluid dynamics within a thin fluid layer in contact with a solid structure via a rough surface. Moreover, the opposing vertical surfaces of the thin layer are in relative motion. This setup is particularly relevant to grinding processes, where cooling lubricants interact with the rough surface of a rotating grinding wheel. The resulting model is nonlinearly coupled through $(i)$ temperature-dependent viscosity and $(ii)$ convective heat transport. The underlying geometry is highly heterogeneous due to the thin rough surface characterized by a small parameter $\epsilon>0$ that represents both the height of the layer and the periodicity of the roughness.
We analyze this nonlinear system for existence, uniqueness, and energy estimates and study the limit behavior $\varepsilon\to 0$ within the framework of two-scale convergence in thin domains. In this limit, we derive an effective interface model in 3D (a line in 2D) for the heat-fluid interactions inside the fluid. We implement the system numerically and validate the limit problem through a direct comparison with the $\varepsilon$-model. Furthermore, we investigate the influence of the temperature-dependent viscosity and various geometrical configurations with simulation experiments. The corresponding numerical code is freely available on GitHub.
Keywords: Homogenization, mathematical modeling, dimension reduction, two-scale convergence, thin fluid films, numerical simulations
MSC 2020: 35B27, 76-10, 76M50, 65N30, 80M40, 80A19 {NoHyper} 1Corresponding author, Email: tomfre@uni-bremen.de
1 Introduction
This research is motivated by grinding processes, which typically involve three main components: the grinding wheel, the workpiece, and a cooling lubricant. Here, the workpiece is the material that is processed or shaped through grinding and the cooling lubricant is a fluid that is used to reduce friction and remove heat during the grinding process [38]. The grinding wheel itself is usually a composite material consisting of abrasive grains (e.g., diamonds) and a cementing matrix called bond. Given the complex granular surface of the grinding wheel, thin gaps and channels emerge in the contact area of the wheel and the workpiece, allowing the coolant to flow through. Understanding the formation and the influence of these microstructures is crucial in optimizing the usage of coolants [54, 56]. Naturally, resolving the granular wheel structure and the potentially rough workpiece surface is numerically not feasible, creating a need for homogenized models that account for microscopic effects in an effective way.
Macroscopic models describing the heat dynamics in the grinding gap via an interface temperature field can be found in the literature; see [33] and the references therein. These models are heuristically derived and the microstructures and fluid flow in the gap are not accounted for. Additionally, fluid pressure estimates are sometimes also investigated and used as a tool to control the supply of the coolant fluid [45, 58]. The fluid dynamics in the gap is strongly influenced by the viscosity of the coolant, which in reality is temperature dependent [30, 50]; and given the large temperature variations during the grinding process, this effect should not be ignored as it can influence the quality of the results [18, 57]. Numerical simulations to predict the fluid pressure in the contact zone are usually done with the help of the Reynolds equation [53, 55], but often without taking into account the temperature variations.
Closely related to grinding are lubrication problems. Here, the goal is not to remove material from an object, but instead to reduce the friction between two moving objects with the help of oils. Often, temperature variations are encountered (either by friction-induced heat or external supply) that influence the viscosity of the fluid [9, 35]. Another related area is fractured media, where thin cracks are present in a larger bulk domain. Solutes and/or heat can be transported through these thin cracks, see [40] and the references therein. We also point to [39] for a numerical study on temperature-dependent viscosity on flows through fractures. Although there are differences, one typically expects much lower velocities than in our case, the analytical results presented for the temperature could be transferred.
Starting with an idealized periodic setup with a small parameter $\varepsilon\ll 1$ characterizing both the length scale of the periodicity as well as the average thickness of the grinding gap, we rigorously derive an effective, homogenized model accounting for the nonlinear coupling between heat and fluid flow dynamics in such a setup. This is accomplished by conducting a limiting procedure $\varepsilon\to 0$ within the framework of two-scale homogenization for thin domains.
From a mathematical point of view, the problem consists of two main aspects: a large bulk domain where the rough interface acts similar to a rough boundary and a thin rough layer. For bulk domains with a rough boundary, homogenization procedures, including diffusion or fluid flow problems, can be found in [10, 16, 44], or in the case of rough interfaces we point to [19]. The case where the amplitude of the roughness does not scale with $\varepsilon$ has also been considered in the literature, see for example [16, 19, 47] and the references therein. Additionally, there are some connections to dimension reductions in fractured media, where inside a bulk domain a thin layer of width $\varepsilon$ with different material properties appears (for example, a locally varying permeability in a porous medium). Both formal as well as rigorous limit procedures have been carried out for such problems and the thin layer reduces to effective interface equations, see [2, 40, 49].
On the other hand, homogenization of diffusion problems in thin layers with periodic structures has been extensively studied before with the concept of two-scale convergence; see for example [29, 46]. Effective fluid flow through thin layers was rigorously investigated in [11] and later extended for a periodic roughness in [12]. Starting with the stationary Stokes equation, they proved the validity of the Reynolds equation [51] when the layer height is small when compared to the length of the roughness period. If the periodicity and the height are of the same order (usually called the Stokes roughness regime), a model similar to a Darcy equation is derived. The Reynolds equation plays an important role in lubrication, which has a setup similar to our underlying scenario. The equation has been further studied and homogenized in the literature [6, 13, 17, 23]; see also [42, Section 1] for a good overview. For layers without roughness, the first derivation of the Reynolds equation with the concept of two-scale convergence was achieved in [43]. Flow through a thin layer, without oscillations, coupled with a temperature-dependent viscosity was studied in [5], where a Reynolds equation with temperature-dependent viscosity was derived via asymptotic expansion.
To be more specific, we consider a complex geometry incorporating a rough, thin layer similar to the setup in [12], where the boundaries of the layer are additionally in relative motion to each other (see Fig. 1). In this geometry, we are considering the coupling between a quasi-stationary Stokes system for the fluid flow and the heat dynamics. Here, strong coupling comes into play via the convection and temperature-dependent viscosity. We assume that the viscosity is a Lipschitz continuous function with respect to the temperature. We show that the resulting nonlinear problem has a weak solution (Theorem 1) via a Schauder fixed point argument where uniqueness follows under additional regularity assumptions at the solution(Theorem 2). In particular, any strong solution is necessarily unique. We then perform a limiting procedure $\varepsilon\to 0$ in the framework of two-scale convergence for thin domains and, as our main result, arrive at a homogenized model which captures the effective interaction between heat dynamics and fluid flow (Theorem 3). A novel aspect of our work is the derivation of $\varepsilon$-independent $L^{\infty}$ bounds of the temperature field, utilizing results of [41]. These are used to obtain strong two-scale convergence in the thin layer, which is required to pass to the limit in the nonlinear terms. To this end, we also show in Section 3.1 multiple ($\varepsilon$-dependent) trace estimates and embedding estimates for both the thin layer and the bulk domain. Please note that the fluid part of this effective model (e.g., when the temperature is constant) is the same as the limit presented in [12, Section 3.3]. Similarly, as for the $\varepsilon$-problem, we are able to establish uniqueness for the homogenized model under additional regularity assumptions (Theorem 4).
A small convergence study $\varepsilon\to 0$ is carried out numerically in two dimensions, investigating the order of convergence of the error between the effective and microscale model averaged over the thin layer. Moreover, we conduct several numerical experiments showcasing and investigating the effects of the temperature-dependent viscosity, the roughness pattern, as well as the inflow condition. The code of our implementation is uploaded and made freely accessible on GitHub [24].
This paper is structured as follows: In Section 2, we start by introducing the geometrical setup, the mathematical model, and the assumptions needed to pass to the limit $\varepsilon\to 0$. In Section 3 the underlying nonlinear model is analyzed regarding existence and uniqueness. Additionally, $\varepsilon$-dependent solution estimates are derived. Section 4 demonstrates the homogenization process and in Theorem 3 the effective model is presented. Finally, in Section 5, we investigate and compare the effective model and the resolved micro model with the help of numerical simulations.
2 Underlying geometry, model and assumptions
2.1 Geometry setup
We start by describing the geometry and the assumptions on the regularity of the interface. In the following, the superscript $f$ denotes that a property belongs to the fluid layer, while $s$ denotes objects belonging to the solid bulk atop the fluid layer. The considered time interval is denoted by $S=(0,T)$, for $T>0$.
The general setup considered is presented in Fig. 1. We consider spatial domains of dimension $d=2,3$. The unit vectors are denoted with $e_{i}\in\mathbb{R}^{d}$ for $i=1,\dots,d$ and the $i$-dimensional unit cube will be denoted by $Y^{i}=(0,1)^{i}$. The overall domain $\Omega\subset\mathbb{R}^{d}$ is assumed to be of the form $\Omega=\Sigma\times(0,H)$ where $\Sigma\subset\mathbb{R}^{d-1}$ is a bounded $d-1$ dimensional cuboid $\Sigma\subset\mathbb{R}^{d-1}$ with corner coordinates in $\mathbb{Z}^{d-1}$. [^2] So $\Omega$ is also a $d$-dimensional cuboid. By a slight abuse of notation, we will understand $\Sigma$ both as a set in $\mathbb{R}^{d-1}$ and the lower boundary part $\Sigma\times\{0\}\subset\partial\Omega$. Any point $x\in\Omega$ will also be denoted by $x=(\tilde{x},x_{d})\subset\Sigma\times(0,H)$, the same notation is used for $y\in Y^{d}$.
Figure 1: Schematic depiction of the geometric setup. Left: The complete domain with the different subdomains and boundaries. Note that $\Omega=\Omega^{s}_{\varepsilon}\cup\Omega^{f}_{\varepsilon}\cup\Gamma_{\varepsilon}$. Right: The reference cell.
The rough interface between solid and fluid is given through a $Y^{d-1}$-periodic $C^{1}$ function $\gamma:\overline{Y^{d-1}}\to[\gamma_{0},1]$ for some $10<\gamma_{0}<$ 1. This lower bound $\gamma_{0}>0$ ensures that the interface does not touch the bottom boundary; see Remark 2 for a discussion regarding this choice and case $\gamma_{0}=0$ and a connection to our motivating application. We introduce the reference cell $Z\subseteq Y^{d}$ and the reference interface $\Gamma$ given by
$Z=\left\{x\in Y^{d}:x_{d}<\gamma(\tilde{x})\right\}\quad\text{ and }\quad\Gamma=\left\{x\in Y^{d}:x_{d}=\gamma(\tilde{x})\right\}.$
Since the corners of $\Sigma$ are at integer coordinates, it can be tiled perfectly with $\varepsilon Y^{d-1}$ for $\varepsilon:=\varepsilon_{n}=n^{-1}$ for any $n\in\mathbb{N}$. In the following, we suppress the $n$ and note that any limit $\varepsilon\to 0$ is understood as a limit $n\to\infty$. With that in mind, we introduce the rough solid domain $\Omega^{s}_{\varepsilon}$ and the rough, thin fluid layer $\Omega^{f}_{\varepsilon}$ by (assuming that $\varepsilon<H$)
$\Omega^{s}_{\varepsilon}=\left\{x\in\Omega:x_{d}>\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right\}\quad\text{and}\quad\Omega^{f}_{\varepsilon}=\left\{x\in\Omega:x_{d}<\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right\}.$
Here, roughness refers to the fast, oscillatory structure (due to $\nicefrac{{x}}{{\varepsilon}}$) of the interface between the fluid and grinding wheel which is given by
$\Gamma_{\varepsilon}=\left\{x\in\Omega:x_{d}=\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right\}.$
Note that $\Omega=\Omega^{s}_{\varepsilon}\cup\Omega^{f}_{\varepsilon}\cup\Gamma_{\varepsilon}$. We denote the outer boundary of the solid domain by $\Sigma^{s}_{\varepsilon}=\partial\Omega^{s}_{\varepsilon}\setminus\Gamma_{\varepsilon}$. Now, let $\Sigma^{\text{in}},\Sigma^{\text{out}}\subset\partial\Sigma$ be two disjoint sets with positive surface measure such that $\Sigma^{\text{in}}\cup\Sigma^{\text{out}}=\partial\Sigma$. We then split the boundary of $\Omega^{f}_{\varepsilon}$ into
$\partial\Omega^{f}_{\varepsilon}=\Sigma\cup\Gamma_{\varepsilon}\cup\Sigma^{\text{in}}_{\varepsilon}\cup\Sigma^{\text{out}}_{\varepsilon}$
where the sets $\Sigma^{\text{in}}_{\varepsilon},\Sigma^{\text{out}}_{\varepsilon}$ are given by
$\Sigma^{\text{in}}_{\varepsilon}=\left\{(\tilde{x},x_{d}):\tilde{x}\in\Sigma^{\text{in}},x_{d}\in\left(0,\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right)\right\}\quad\text{ and }\quad\Sigma^{\text{out}}_{\varepsilon}=\left\{(\tilde{x},x_{d}):\tilde{x}\in\Sigma^{\text{out}},x_{d}\in\left(0,\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right)\right\}.$
Finally, with $\Omega_{\varepsilon}=\Sigma\times(0,\varepsilon)$, we denote the thin layer without any roughness. Note, that both $H$ and $|\Sigma|$ are independent of $\varepsilon$ and are therefore of order $\mathcal{O}(1)$. The volume of the fluid layer $|\Omega^{f}_{\varepsilon}|$ and the area of the lateral boundary $|\Sigma^{\text{in}}_{\varepsilon}|,|\Sigma^{\text{out}}_{\varepsilon}|$ are of order ${O}(\varepsilon)$. Furthermore, for the length of the rough interface and the remaining boundary of the solid domain, it holds $|\Gamma_{\varepsilon}|,|\Sigma^{s}_{\varepsilon}|=\mathcal{O}(1)$.
In our motivating example of a grinding process, the grinding wheel moves over the workpiece, with a given velocity $u_{\text{motion}}$. Since we consider only the section of the grinding gap, it is reasonable to assume that $u_{\text{motion}}\in\mathbb{R}^{d}$ and $u_{\text{motion}}\cdot e_{d}=0$. Because of the movement both domains $\Omega^{s}_{\varepsilon}$ and $\Omega^{f}_{\varepsilon}$ would be time-dependent. To circumvent the homogenization of a moving domain we will view the problem from the reference frame of the grinding wheel and obtain a fixed geometry, see also the model description in the next section.
Remark 1.
In the case of grinding, the lower boundary of the fluid layer would also be given by a rough surface describing the workpiece. In our setup, this roughness may be characterized by another function $\gamma_{2}$, but with the same period of $\varepsilon$. It is also reasonable to assume that the workpiece roughness is constant in the direction of $u_{\text{motion}}$, therefore one can again work in a fixed reference frame. To simplify the notation we only consider one oscillating interface, but the following analysis can be transferred to the more general case. The smoothness assumption of $\gamma$ stems from technical reasons and is needed to estimate the fluid pressure.
With the subscript ##\## we will indicate a space of periodic functions in the directions $11,\dots,d-$ 1, for example
$H^{1}_{\#}(Z)\coloneqq\left\{u\in H^{1}_{\text{loc}}(\mathbb{R}^{d}):u_{|Z}\in H^{1}(Z),u(y+e_{i})=u(y)\text{ for almost all }y\in Z\text{ and }i=1,\dots,d-1\right\}.$
Additionally, we define $\nabla_{\tilde{x}}u=(\partial_{x_{1}}u,\dots,\partial_{x_{d-1}}u,0)$, where a 00 is appended such that the gradient it compatible with the usual gradient in $\mathbb{R}^{d}$. With $\|\cdot\|_{H^{1}(\Omega^{f}_{\varepsilon}\cup\Omega^{s}_{\varepsilon})}$ we will denote the combined $H^{1}$ norm on both subdomains, e.g.,
$\|\cdot\|_{H^{1}(\Omega^{f}_{\varepsilon}\cup\Omega^{s}_{\varepsilon})}\coloneqq\sqrt{\|\cdot\|_{H^{1}(\Omega^{f}_{\varepsilon})}^{2}+\|\cdot\|_{H^{1}(\Omega^{s}_{\varepsilon})}^{2}}.$
2.2 Underlying model equations
For a function $\varphi\in L^{2}(\Omega)$, we use the notation $\varphi_{\varepsilon}^{k}:=\varphi_{|\Omega_{\varepsilon}^{k}}$ ($k=f,s$) for the restriction of $\varphi$ to the subdomain $\Omega_{\varepsilon}^{k}$. A similar notation holds for the coefficients of the models, which are assumed to be constant in each subdomain. We suppress superscripts and the index $\varepsilon$ when possible to improve readability.
We consider a reference frame of the grinding wheel, where instead of a moving interface $\Gamma_{\varepsilon}$ we have a prescribed velocity at the bottom boundary $\Sigma$. To describe the temperature $\theta_{\varepsilon}=(\theta^{s}_{\varepsilon},\theta^{f}_{\varepsilon})$, we utilize heat equations in both domains
[…]
Here, $\kappa=k_{f}$ in $\Omega^{f}_{\varepsilon}$ and $\kappa=\kappa_{s}$ in $\Omega^{s}_{\varepsilon}$ are the heat conductivities in each subdomain and $u_{\varepsilon}$ is the fluid velocity given via a Stokes system below. As they have no specific bearing on the mathematical analysis or the homogenization result, we have set the volumetric heat capacity to 1 in both phases.
At the outer boundary, we apply a Dirichlet condition on $\Sigma^{\text{in}}_{\varepsilon}$ for a prescribed cooling temperature (for simplification set to 0, e.g., the temperature is seen relative to the inflow temperature, which is possible in this case since we always assume $\operatorname{div}{u_{\varepsilon}}=0$), and on the other boundary sections we apply a homogeneous Neumann condition:
[…]
The temperature profiles of the two domains are coupled with a Robin exchange condition:
[…]
where $\llbracket\theta\rrbracket=(\theta^{f}-\theta^{s})$ denotes the usual jump across an interface, $\alpha$ denotes the heat exchange parameter, and $f_{\varepsilon}$ describes a heat source. Here, the chosen heat source is motivated by our underlying application problem of grinding, there the system heats up through friction between the workpiece and the grinding wheel. Therefore, the heat is produced at the interface. To simplify the notation the heat source is only located on the interface side of the solid domain. Through the Robin condition heat is then also transferred into the fluid.
Note, the specific $\varepsilon$-scaling can be reasoned from a non-dimensionalization step and can be found in similar problems [20, 29] and originates from the scaling of the physical parameters (e.g., a fast heat conductivity in the thin layer compared to the height of the layer and the scaling of the density keeps the total amount of heat energy constant with respect to $\varepsilon$). Clearly Eq. 2.1b could be multiplied with $\varepsilon$ and the scaling would then only enter on the right-hand side of (2.1f). For the following analysis, the above form is more appropriate.
Next to the temperature, we also model the fluid flow inside $\Omega^{f}_{\varepsilon}$. Fluid velocity $u_{\varepsilon}$ and pressure $p_{\varepsilon}$ are given as the solution of a Stokes system with temperature-dependent viscosity $\mu$,
[…]
2.3 Assumptions on data
To ensure the existence of solutions and to pass to the limit $\varepsilon\to 0$, we need multiple assumptions on the data; these are collected in this section.
  1. (A1)
    For the coefficients we assume $\alpha,\kappa^{f},\kappa^{s}>0$ and that they are of order $\mathcal{O}(1)$.
  2. (A2)
    The temperature dependent viscosity is given by a Lipschitz continuous function $\mu:\mathbb{R}\to\mathbb{R}$ that fulfills $\underline{\smash{\mu}}\leq\mu(x)\leq\overline{\mu}$ for some $0<\underline{\smash{\mu}}\leq\overline{\mu}<\infty$, independent of $\varepsilon$.
  3. (A3)
    The initial conditions fulfill that $\theta_{\varepsilon,0}\in L^{\infty}(\Omega)$ and $C_{0}\coloneqq\sup_{\varepsilon>0}\|\theta_{\varepsilon,0}\|_{L^{\infty}(\Omega)}<\infty$.
  4. (A4)
    For the heat source it holds $f_{\varepsilon}\in L^{\infty}(S\times\Gamma_{\varepsilon})$ and $C_{f}\coloneqq\sup_{\varepsilon>0}\|f_{\varepsilon}\|_{L^{\infty}(S\times\Gamma_{\varepsilon})}<\infty$.
  5. (A5)
    For the prescribed velocity on the lateral fluid boundaries, we assume there is $u_{BC}\in W^{1,\infty}(\partial\Sigma\times(0,1))^{d}$ with the properties:
    • $u_{BC}\cdot e_{d}=0$,
    • $u_{BC}(\tilde{x},0)=u_{\text{motion}}$, where it still holds that $u_{\text{motion}}\cdot e_{d}=0$,
    • $u_{BC}(\tilde{x},x_{d})=0$ for almost all $\tilde{x}\in\partial\Sigma$ and $x_{d}\geq\gamma_{0}$,
    • inflow and outflow conditions $\left(u_{BC}\cdot n_{\Sigma}\right)_{|\Sigma^{\text{in}}}<0\text{ and }\left(u_{BC}\cdot n_{\Sigma}\right)_{|\Sigma^{\text{out}}}\geq 0$.
    In addition, we require
    $\int_{\partial\Sigma\times(0,1)}u_{BC}\cdot n_{\Sigma}\,\mathrm{d}\sigma_{x}=0$ (2.2)
    to ensure the existence of solutions. Here, $n_{\Sigma}$ is the normal vector that points outward from $\Sigma\times(0,1)$. The boundary condition on the thin layer is given by $u_{\varepsilon,BC}(x)=u_{BC}(\tilde{x},x_{d}/\varepsilon)$ for $x\in\Sigma^{\text{in}}_{\varepsilon}\cup\Sigma^{\text{out}}_{\varepsilon}$.
  6. (A6)
    For the initial conditions there are limit functions $\theta^{s}_{0}\in L^{2}(\Omega)$ and $\theta^{f}_{0}\in L^{2}(\Sigma)$ such that $\chi_{\Omega^{s}_{\varepsilon}}\theta_{\varepsilon,0}\to\theta^{s}_{0}$ in $L^{2}(\Omega^{s}_{\varepsilon})$ and $\chi_{\Omega^{f}_{\varepsilon}}\theta^{f}_{\varepsilon,0}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\theta^{f}_{0}$ in $L^{2}(\Omega_{\varepsilon})$, for $\varepsilon\to 0$. Here, $\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}$ denotes two-scale convergence in the thin layer which will be introduced in Definition 1.
  7. (A7)
    For the heat source there is a $f\in L^{2}(S\times\Sigma\times\Gamma)$ such that $f_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}f$, for $\varepsilon\to 0$.
The Assumptions (A1) and (A2) are often encountered in similar problems, especially the properties of $\mu$ are needed to ensure the existence of a solution.
Given the assumption (A5), we can always construct a continuation $U\in H^{1}(\Sigma\times(0,1))$ of $u_{BC}$, with $U(\tilde{x},y)=0$ for $y\geq\gamma_{0}$ and $\operatorname{div}{U}=0$, [32, Theorem. 3.5]. By setting $U_{\varepsilon}(x)=(U_{1},\dots,\varepsilon U_{d})(\tilde{x},\frac{x_{d}}{\varepsilon})$ the extension can also be transferred to the thin rough layer $\Omega^{f}_{\varepsilon}$, this construction was already derived in [12]. We expect that the condition $u_{BC}=0$ for $x_{d}\geq\gamma_{0}$ could be removed and the extension still be carried out by first building an extension around the boundary falling back to a rectangle of height $\varepsilon\gamma_{0}$ and then proceeding as above, but this is not further investigated here. Examples of functions that fulfill the assumptions can be found in Section 5. We could also utilize a time-dependent boundary function $u_{BC}\in L^{\infty}(S;W^{1,\infty}(\partial\Sigma\times(0,1)))^{d}$, this would not change the following results. To simplify the notation we keep a time-independent $u_{BC}$. Additionally, given the prescribed boundary conditions of $u_{BC}$, it holds $u_{\varepsilon,BC}=u_{\text{motion}}$ on $\Sigma$ and we and we have inflow at $\Sigma^{\text{in}}_{\varepsilon}$ and outflow at $\Sigma^{\text{out}}_{\varepsilon}$, to be precise
$\left(u_{\varepsilon,BC}\cdot n_{\varepsilon}\right)_{|\Sigma^{\text{in}}_{\varepsilon}}<0\quad\text{ and }\quad\left(u_{\varepsilon,BC}\cdot n_{\varepsilon}\right)_{|\Sigma^{\text{out}}_{\varepsilon}}\geq 0.$ (2.3)
Remark 2.
We provide some additional comments regarding our model and our mathematical assumptions specifically in regards to the application case of grinding:
  • $(i)$
    During the actual physical process there is, of course, contact between the grinding wheel and the workpiece. In our model, that would imply that $\Omega^{s}_{\varepsilon}$ and $\Sigma$ are touching (i.e., $\gamma_{0}=0$). Note that for dimension $d=2$, the model would no longer be reasonable as any contact would necessarily disrupt fluid flow; while for $d=3$ it would still be possible for the fluid to flow around the contact points. In other works where the behavior of the fluid pressure inside the grinding gap is studied, an artificial gap is introduced such that the pressure can be modeled by the Reynolds equation [53, 55]. Their findings demonstrate that this is a reasonable approximation to the real dynamics. Therefore, we keep this simplification of no direct contact and focus on how far a temperature-dependent viscosity influences the model.
  • $(ii)$
    The heat which is released during grinding is caused by the friction between the grains of the grinding wheel and the workpiece and only part of this heat is released into the system directly (e.g., part of it is also transported with the formation of small chips [48]). Here, we consider a simplified model where we assume a prescribed heat source at the interface but where we do not account for any mechanical effects in the solid part (in particular: no changing geometries, no chips, and no strains).
  • $(iii)$
    The relationship between temperature and viscosity of fluids is complex and often modeled heuristically. A typical example is the Vogel-Fulcher-Tammann equation given as
    $\mu(\theta)=\mu_{0}\exp\left(\frac{\alpha}{\theta-T_{0}}\right)$
    where $\mu_{0},\alpha,T_{0}$ are fitting parameters, we refer to [30] and references therein. This, of course, is only valid in a certain temperature range (usually a proper subset of the liquid phase range of the substance). However, it suggests that the relationship tends to be very smooth and that the Lipschitz continuity is a reasonable assumption.
  • $(iv)$
    The specific assumptions regarding $u_{BC}$ and, as a consequence, $u_{\varepsilon,BC}$ are a bit technical, but they ensure that there is an inflow and an outflow region and that $\Sigma$ and $\Gamma_{\varepsilon}$ are in relative motion to each other. The regularity assumption ensures compatibility across the different boundary parts. The inflow condition models the continuous injection of cooling liquid where the corresponding Dirichlet condition Eq. 2.1d accounts for the cooling effect.
  • $(v)$
    We do not include inertia, turbulence, and buoyancy effects by using the quasi-stationary Stokes system. This is somewhat justified by the thinness of the contact area where the fluid flow does not have enough space to create strong turbulence and is mainly pulled along by the moving wheel as demonstrated in the experiments in [21]. We also neglect the real and important phenomenon of air entrainment into the fluid layer [34, 36]. Additionally, the fluid movement adjusts on a shorter time interval than the temperature.
3 The microscale model
We start by analyzing the model on the $\varepsilon$-scale to ensure that it is well-posed and to derive solution bounds needed for the limiting procedure. To this end, we first list some technical lemmas needed in the following sections.
3.1 Auxiliary results
Lemma 1 (Extension operators and transformation).
  • $(i)$
    For all $\varepsilon>0$, there exist a linear extension operator $\mathbb{E}_{\varepsilon}:W^{1,p}(\Omega^{f}_{\varepsilon})\to W^{1,p}(\Omega_{\varepsilon})$ such that
    $\|\mathbb{E}_{\varepsilon}\psi\|_{W^{1,p}(\Omega_{\varepsilon})}\leq C\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}\quad\text{ for all }\psi\in W^{1,p}(\Omega^{f}_{\varepsilon})$
    and for all $p\in[1,\infty)$. Here, $C>0$ may depend on $p$ but is independent of $\varepsilon$.
  • $(ii)$
    There exists a diffeomorphism $\Phi:\Omega\to\Omega^{s}_{\varepsilon}$ such that
    $\frac{1}{C}\|\varphi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}\leq\|\varphi\circ\Phi\|_{W^{1,p}(\Omega)}\leq C\|\varphi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}\quad\text{ for all }\varphi\in W^{1,p}(\Omega^{s}_{\varepsilon})$
    and for all $p\in[1,\infty)$. Here, $C>0$ may depend on $p$ but is independent of $\varepsilon$.
Proof.
$(i)$. The extension operators can be constructed using a classical extension procedure [1]. For thin layers with a more general geometry, this was carried out in detail and further studied in [28, Theorem 1]. The general idea is to first construct extension operators for a single cell, in our case from $W^{1,p}(Z)$ to $W^{1,p}(Y^{d})$. Those exist since $\gamma$ is Lipschitz-continuous, and they are naturally $\varepsilon$-independent (as there is no $\varepsilon$ involved yet). The statement itself then follows by gluing these extension operators together via a periodic partition of unity in conjunction with a scaling argument.
$(ii)$. For the transformation of $\Phi$ to the complete domain $\Omega$, we define the functions
$\phi_{1}(x)\coloneqq\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})+x_{d}\frac{H-\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}{H}\quad\text{ and }\quad\phi_{2}(x)\coloneqq H\frac{x_{d}-\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}{H-\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})},$
where for a fixed $\tilde{x}\in\Sigma$ it holds
$\phi_{1}(\tilde{x},\cdot):(0,H)\to(\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}}),H)\quad\text{ and }\quad\phi_{2}(\tilde{x},\cdot):(\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}}),H)\to(0,H).$
With $\phi_{1}$ we obtain the transformation
$\Phi:\Omega\to\Omega^{s}_{\varepsilon},\quad x\mapsto(\tilde{x},\phi_{1}(\tilde{x},x_{d}))$
Since $\gamma$ is in $C^{1}(\overline{Y^{d-1}};[\gamma_{0},1])$ and $H>\varepsilon$ it holds that $G$ is bijective and continuously differentiable. The inverse is given by $\Phi^{-1}(x)=(\tilde{x},\phi_{2}(\tilde{x},x_{d}))$. Given the properties of $\gamma$, it holds
$|\phi_{1}(x)|,|\phi_{2}(y)|\leq C,\quad|\nabla\phi_{1}(x)|,|\nabla\phi_{2}(y)|\leq C\quad\text{ and }\quad c\leq\det(\nabla\Phi),\det(\nabla\Phi^{-1})\leq C$
for all $x\in\Omega,y\in\Omega^{s}_{\varepsilon}$, where $C,c>0$ can be chosen independent of $\varepsilon$. From these bounds we also obtain $\varepsilon$ independent bounds of $|\nabla\Phi|$ and $|\nabla\Phi^{-1}|$. The estimates in the $W^{1,p}$ norm for $\varphi$ and $\varphi\circ\Phi$ then follow via integration by substitution. ∎
Lemma 2 (Trace estimates).
  • $(i)$
    For every $p\in[1,\infty)$, there is a constant $C>0$ (independent of $\varepsilon$) such that
    $\|\psi\|_{L^{p}(\Gamma_{\varepsilon})}\leq\frac{C}{\varepsilon^{1/p}}\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}\quad\text{for all}\quad\psi\in W^{1,p}(\Omega^{f}_{\varepsilon})$
  • $(ii)$
    For every $p\in[1,\infty)$, there is $C>0$ (independent of $\varepsilon$) such that
    $\|\phi\|_{L^{p}(\Gamma_{\varepsilon})}\leq C\|\phi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}\quad\text{for all}\quad\varphi\in W^{1,p}(\Omega^{s}_{\varepsilon})$
  • $(iii)$
    Let $\Sigma^{f}_{\varepsilon}\coloneqq\partial\Omega^{f}_{\varepsilon}\setminus(\Sigma\cup\Gamma_{\varepsilon})$ denote the lateral boundary of the fluid domain. For every $p\in[1,\infty)$, there is a constant $C>0$ (independent of $\varepsilon$) such that
    $\|\psi\|_{L^{p}(\Sigma^{f}_{\varepsilon})}\leq C\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}\quad\text{for all}\quad\psi\in W^{1,p}(\Omega^{f}_{\varepsilon}).$
Proof.
$(i)$. This estimate for $\psi\in W^{1,p}(\Omega^{f}_{\varepsilon})$ is quite standard and follows by a splitting argument, which we shortly sketch. Split the integral over $\Gamma_{\varepsilon}$ into integrals over the boundary of the cell $\varepsilon Z$, applying a scaling argument followed by a trace estimate in the rescaled domain and then transforming back to the $\varepsilon$ domain. Note, that the estimate could be improved by a $\varepsilon^{1/p}$ in front of the gradient terms, but this is not needed in the following analysis.
$(ii)$. A similar estimate was derived in [19, Proposition 2], albeit with homogeneous boundary conditions at the outer boundary $\partial\Omega^{s}_{\varepsilon}\setminus\Gamma_{\varepsilon}$ for $\varphi\in H^{1}(\Omega^{s}_{\varepsilon})$. By construction of the transformation $\Phi:\Omega\to\Omega^{s}_{\varepsilon}$ from Lemma 1, it holds $(\varphi\circ\Phi)_{|\Sigma}=\varphi_{|\Gamma_{\varepsilon}}$ and
$\int_{\Gamma_{\varepsilon}}|\varphi(x)|^{p}\,\mathrm{d}\sigma_{x}$ $=\int_{\Sigma}\left|\varphi(\tilde{x},\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}}))\right|^{p}\sqrt{1+|\varepsilon\nabla_{\tilde{x}}\left[\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})\right]|^{2}}\,\mathrm{d}\sigma_{\tilde{x}}$
$=\int_{\Sigma}\left|(\varphi\circ\Phi)(\tilde{x},0)\right|^{p}\sqrt{1+|\left[\nabla_{\tilde{x}}\gamma\right](\nicefrac{{\tilde{x}}}{{\varepsilon}})|^{2}}\,\mathrm{d}\sigma_{\tilde{x}}$
$\leq C\int_{\Sigma}\left|(\varphi\circ\Phi)(\tilde{x},0)\right|^{p}\,\mathrm{d}\sigma_{\tilde{x}}.$
On $\Omega$ we can use a trace theorem independent of $\varepsilon$ to obtain with Lemma 1.($ii$)
$\|\varphi\|_{L^{p}(\Gamma_{\varepsilon})}\leq C\|\varphi\circ\Phi\|_{L^{p}(\Sigma)}\leq C\|\varphi\circ\Phi\|_{W^{1,p}(\Omega)}\leq C\|\varphi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}.$
$(iii).$ This follows by a scaling argument. With the extension operator from Lemma 1, it holds
$\|\psi\|_{L^{p}(\Sigma^{f}_{\varepsilon})}$ $=\|\mathbb{E}\psi\|_{L^{p}(\Sigma^{f}_{\varepsilon})}\leq\|\mathbb{E}\psi\|_{L^{p}(\partial\Sigma\times(0,\varepsilon))}=\varepsilon^{\frac{1}{p}}\|\mathbb{E}\psi(\tilde{x},\varepsilon x_{d})\|_{L^{p}(\partial\Sigma\times(0,1))}$
$\leq C\varepsilon^{\frac{1}{p}}\|\mathbb{E}\psi(\tilde{x},\varepsilon x_{d})\|_{W^{1,p}(\Omega)}\leq C\|\mathbb{E}\psi\|_{W^{1,p}(\Omega_{\varepsilon})}\leq C\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}.$
Where we used the trace theorem on $\Omega$. Note, as before, the estimate could be improved by a $\varepsilon^{1/p}$ in front of the term with the $\partial_{x_{d}}$ norm, but this is not needed in the following analysis. ∎
We later need to show boundedness of the temperature solutions to get better estimates for the time derivatives. There we rely on the following lemma, which follows by standard results via a natural scaling argument:
Lemma 3 (Interpolation embedding for thin domains).
There is $C>0$ such that, for any $\psi\in L^{\infty}(S;L^{2}(\Omega_{\varepsilon}))\cap L^{2}(S;H^{1}(\Omega_{\varepsilon}))$,
$\|\psi\|_{L^{r_{d}}(S\times\Omega_{\varepsilon})}\leq C\varepsilon^{-\frac{1}{d+2}}\left(\|\psi\|_{L^{\infty}(S;L^{2}(\Omega_{\varepsilon}))}+\|\psi\|_{L^{2}(S;H^{1}(\Omega_{\varepsilon}))}\right).$
Here, $r_{d}=\frac{2(d+2)}{d}$, $d=2,3$.
Proof.
For any $\psi\in L^{\infty}(S;L^{2}(\Omega_{\varepsilon})\cap L^{2}(S;H^{1}(\Omega_{\varepsilon}))$, let $\bar{\psi}\in L^{\infty}(S;L^{2}(\Omega))\cap L^{2}(S;H^{1}(\Omega))$ be the corresponding upscaled function given by $\bar{\psi}(t,\tilde{x},x_{d})=\psi(t,\tilde{x},\varepsilon x_{d})$. Using standard interpolation embeddings (see, e.g., [41, Chapter 2, Inequality 3.8]), it holds
$\|\bar{\psi}\|_{L^{r_{d}}(S\times\Omega)}\leq C\left(\|\bar{\psi}\|_{L^{\infty}(S;L^{2}(\Omega))}+\|\bar{\psi}\|_{L^{2}(S;H^{1}(\Omega))}\right)$
for some constant $C>0$ which does not depend on $\varepsilon$. With a standard scaling argument, we also see that
$\|\bar{\psi}\|_{L^{r_{d}}(S\times\Omega)}=\varepsilon^{\frac{-1}{r_{d}}}\|\psi\|_{L^{r_{d}}(S\times\Omega_{\varepsilon})},\quad\|\bar{\psi}\|_{L^{\infty}(S;L^{2}(\Omega))}=\varepsilon^{\frac{-1}{2}}\|\psi\|_{L^{\infty}(S;L^{2}(\Omega_{\varepsilon}))}.$
Similarly, for the term including the $H^{1}$-norm, we estimate (note that $\varepsilon<1$)
$\|\bar{\psi}\|_{L^{2}(S;H^{1}(\Omega))}^{2}=\varepsilon^{-1}\left(\|\psi\|_{L^{2}(S\times\Omega_{\varepsilon})}^{2}+\|\nabla_{\tilde{x}}\psi\|_{L^{2}(S\times\Omega_{\varepsilon})}^{2}+\varepsilon\|\partial_{x_{d}}\psi\|_{L^{2}(S\times\Omega_{\varepsilon})}^{2}\right)\leq\varepsilon^{-1}\|\psi\|_{L^{2}(S;H^{1}(\Omega_{\varepsilon}))}^{2}.$
Since $\nicefrac{{1}}{{r_{d}}}-\nicefrac{{1}}{{2}}=\nicefrac{{-1}}{{d+2}}$, the statement follows. ∎
3.2 Analysis of the micro model
Next, we investigate the underlying model on the microscale. First, we define
$L^{2}_{0}(\Omega^{f}_{\varepsilon})\coloneqq\left\{\varphi\in L^{2}(\Omega^{f}_{\varepsilon}):\int_{\Omega^{f}_{\varepsilon}}\varphi\,\mathrm{d}x=0\right\}\quad\text{and}\quad H^{1}_{A}(\Omega^{f}_{\varepsilon})\coloneqq\left\{\varphi\in H^{1}(\Omega^{f}_{\varepsilon}):\varphi_{|A}=0\right\},$
where $A\subseteq\partial\Omega^{f}_{\varepsilon}$. To shorten the notation we use, when applicable, the usual notation $H^{1}_{0}(\Omega^{f}_{\varepsilon})=H^{1}_{\partial\Omega^{f}_{\varepsilon}}(\Omega^{f}_{\varepsilon})$. For the solutions of the temperature space we define $\Theta_{\varepsilon}\coloneqq H^{1}(\Omega^{s}_{\varepsilon})\times H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})$. For the fluid velocity and pressure we consider
$V_{\varepsilon}\coloneqq\left\{v\in L^{\infty}(S;H^{1}(\Omega^{f}_{\varepsilon})^{d}):v=u_{\varepsilon,BC}\text{ on }\partial\Omega^{f}_{\varepsilon}\text{ and }\operatorname{div}{v}=0\text{ a.e. in }\Omega^{f}_{\varepsilon}\right\}\quad\text{ and }\quad Q_{\varepsilon}\coloneqq L^{\infty}(S;L^{2}_{0}(\Omega^{f}_{\varepsilon})).$
The weak formulation of the heat equation, for a fixed $v\in V_{\varepsilon}$, is given by: Find $\theta_{\varepsilon}=(\theta^{s}_{\varepsilon},\theta^{f}_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})$ such that $\partial_{t}\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$ and for all $\varphi\in L^{2}(S;\Theta_{\varepsilon})$ and almost all $t\in S$
$\begin{split}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{s}_{\varepsilon}}&+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{f}_{\varepsilon}}\\ &+\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}+\alpha(\llbracket\theta_{\varepsilon}\rrbracket,\llbracket\varphi\rrbracket)_{\Gamma_{\varepsilon}}=(f_{\varepsilon},\varphi^{s})_{\Gamma_{\varepsilon}}.\end{split}$ (3.1)
[…]
$\begin{split}&(\mu(\psi_{\varepsilon})\nabla u_{\varepsilon},\nabla v)_{\Omega^{f}_{\varepsilon}}-(p_{\varepsilon},\operatorname{div}{(v)})_{\Omega^{f}_{\varepsilon}}=0.\end{split}$ (3.2)
For a fixed $\varepsilon>0$ we now prove that there exists a solution of the coupled nonlinear system given by Eq. 3.1 and (3.2). To show this, we follow a similar strategy as in [50], where we split up the problem, first consider both weak formulations on their own, and then show a solution of the coupled system by a fix-point argument of an iterative scheme.
Lemma 4.
Let the Assumptions (A1), (A3) and (A4) be fulfilled. For a fixed $\varepsilon>0$ and a given $v\in V_{\varepsilon}$ there exists a unique solution $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ satisfying $\partial_{t}\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$, $\theta_{\varepsilon}(0,\cdot)=\theta_{\varepsilon,0}$ a.e. and Eq. 3.1. In addition, it holds
$\begin{split}\|\theta_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{s}_{\varepsilon}))}+\frac{1}{\sqrt{\varepsilon}}\|\theta_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\|\nabla\theta_{\varepsilon}\|_{L^{2}(S;L^{2}(\Omega^{s}_{\varepsilon}))}+\frac{1}{\sqrt{\varepsilon}}\|\nabla\theta_{\varepsilon}\|_{L^{2}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\|\llbracket\theta_{\varepsilon}\rrbracket\|_{L^{2}(S\times\Gamma_{\varepsilon})}\leq C,\end{split}$ (3.3)
where $C<\infty$ can be chosen independent of $\varepsilon$ and $v$.
Proof.
We start by showing the existence and uniqueness of a solution, for a fixed $v\in V_{\varepsilon}$. First, note that our problem is of the form
$\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+A_{\varepsilon}(\theta_{\varepsilon},\varphi;t)=(f_{\varepsilon},\varphi^{s})_{\Gamma_{\varepsilon}},$ (3.4)
with the time-dependent bilinear form
$A_{\varepsilon}(\cdot,\cdot,t):L^{2}(S;\Theta_{\varepsilon})\times L^{2}(S;\Theta_{\varepsilon})\to\mathbb{R},\\ A_{\varepsilon}(\psi,\varphi;t)\coloneqq(\kappa\nabla\psi(t),\nabla\varphi(t))_{\Omega^{s}_{\varepsilon}}+\frac{1}{\varepsilon}(\kappa\nabla\psi(t),\nabla\varphi(t))_{\Omega^{f}_{\varepsilon}}+\frac{1}{\varepsilon}(v\cdot\nabla\psi(t),\varphi(t))_{\Omega^{f}_{\varepsilon}}+\alpha(\llbracket\psi(t)\rrbracket,\llbracket\varphi(t)\rrbracket)_{\Gamma_{\varepsilon}}.$
By the embedding $H^{1}(\Omega)\hookrightarrow L^{4}(\Omega)$, the convective term $(v\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}$ is well defined. We obtain that $A_{\varepsilon}$ is bounded, to be precise for all $\psi,\varphi\in L^{2}(S;\Theta_{\varepsilon})$ and almost all $t\in S$
$|A_{\varepsilon}(\psi,\varphi;t)|\leq C\left(1+\frac{1}{\varepsilon}+\|v\|_{L^{\infty}(S;H^{1}(\Omega^{f}_{\varepsilon}))}\right)\|\psi(t)\|_{H^{1}(\Omega^{f}_{\varepsilon}\cup\Omega^{s}_{\varepsilon})}\|\phi(t)\|_{H^{1}(\Omega^{f}_{\varepsilon}\cup\Omega^{s}_{\varepsilon})},$
for a $C<\infty$. Additionally, via integration by parts we have (suppressing the time variable $t$)
$(v\cdot\nabla\psi,\varphi)_{\Omega^{f}_{\varepsilon}}=-(v\cdot\nabla\varphi,\psi)_{\Omega^{f}_{\varepsilon}}-(\operatorname{div}{(v)}\psi,\varphi)_{\Omega^{f}_{\varepsilon}}+(v\cdot n_{\varepsilon},\psi\varphi)_{\partial\Omega^{f}_{\varepsilon}}.$ (3.5)
Plugging in $\varphi=\psi$, that $\operatorname{div}{(v)}=0$ and the boundary conditions on $\partial\Omega^{f}_{\varepsilon}$ into Eq. 3.5 leads to
$(v\cdot\nabla\psi,\psi)_{\Omega^{f}_{\varepsilon}}=\frac{1}{2}\int_{\Sigma^{\text{out}}_{\varepsilon}}u_{\varepsilon,BC}\cdot n_{\varepsilon}\,\psi^{2}\,\mathrm{d}\sigma_{x}\geq 0$
for almost all $t\in S$. Therefore, there exists a $\lambda>0$ such that
$A_{\varepsilon}(\psi,\psi;t)+\lambda\left(\|\psi(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\|\psi(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}\right)\geq C\|\psi(t)\|^{2}_{H^{1}(\Omega^{f}_{\varepsilon}\cup\Omega^{s}_{\varepsilon})}$
for all $\psi\in L^{2}(S;\Theta_{\varepsilon})$ and almost all $t\in S$. Given the representation (3.4) and the above estimates for the bilinear form $A_{\varepsilon}$, there exists a unique solution $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ of the problem [52, Chapter 3.2].
For the norm estimate, we utilize the above computations of the convection term, use $\varphi=\theta_{\varepsilon}$ in the weak formulation and integrate afterward over the time interval $(0,t)$, for $t>0$, to obtain
$\begin{split}\frac{1}{2}\|\theta_{\varepsilon}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}&+\frac{1}{2\varepsilon}\|\theta_{\varepsilon}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon}\|_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}^{2}+\alpha\|\llbracket\theta_{\varepsilon}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\\ &\leq C\|f_{\varepsilon}\|_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\|\theta_{\varepsilon}\|_{L^{2}((0,t);H^{1}(\Omega^{s}_{\varepsilon}))}+\frac{1}{2}\|\theta_{\varepsilon,0}\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}+\frac{1}{2\varepsilon}\|\theta_{\varepsilon,0}\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}.\end{split}$
Applying Young’s inequality followed by Gronwall’s inequality leads to the desired estimate (3.3), where $C$ is independent of $\varepsilon$. ∎
Remark 3.
Note, the condition in (2.3) is used in the previous Lemma to derive that $(v\cdot\nabla\psi,\psi)_{\Omega^{f}_{\varepsilon}}\geq 0$, this is not a necessary condition to derive coercivity and carry out the energy estimates. Given the trace inequality and the compact embedding of $H^{1}(\Omega^{f}_{\varepsilon})$ into $L^{2}(\partial\Omega^{f}_{\varepsilon})$, the terms stemming from the convection terms (after integration by part) could be included in the already present gradient terms and $H^{1}$-norm. But (2.3) is a reasonable assumption in our case and let us simplify the notation.
Lemma 5.
Let the same assumptions be fulfilled as in Lemma 4 and let $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ be the solution of Eq. 3.1. Then there exists a $C>0$ independent of $\varepsilon$ such that $\|\theta_{\varepsilon}\|_{L^{\infty}(S\times\Omega)}\leq C.$
Proof.
For any $k>0$ construct the functions
$\theta_{\varepsilon,k}^{s}\coloneqq\begin{cases}\theta_{\varepsilon}^{s}-k\quad&\text{if }\theta_{\varepsilon}^{s}>k,\\ 0\quad&\text{if }|\theta_{\varepsilon}^{s}|\leq k,\\ \theta_{\varepsilon}^{s}+k\quad&\text{if }\theta_{\varepsilon}^{s}<-k,\end{cases}\quad\text{and}\quad\theta_{\varepsilon,k}^{f}\coloneqq\begin{cases}\theta_{\varepsilon}^{f}-k\quad&\text{if }\theta_{\varepsilon}^{f}>k,\\ 0\quad&\text{if }|\theta_{\varepsilon}^{f}|\leq k,\\ \theta_{\varepsilon}^{f}+k\quad&\text{if }\theta_{\varepsilon}^{f}<-k,\end{cases}$
and the sets
$A_{\varepsilon,k}^{s}(\tau)\coloneqq\left\{x\in\Omega_{\varepsilon}^{s}:|\theta_{\varepsilon}^{s}(\tau,x)|>k\right\}\quad\text{and}\quad A_{\varepsilon,k}^{f}(\tau)\coloneqq\left\{x\in\Omega_{\varepsilon}^{f}:|\theta_{\varepsilon}^{f}(\tau,x)|>k\right\}.$
Note that it holds $\partial_{t}\theta_{\varepsilon,k}^{s}=\chi_{A_{\varepsilon,k}^{s}}\partial_{t}\theta_{\varepsilon}^{s}$ and $\nabla\theta_{\varepsilon,k}^{s}=\chi_{A_{\varepsilon,k}^{s}}\nabla\theta_{\varepsilon}^{s}$ (similiar for $\theta_{\varepsilon,k}^{f}$). By Assumption (A3) we set $k>C_{0}=\sup_{\varepsilon>0}\|\theta_{\varepsilon,0}\|_{L^{\infty}(\Omega)}$. Next, we take $\theta_{\varepsilon,k}=(\theta_{\varepsilon,k}^{s},\theta_{\varepsilon,k}^{f})$ as a test function in Eq. 3.1. Integrating over the time interval $(0,t)\subset S$, we obtain:
$\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}$ $+\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}$
$+\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}+\alpha\|\llbracket\theta_{\varepsilon,k}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}=(f_{\varepsilon},\theta_{\varepsilon,k}^{s})_{(0,t)\times\Gamma_{\varepsilon}}$
Since $\nabla\theta_{\varepsilon,k}^{f}=\chi_{A_{\varepsilon,k}^{f}}\nabla\theta_{\varepsilon}^{f}$, the convection is
$\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}=\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}.$
Eq. 3.5 and $\operatorname{div}{v}=0$ yield
$\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}=\frac{1}{2\varepsilon}(v\cdot\nu\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\partial\Omega^{f}_{\varepsilon}}=\frac{1}{2\varepsilon}(v\cdot\nu\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\partial\Sigma^{\text{out}}_{\varepsilon}}\geq 0.$
Therefore we obtain
$\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}$ $+\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}$
$+\alpha\|\llbracket\theta_{\varepsilon,k}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\leq(f_{\varepsilon},\theta_{\varepsilon,k}^{s})_{(0,t)\times\Gamma_{\varepsilon}}.$
By Assumption (A4) and the trace estimate in Lemma 2 the source term on the right-hand side can be estimated by
$\begin{split}\left|(f_{\varepsilon},\theta_{\varepsilon,k}^{s})_{(0,t)\times\Gamma_{\varepsilon}}\right|&\leq\|f_{\varepsilon}\|_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}\int_{0}^{t}\int_{\Gamma_{\varepsilon}}|\theta_{\varepsilon,k}^{s}|\,\mathrm{d}x\,\mathrm{d}\tau\\ &\leq C\|f_{\varepsilon}\|_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}\left(\|\theta_{\varepsilon,k}^{s}\|_{L^{1}((0,t)\times\Omega^{s}_{\varepsilon})}+\|\nabla\theta_{\varepsilon,k}^{s}\|_{L^{1}((0,t)\times\Omega^{s}_{\varepsilon})}\right)\\ &\leq\|f_{\varepsilon}\|^{2}_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}C_{\delta}\|1\|_{L^{2}((0,t)\times A_{\varepsilon,k}^{s})}^{2}+C\|\theta_{\varepsilon,k}^{s}\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}+\delta\|\nabla\theta_{\varepsilon,k}^{s}\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}.\end{split}$ (3.6)
In the last step, Young’s inequality with a $\delta>0$ was used. We choose $\delta$ such that we can include the gradient term on the left side and additionally assume $k\geq\|f_{\varepsilon}\|_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}$, to arrive at
$\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}$ $+\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}$
$\leq C\left[\|\theta_{\varepsilon,k}^{s}\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|A_{\varepsilon,k}^{s}(\tau)|\,\mathrm{d}\tau\right]$
Isolating the terms that are connected to the temperature of the solid, we have
$\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}+\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}$ $\leq C\left[\|\theta_{\varepsilon,k}^{s}\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|A_{\varepsilon,k}^{s}(\tau)|\,\mathrm{d}\tau\right].$ (3.7)
With the transformation $\Phi$ from Lemma 1 we can switch to the fixed domain $\Omega$. Note that $x\in\Omega\setminus\Phi(A_{\varepsilon,k}^{s}(\tau))$ iff $\theta_{\varepsilon,k}^{s}\circ\Phi(x,\tau)=0$ for a.a. $\tau\in(0,t)$. And since $\det{(\nabla\Phi)}(x)=({H-\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})})/{H}>0$, we conclude from Eq. 3.7, that also
$\|(\theta_{\varepsilon,k}^{s}\circ\Phi)(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}+\|\nabla(\theta_{\varepsilon,k}^{s}\circ\Phi)\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}$ $\leq C\left[\|\theta_{\varepsilon,k}^{s}\circ\Phi\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|\Phi(A_{\varepsilon,k}^{s}(\tau))|\,\mathrm{d}\tau\right].$
By Gronwall’s inequality, we arrive at
$\|\theta_{\varepsilon,k}^{s}\circ\Phi\|_{L^{\infty}(S;L^{2}(\Omega))\cap L^{2}(S;H^{1}(\Omega))}\leq Ck\left(\int_{S}|\Phi(A_{\varepsilon,k}^{s}(\tau))|\,\mathrm{d}\tau\right)^{\frac{1}{2}}$
The domain $\Omega$ is independent of $\varepsilon$, therefore applying the results of [41, Chapter 2, Theorem 6.1] we obtain $\|\theta^{s}_{\varepsilon}\circ\Phi\|_{L^{\infty}(S\times\Omega)}\leq C$, where $C>0$ can be chosen independent of $\varepsilon$. By the properties of $\Phi$ this immediately leads to $\|\theta^{s}_{\varepsilon}\|_{L^{\infty}(S\times\Omega^{s}_{\varepsilon})}\leq C$ and by the properties of the trace also to $|\theta^{s}_{\varepsilon|\Gamma_{\varepsilon}}|=|(\theta^{s}_{\varepsilon}\circ\Phi)_{|\Sigma}|\leq C$.
For the estimate in the fluid domain, test with $(0,\theta^{f}_{\varepsilon,k})$, with similar computations as above we arrive at
$\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}+\alpha\|\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}=(\theta^{s}_{\varepsilon},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Gamma_{\varepsilon}}$
The right-hand side can bounded by the same steps as in Eq. 3.6. Note, that the trace estimate from Lemma 2 introduces an additional $\nicefrac{{1}}{{\varepsilon}}$. Setting $k\geq\|\theta^{s}_{\varepsilon}\|_{L^{\infty}(S\times\Omega^{s}_{\varepsilon})}$, we have
$\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\frac{1}{\varepsilon}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}\leq\frac{C}{\varepsilon}\left[\|\theta_{\varepsilon,k}^{f}\|_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|A_{\varepsilon,k}^{f}(\tau)|\,\mathrm{d}\tau\right].$
With the extension operator $\mathbb{E}_{\varepsilon}$ from Lemma 1, we transform to the non-rough thin layer. Define $\mathbb{E}_{\varepsilon}{A}^{f}_{\varepsilon,k}(\tau)=\{x\in\Omega_{\varepsilon}:\mathbb{E}_{\varepsilon}\theta^{f}_{e,k}(t,x)\neq 0\}$, clearly $|A^{f}_{\varepsilon,k}(\tau)|\leq|\mathbb{E}_{\varepsilon}A^{f}_{\varepsilon,k}(\tau)|$. Multiplying by $\varepsilon$, we obtain that on the non-rough thin layer, it holds
$\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega_{\varepsilon})}+\kappa^{f}\|\nabla\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega_{\varepsilon})}\leq C\left[\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{2}((0,t)\times\Omega_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|\mathbb{E}_{\varepsilon}A_{\varepsilon,k}^{f}(\tau)|\,\mathrm{d}\tau\right],$
and consequently by Gronwall’s inequality
$\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S;L^{2}(\Omega_{\varepsilon}))\cap L^{2}(S;H^{1}(\Omega_{\varepsilon}))}\leq Ck\left(\int_{S}|\mathbb{E}_{\varepsilon}A_{\varepsilon,k}^{f}(\tau)|\,\mathrm{d}\tau\right)^{\frac{1}{2}}.$
Applying again [41, Chapter 2, Theorem 6.1], we now have to be careful regarding the $\varepsilon$ dependency of the estimate. First we obtain
$\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega_{\varepsilon})}\leq C\left(1+C\beta_{\varepsilon}^{\frac{d+2}{2}}|S|^{\frac{1}{2}}|\Omega_{\varepsilon}|^{\frac{1}{2}}\right)\leq C\left(1+C\varepsilon^{\frac{1}{2}}\beta_{\varepsilon}^{\frac{d+2}{2}}\right),$
where $\beta_{\varepsilon}$ is the embedding constant from Lemma 3. By Lemma 3, $\beta_{\varepsilon}=\mathcal{O}(\varepsilon^{-\frac{1}{d+2}})$. Therefore $\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega_{\varepsilon})}\leq C$ for $C>0$ independent of $\varepsilon$. Lastly, since $\|\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega^{f}_{\varepsilon})}=\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega^{f}_{\varepsilon})}\leq C$, we have the desired estimate. ∎
Lemma 6.
Let the same assumptions be fulfilled as in Lemma 4 and let $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ be the solution of Eq. 3.1, with $\partial_{t}\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$. If $\|v_{\varepsilon}\|_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\leq C\sqrt{\varepsilon}$, then there exists a $C>0$ independent of $\varepsilon$ such that
$\begin{split}\|\partial_{t}\theta_{\varepsilon}\|_{L^{2}(S;H^{1}(\Omega^{s}_{\varepsilon})^{\prime})}\leq C\quad\text{ and }\quad\|\partial_{t}\theta_{\varepsilon}\|_{L^{2}(S;H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime})}\leq C\sqrt{\varepsilon}.\end{split}$ (3.8)
Proof.
By testing Eq. 3.1 with arbitrary $\varphi\in L^{2}(S;\Theta_{\varepsilon})$ one obtains for almost all $t\in S$
$\begin{split}\bigl{|}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}&+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}\bigr{|}\\ &\leq|(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{s}_{\varepsilon}}|+\frac{1}{\varepsilon}|(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{f}_{\varepsilon}}|+\alpha|(\llbracket\theta_{\varepsilon}\rrbracket,\llbracket\varphi\rrbracket)_{\Gamma_{\varepsilon}}|+|(f_{\varepsilon},\varphi^{s})_{\Gamma_{\varepsilon}}|+|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}|\\ &\leq\kappa^{s}\|\nabla\theta_{\varepsilon}\|_{L^{2}(\Omega^{s}_{\varepsilon})}\|\nabla\varphi^{g}\|_{L^{2}(\Omega^{s}_{\varepsilon})}+\frac{\kappa^{f}}{\varepsilon}\|\nabla\theta_{\varepsilon}\|_{L^{2}(\Omega^{f}_{\varepsilon})}\|\nabla\varphi^{f}\|_{L^{2}(\Omega^{f}_{\varepsilon})}+\alpha\|\llbracket\theta_{\varepsilon}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}(\|\varphi^{f}\|_{L^{2}(\Gamma_{\varepsilon})}+\|\varphi^{s}\|_{L^{2}(\Gamma_{\varepsilon})})\\ &\hskip 20.0pt+C\|f_{\varepsilon}\|_{L^{2}(\Gamma_{\varepsilon})}\|\varphi^{s}\|_{H^{1}(\Omega^{s}_{\varepsilon})}+|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}|.\end{split}$
Choosing $\varphi^{f}=0$, integrating over the time interval $S$, followed by Hölder’s inequality and applying (3.3) gives the desired estimate for $\partial_{t}\theta_{\varepsilon}^{s}$.
For estimating $\partial_{t}\theta^{f}_{\varepsilon}$ we set $\varphi^{s}=0$ then the only problematic part is the convection term (the trace over $\Gamma_{\varepsilon}$ can be estimated by Lemma 2). By Eq. 3.5 and $\operatorname{div}{v_{\varepsilon}}=0$
$\begin{split}\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi^{f})_{S\times\Omega^{f}_{\varepsilon}}=-\frac{1}{\varepsilon}(v_{\varepsilon}\theta_{\varepsilon},\nabla\varphi^{f})_{S\times\Omega^{f}_{\varepsilon}}+\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}.\end{split}$
For the first term on the right-hand side, we use the results from Lemma 5 and the assumption on $v_{\varepsilon}$ to obtain
$\left|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\theta_{\varepsilon},\nabla\varphi^{f})_{S\times\Omega^{f}_{\varepsilon}}\right|\leq\frac{1}{\varepsilon}\|\theta\|_{L^{\infty}(S\times\Omega^{f}_{\varepsilon})}\|v_{\varepsilon}\|_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\|\nabla\varphi^{f}\|_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\leq\frac{C}{\sqrt{\varepsilon}}\|\varphi^{f}\|_{L^{2}(S;H^{1}(\Omega^{f}_{\varepsilon}))}.$
For the second term, we again use our choice of boundary conditions, leading to
$\begin{split}\left|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}\right|\leq\frac{1}{\varepsilon}\|v_{\varepsilon}\|_{L^{\infty}(S\times\partial\Omega^{f}_{\varepsilon})}\|\theta_{\varepsilon}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\|\phi^{f}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}.\end{split}$
With Lemma 2.$(iii)$ and the estimates of $\theta^{f}_{\varepsilon}$ from Lemma 4, we obtain
$\begin{split}\left|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}\right|\leq\frac{C}{\varepsilon}\|\theta_{\varepsilon}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\|\varphi^{f}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\leq\frac{C}{\sqrt{\varepsilon}}\|\varphi^{f}\|_{L^{2}(S;H^{1}(\Omega^{f}_{\varepsilon}))}.\end{split}$
Multiplying the complete inequality by $\varepsilon$ we get the desired result. ∎
The additional assumption $\|v_{\varepsilon}\|_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\leq C\sqrt{\varepsilon}$ is not a new requirement to the problem, as the solution of the Stokes problem will fulfill this estimate naturally.
Lemma 7.
Let the Assumptions (A2) and (A5) be fulfilled. Then for all $\varepsilon>0$ and $\psi_{\varepsilon}\in L^{2}(S\times\Omega^{f}_{\varepsilon})$ there exists a unique solution $(u_{\varepsilon},p_{\varepsilon})\in V_{\varepsilon}\times Q_{\varepsilon}$ of Eq. 3.2. In addition, it holds
$\frac{1}{\sqrt{\varepsilon}}\|u_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\sqrt{\varepsilon}\|\nabla u_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}\leq C\quad\text{ and }\quad\sqrt{\varepsilon}\|p_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}\leq\frac{C}{\varepsilon}$ (3.9)
for an $C<\infty$ independent of $\varepsilon$ and $\psi_{\varepsilon}$.
Proof.
Given the bounds of the viscosity (A2) and Assumption (A5), this is analog to the steps in [12, Section 3.2 and 3.4]. ∎
We introduce the operators $\mathbb{T}_{\varepsilon}:V_{\varepsilon}\to L^{2}(S\times\Omega)$, which maps a prescribed velocity field to the corresponding temperature field via Eq. 3.1, and the Stokes operator $\mathbb{S}_{\varepsilon}:L^{2}(S\times\Omega)\to V_{\varepsilon}$ of Eq. 3.2, which returns the velocity field for a given temperature distribution. We are looking for fixed points of the operator
$\mathbb{F}_{\varepsilon}:L^{2}(S\times\Omega)\to L^{2}(S\times\Omega),\quad\psi_{\varepsilon}\mapsto\mathbb{T}_{\varepsilon}\left(\mathbb{S}_{\varepsilon}(\psi_{\varepsilon})\right).$ (3.10)
and point out that any such fixed point is automatically a weak solution to the full problem. In the following, we first investigate the continuity properties of $\mathbb{T}_{\varepsilon}$ and $\mathbb{S}_{\varepsilon}$.
Lemma 8.
Let the Assumptions (A2) and (A5) be fulfilled. Given a sequence $\{\psi_{k}\}_{k\in\mathbb{N}}\subset L^{2}(S\times\Omega)$ with $\psi_{k}\xrightarrow[]{k\to\infty}\psi$ in $L^{2}(S\times\Omega)$, it holds
$\mathbb{S}_{\varepsilon}(\psi_{k})\xrightharpoonup{k\to\infty}\mathbb{S}_{\varepsilon}(\psi)\quad\text{ in }\quad L^{2}(S;H^{1}(\Omega^{f}_{\varepsilon})^{d}).$
Proof.
For each $\psi_{k}$, there exists a unique solution $(u_{k},p_{k})\in V_{\varepsilon}\times Q_{\varepsilon}$ with $\|u_{k}\|_{L^{\infty}(S;H^{1}(\Omega^{f}_{\varepsilon}))}+\|p_{k}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon})}\leq C_{\varepsilon}$, where $C_{\varepsilon}$ denotes a constant that may depend on $\varepsilon$. Given this bound, we can find a $u\in L^{2}(S;H^{1}(\Omega^{f}_{\varepsilon})^{d})$ and a $p\in L^{2}(S\times\Omega^{f}_{\varepsilon})$ such that there is a subsequence of $(u_{k},p_{k})$, still denoted with $k$, that weakly converges:
$u_{k}\rightharpoonup u\quad\text{in }L^{2}(S;H^{1}(\Omega^{f}_{\varepsilon})^{d})\quad\text{ and }\quad p_{k}\rightharpoonup p\quad\text{in }L^{2}(S;\Omega^{f}_{\varepsilon}).$
Given the Lipschitz continuity and bounds of $\mu$, it holds $\mu(\psi_{k})\to\mu(\psi)$ in $L^{2}(S\times\Omega)$. Therefore, we obtain the weak convergence
$\mu(\psi_{k})\nabla u_{k}\rightharpoonup\mu(\psi)\nabla u\quad\text{in }L^{1}(S\times\Omega^{f}_{\varepsilon})^{d}.$
Since the viscosity is in $L^{\infty}(S\times\Omega^{f}_{\varepsilon})$, we have that $\mu(\psi_{k})\nabla u_{k}$ and $\mu(\psi)\nabla u$ are bounded in $L^{2}(S\times\Omega^{f}_{\varepsilon})^{d}$. This lets us conclude
$\mu(\psi_{k})\nabla u_{k}\rightharpoonup\mu(\psi)\nabla u\quad\text{in }L^{2}(S\times\Omega^{f}_{\varepsilon})^{d},$
for a subsequence. We are now able to pass to the limit in the weak formulation and obtain
00 0 $=\lim_{k\to\infty}\left[(\mu(\psi_{k})\nabla u_{k},\nabla v)_{S\times\Omega^{f}_{\varepsilon}}-(p_{k},\operatorname{div}{(v)})_{S\times\Omega^{f}_{\varepsilon}}\right]=(\mu(\psi)\nabla u,\nabla v)_{S\times\Omega^{f}_{\varepsilon}}-(p,\operatorname{div}{(v)})_{S\times\Omega^{f}_{\varepsilon}},$
for all $v\in H^{1}_{0}(\Omega^{f}_{\varepsilon})^{d}$. By localization in time, we see that, for almost all $t\in S$, $(u,p)$ is the solution to the Stokes problem (3.2) in $V_{\varepsilon}\times Q_{\varepsilon}$ corresponding to $\psi$. Because the solution is unique, the whole sequence weakly converges to the same limit. ∎
Lemma 9.
Let Assumptions (A1) and (A3)(A5) be satisfied. Given a sequence $\{u_{k}\}_{k\in\mathbb{N}}\subset V_{\varepsilon}$ and $u\in V_{\varepsilon}$ such that $\sup_{k\in\mathbb{N}}\|u_{k}\|_{V_{\varepsilon}}<\infty$ and $u_{k}\xrightharpoonup{k\to\infty}u$ in $L^{2}(S;H^{1}(\Omega^{f}_{\varepsilon}))$, it holds
$\mathbb{T}_{\varepsilon}(u_{k})\xrightarrow[]{k\to\infty}\mathbb{T}_{\varepsilon}(u)\quad\text{ in }\quad L^{2}(S\times\Omega).$
Proof.
Given the assumptions, the sequences $\{\theta_{k}\}_{k\in\mathbb{N}}:=\{\mathbb{T}_{\varepsilon}(u_{k})\}_{k\in\mathbb{N}}\subset L^{2}(S;\Theta_{\varepsilon})$ and $\{\partial_{t}\theta_{k}\}_{k\in\mathbb{N}}\subset L^{2}(S;\Theta_{\varepsilon}^{\prime})$ are bounded uniformly in $k\in\mathbb{N}$ (see Lemma 4). Therefore, there is a subsequence, still indexed with $k$, and a weak limit $\theta\in L^{2}(S;\Theta_{\varepsilon})$ with $\partial_{t}\theta\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$ such that $\theta_{k}\rightharpoonup\theta$ in $L^{2}(S;\Theta_{\varepsilon})$ and $\partial_{t}\theta_{k}\rightharpoonup\partial_{t}\theta$ in $L^{2}(S;\Theta_{\varepsilon}^{\prime})$.
Note that the embeddings $H^{1}(\Omega^{j}_{\varepsilon})\hookrightarrow L^{2}(\Omega^{j}_{\varepsilon})$ are compact and the embeddings $L^{2}(\Omega^{j}_{\varepsilon})\hookrightarrow H^{1}(\Omega^{j}_{\varepsilon})^{\prime}$ are continuous for $j=f,g$. Therefore, we obtain strong convergence $\theta_{k}\to\theta$ in $L^{2}(S\times\Omega)$ by the Aubin-Lions lemma.
It remains to be shown that $\theta=\mathbb{T}_{\varepsilon}(u)$. Looking at the weak form Eq. 3.1, we can immediately pass to the limit in most terms with only the convection term as a potential issue. Now, since $\theta_{k}\to\theta$ strongly in $L^{2}(S\times\Omega_{\varepsilon}^{f})$ and $u_{k}\rightharpoonup u$ weakly in $L^{2}(S\times\Omega_{\varepsilon}^{f})^{d}$, we have $u_{k}\theta_{k}\rightharpoonup u\theta$ weakly in $L^{1}(S\times\Omega_{\varepsilon}^{f})^{d}$. Via the embedding $H^{1}(\Omega^{f}_{\varepsilon})\hookrightarrow L^{4}(\Omega^{f}_{\varepsilon})$ and $\sup_{k\in\mathbb{N}}\|u_{k}\|_{V_{\varepsilon}}<\infty$, we also have
$\|u_{k}\theta_{k}\|_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\leq C_{\varepsilon}\|u_{k}\|_{L^{\infty}(S;H^{1}(\Omega^{f}_{\varepsilon}))}\|\theta_{k}\|_{L^{2}(S;H^{1}(\Omega^{f}_{\varepsilon}))}\leq C_{\varepsilon}$
where $C_{\varepsilon}>0$ may depend on $\varepsilon$ but not on $k$. This implies that, at least along a subsequence, $u_{k}\theta_{k}$ also converges weakly in $L^{2}(S\times\Omega^{f}_{\varepsilon})^{d}$ which then implies $u_{k}\theta_{k}\rightharpoonup u\theta$ in $L^{2}(S\times\Omega^{f}_{\varepsilon})^{d}$.
This allows us to pass to the limit $k\to\infty$ in the weak formulation and deduce that $\theta$ is the solution corresponding to the velocity field $u$, i.e., $\theta=\mathbb{T}_{\varepsilon}(u)$. Finally, since the solution to the temperature problem is unique, the whole sequence converges. ∎
Combining the previous two lemmas, we obtain the continuity of the concatenated operator $\mathbb{F}_{\varepsilon}$.
Corollary 1.
The operator $\mathbb{F}_{\varepsilon}$ defined in Eq. 3.10 is continuous from $L^{2}(S\times\Omega)$ to $L^{2}(S\times\Omega)$.
With the continuity, we can now tackle the existence of a fixed point.
Theorem 1 (Existence).
Let Assumptions (A1)-(A5) be satisfied. Then the operator $\mathbb{F}_{\varepsilon}$ defined in Eq. 3.10 has a fixed point in $L^{2}(S;\Theta_{\varepsilon})$.
Proof.
We observe that $L^{2}(S;\Theta_{\varepsilon})\subset L^{2}(S\times\Omega)$ and $\mathbb{F}_{\varepsilon}[L^{2}(S\times\Omega)]\subset L^{2}(S;\Theta_{\varepsilon})$ by Lemma 4 and 7. Furthermore, $\partial_{t}\psi\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$ for all $\psi\in\mathbb{F}_{\varepsilon}[L^{2}(S\times\Omega)]$ with uniform bounds from Lemma 4 and 6 which are independent of $\psi$. As in the proof of Lemma 9, we can use the Aubin-Lions lemma to conclude that $\mathbb{F}_{\varepsilon}[L^{2}(S\times\Omega)]$ is compact in $L^{2}(S\times\Omega)$. Finally, $\mathbb{F}_{\varepsilon}$ is $L^{2}(S\times\Omega)$-continuous (Corollary 1). Therefore, we apply Schauder’s fixed point theorem and obtain the existence of at least one fixed point of the operator $\mathbb{F}_{\varepsilon}$ in $L^{2}(S;\Theta)$. ∎
The previous theorem ensures that, under the given assumptions, there exists a weak solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})$ of the problem (2.1) for all $\varepsilon>0$. By construction of the operator $\mathbb{F}_{\varepsilon}$ the fixed point also fulfills the estimates of Lemma 4 and Lemma 7. To ensure uniqueness, further assumptions on the regularity of the solution are required, as denoted in the following theorem.
Theorem 2 (Uniqueness).
Let Assumptions (A1)(A5) be satisfied. If there is a solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})\times V_{\varepsilon}\times Q_{\varepsilon}$ of the coupled system Eq. 3.1 and (3.2), such that
  • if $d=2:$ $\nabla\theta_{\varepsilon}\in L^{\infty}(S;L^{1+\delta}(\Omega_{\varepsilon}^{f})^{2})$ and $\nabla u_{\varepsilon}\in L^{\infty}(S;L^{2+\delta}(\Omega_{\varepsilon}^{f})^{2\times 2})$,
  • if $d=3:$ $\nabla\theta_{\varepsilon}\in L^{\infty}(S;L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f})^{3})$ and $\nabla u_{\varepsilon}\in L^{\infty}(S;L^{3+\delta}(\Omega_{\varepsilon}^{f})^{3\times 3})$,
for arbitrary $\delta>0$. Then, this is the only solution to the problem.
Proof.
We provide the proof in the case $d=3$, the other case follows by the same steps but with a better estimate in the Gagliardo-Nierenberg inequality. Suppose there are two sets of solutions $(\theta_{\varepsilon}^{i},u_{\varepsilon}^{i},p_{\varepsilon}^{i})\in L^{2}(S;\Theta_{\varepsilon})\times V_{\varepsilon}\times Q_{\varepsilon}$ ($i=1,2$) with their differences denoted by $(\overline{\theta_{\varepsilon}},\overline{u_{\varepsilon}},\overline{p_{\varepsilon}})$. Furthermore, let $(\theta_{\varepsilon}^{1},u_{\varepsilon}^{1})$ satisfy the additional gradient regularity. Using the weak formulations, it holds
$\langle\partial_{t}\overline{\theta_{\varepsilon}},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+(\kappa^{s}\nabla\overline{\theta_{\varepsilon}},\nabla\varphi)_{\Omega^{s}_{\varepsilon}}+\frac{1}{\varepsilon}\langle\partial_{t}\overline{\theta_{\varepsilon}},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}(\kappa^{f}\nabla\overline{\theta_{\varepsilon}},\nabla\varphi)_{\Omega^{s}_{\varepsilon}}\\ +\frac{1}{\varepsilon}(\overline{u_{\varepsilon}}\cdot\nabla{\theta_{\varepsilon}}^{1},\varphi)_{\Omega^{f}_{\varepsilon}}+\frac{1}{\varepsilon}(u_{\varepsilon}^{2}\cdot\nabla\overline{\theta_{\varepsilon}},\varphi)_{\Omega^{f}_{\varepsilon}}+\alpha(\llbracket\overline{\theta_{\varepsilon}}\rrbracket,\llbracket\varphi\rrbracket)_{\Gamma_{\varepsilon}}=0,$
$\begin{split}&(\mu(\theta_{\varepsilon}^{2})\nabla\overline{u_{\varepsilon}},\nabla v)_{\Omega^{f}_{\varepsilon}}+((\mu(\theta_{\varepsilon}^{2})-\mu(\theta_{\varepsilon}^{1})))\nabla u_{\varepsilon}^{1},\nabla v)_{\Omega^{f}_{\varepsilon}}-(\overline{p_{\varepsilon}},\operatorname{div}{v})_{\Omega^{f}_{\varepsilon}}=0.\end{split}$
for all test functions $(\varphi,v)\in L^{2}(S;\Theta_{\varepsilon})\times H^{1}_{0}(\Omega^{f}_{\varepsilon})^{d}$ and almost all $t\in S$. Now, choosing $(\varphi,v)=(\overline{\theta_{\varepsilon}},\overline{u_{\varepsilon}})$ as test functions, we obtain
$\frac{1}{2}\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+c_{g}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\frac{1}{2\varepsilon}\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\frac{c_{f}}{\varepsilon}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\alpha\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Gamma_{\varepsilon})}^{2}+\frac{1}{\varepsilon}(u_{\varepsilon}^{2}\cdot\nabla\overline{\theta_{\varepsilon}},\overline{\theta_{\varepsilon}})_{\Omega^{f}_{\varepsilon}}\\ \leq-\frac{1}{\varepsilon}(\overline{u_{\varepsilon}}\cdot\nabla\theta_{\varepsilon}^{1},\overline{\theta_{\varepsilon}})_{\Omega^{f}_{\varepsilon}},$
$\begin{split}&\underline{\mu}\|\nabla\overline{u_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}\leq((\mu({\theta_{\varepsilon}}^{1})-\mu({\theta_{\varepsilon}}^{2}))\nabla u_{\varepsilon}^{1},\nabla\overline{u_{\varepsilon}})_{\Omega^{f}_{\varepsilon}}.\end{split}$
Take $r=\nicefrac{{2(3+\delta)}}{{1+\delta}}$, then $1=\nicefrac{{1}}{{2}}+\nicefrac{{1}}{{r}}+\nicefrac{{1}}{{3+\delta}}$. Using the Lipschitz property of the viscosity and Hölder’s and the Poincaré inequality, we are led to
$\|\overline{u_{\varepsilon}}\|_{H^{1}(\Omega_{\varepsilon}^{f})}^{2}\leq C\|\overline{\theta_{\varepsilon}}\|_{L^{r}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}.$ (3.11)
For the heat problem, we first note that
$(u_{\varepsilon}^{2}\cdot\nabla\overline{\theta_{\varepsilon}},\overline{\theta_{\varepsilon}})_{\Omega^{f}_{\varepsilon}}\geq 0$
via Eq. 3.5 resulting in (for $q=\nicefrac{{18+12\delta}}{{3+10\delta}}$)
$\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\llbracket\overline{\theta_{\varepsilon}}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}^{2}\\ \leq C_{\varepsilon}\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{L^{q}(\Omega_{\varepsilon}^{f})}\|\nabla{\theta_{\varepsilon}}^{1}\|_{{L^{\nicefrac{{3}}{{2}}+\delta}}(\Omega_{\varepsilon}^{f})}.$ (3.12)
Summing Eqs. 3.11 and 3.12 and integrating over time, we get
$\|\overline{\theta_{\varepsilon}}(t)\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\overline{\theta_{\varepsilon}}(t)\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\int_{0}^{t}\left(\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\llbracket\overline{\theta_{\varepsilon}}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}^{2}+\|\overline{u_{\varepsilon}}\|_{H^{1}(\Omega_{\varepsilon}^{f})}^{2}\right)\,\mathrm{d}\tau\\ \leq C_{\varepsilon}\int_{0}^{t}\left(\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{L^{q}(\Omega_{\varepsilon}^{f})}\|\nabla{\theta_{\varepsilon}}^{1}\|_{{L^{\nicefrac{{3}}{{2}}+\delta}}(\Omega_{\varepsilon}^{f})}+\|\overline{\theta_{\varepsilon}}\|_{{L^{r}}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}\right)\,\mathrm{d}\tau.$
Focusing on the second term on the right-hand side, we have
$\int_{0}^{t}\|\overline{\theta_{\varepsilon}}\|_{{L^{r}}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}\,\mathrm{d}\tau\leq\|\nabla u_{\varepsilon}^{1}\|_{L^{\infty}(S;{L^{3+\delta}}(\Omega_{\varepsilon}^{f}))}^{2}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(S;{L^{r}}(\Omega_{\varepsilon}^{f}))}^{2}.$
The Gagliardo-Nierenberg interpolation inequality with $\lambda=\nicefrac{{3}}{{3+\delta}}$, yields
$\|\overline{\theta_{\varepsilon}}\|_{L^{2}(S;{L^{r}}(\Omega_{\varepsilon}^{f}))}\leq C\|\nabla\overline{\theta_{\varepsilon}}\|^{\lambda}_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\|\overline{\theta_{\varepsilon}}\|^{1-\lambda}_{L^{2}(S\times\Omega^{f}_{\varepsilon})}.$
Using $\nabla u_{\varepsilon}^{1}\in L^{\infty}(S;L^{3}(\Omega_{\varepsilon}^{f}))^{d}$ and the scaled Young’s inequality leads to
$\int_{0}^{t}\|\overline{\theta_{\varepsilon}}\|_{{L^{r}}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}\,\mathrm{d}\tau\leq\frac{1}{4}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(S\times\Omega_{\varepsilon}^{f})}^{2}+C_{\varepsilon}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(S\times\Omega_{\varepsilon}^{f})}^{2}.$
For the first term on the right-hand side, we use $\nabla\theta_{\varepsilon}^{1}\in L^{\infty}(S;{L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f})})$ to estimate
$\int_{0}^{t}\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{{L^{q}}(\Omega_{\varepsilon}^{f})}\|\nabla{\theta_{\varepsilon}}^{1}\|_{L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f})}\,\mathrm{d}\tau\leq C_{\varepsilon}\|\overline{u_{\varepsilon}}\|_{L^{2}(S;H^{1}(\Omega_{\varepsilon}^{f}))}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(S;L^{6}(\Omega_{\varepsilon}^{f}))}$
which (again using Young’s inequality and the Gagliardo-Nierenberg interpolation) gives
$C_{\varepsilon}\int_{0}^{t}\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{{L^{q}}(\Omega_{\varepsilon}^{f})}\|\nabla{\theta_{\varepsilon}}^{1}\|_{L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f})}\,\mathrm{d}\tau\leq\frac{1}{2}\|\overline{u_{\varepsilon}}\|_{L^{2}(S;H^{1}(\Omega_{\varepsilon}^{f}))}^{2}+\frac{1}{4}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(S\times\Omega_{\varepsilon}^{f})}^{2}+C_{\varepsilon}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(S\times\Omega_{\varepsilon}^{f})}^{2}.$
Putting everything together and utilizing Gronwall’s inequality shows $\overline{u_{\varepsilon}}=0$ and $\overline{\theta_{\varepsilon}}=0$.
In particular, any classical solution must be unique via Theorem 2. We point to an analogous statement for the limit problem (Theorem 4).
4 Homogenization
To pass to the limit $\varepsilon\to 0$ we utilize the concept of two-scale convergence for thin layers first introduced in [46, Definition 4.1] and further developed in following articles, see for example [14, 29] and the references therein. We start by introducing the definition, then consider the existence of limit functions, and then pass to the limit in the weak formulation. Lastly, we analyze the existence and uniqueness of solutions to the effective model.
4.1 Two scale convergence
Definiton 1 (Two-scale convergence on thin domains).
  • i)
    A sequence $u_{\varepsilon}\in L^{2}(S\times\Omega_{\varepsilon})$ is said to weakly two-scale converge to a function $u\in L^{2}(S\times\Sigma\times Y^{d})$ (notation $u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}u$) if
    $\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{S}\int_{\Omega_{\varepsilon}}u_{\varepsilon}(t,x)\varphi\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}x\,\mathrm{d}t=\int_{S}\int_{\Sigma}\int_{Y^{d}}u(t,\tilde{x},y)\varphi(t,\tilde{x},y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$ (4.1)
    for all $\varphi\in C(\overline{S\times\Sigma};C_{\#}(\overline{Y^{d}}))$. If in addition
    $\lim_{\varepsilon\to 0}\frac{1}{\sqrt{\varepsilon}}\|u_{\varepsilon}\|_{L^{2}(S\times\Omega_{\varepsilon})}=\|u\|_{L^{2}(S\times\Sigma\times Y^{d})}$
    then the sequence is said to strongly two-scale converge and we write $u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightarrow}}u$.
  • ii)
    A sequence $u_{\varepsilon}\in L^{2}(S\times\Gamma_{\varepsilon})$ is said to weakly two-scale converge to a function $u\in L^{2}(S\times\Sigma\times\Gamma)$ if
    $\lim_{\varepsilon\to 0}\int_{S}\int_{\Gamma_{\varepsilon}}u_{\varepsilon}(t,x)\varphi\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}\sigma_{x}\,\mathrm{d}t=\int_{S}\int_{\Sigma}\int_{\Gamma}u(t,\tilde{x},y)\varphi(t,\tilde{x},y)\,\mathrm{d}\sigma_{y}\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$ (4.2)
    for all $\varphi\in C(\overline{S\times\Sigma};C_{\#}(\Gamma))$.
Note, that due to the Assumption (A5) it holds for the boundary condition of the velocity
$\frac{1}{\sqrt{\varepsilon}}\|u_{\varepsilon,BC}\|_{L^{2}(\partial\Sigma\times(0,\varepsilon))}=\|u_{BC}\|_{L^{2}(\partial\Sigma\times(0,1))},$
as well as
$\begin{split}\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{\partial\Sigma\times(0,\varepsilon)}u_{\varepsilon,BC}\varphi\left(\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}x&=\lim_{\varepsilon\to 0}\int_{\partial\Sigma}\int_{0}^{1}u_{BC}(\tilde{x},x_{d})\varphi\left(\tilde{x},\nicefrac{{\tilde{x}}}{{\varepsilon}},x_{d}\right)\,\mathrm{d}x_{d}\,\mathrm{d}\tilde{x}\\ &=\int_{\partial\Sigma}\int_{Y^{d}}u_{BC}(\tilde{x},y_{d})\varphi(\tilde{x},y)\,\mathrm{d}y\,\mathrm{d}\tilde{x},\end{split}$ (4.3)
for all $\varphi\in C(\partial\Sigma;C_{\#}(\overline{Y^{d}}))$. The two above properties imply the strong two-scale convergence of the boundary function, $u_{\varepsilon,BC}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}u_{BC}$. Where (given that the underlying domain is rectangular with integer corners) the strong two-scale convergence on the lateral boundary is defined as Definition 1 $i)$, with $\Omega_{\varepsilon}$ and $\Sigma$ replaced by $\partial\Sigma\times(0,\varepsilon)$ and $\partial\Sigma$ respectively.
Given the assumptions on the data and the estimate derived in Section 3.2 we obtain the following limit functions.
Proposition 1 (Existence of limit functions).
Under the Assumption (A1)-(A5), it holds for the weak solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})\times V_{\varepsilon}\times Q_{\varepsilon}$ of Eq. 2.1 that there exists a subsequence still denoted by $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})$ and:
  • i)
    a $\theta^{s}\in L^{2}(S;H^{1}(\Omega))\cap H^{1}(S;H^{1}(\Omega)^{\prime})$ with
    $\chi_{\Omega^{s}_{\varepsilon}}\theta_{\varepsilon}\to\theta^{s}\text{ in }L^{2}(S\times\Omega)\quad\text{ and }\quad\chi_{\Omega^{s}_{\varepsilon}}\nabla\theta_{\varepsilon}\rightharpoonup\nabla\theta^{s}\text{ in }L^{2}(S\times\Omega)^{d},$
  • ii)
    a $\theta^{f}\in L^{2}(S;H^{1}(\Sigma))\cap H^{1}(S;H^{1}(\Sigma)^{\prime})$ and a $\theta_{1}^{f}\in L^{2}(S\times\Sigma;H^{1}_{\#}(Y^{d}))$ with
    $\chi_{\Omega^{f}_{\varepsilon}}\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightarrow}}\chi_{Z}\theta^{f}\quad\text{ and }\quad\chi_{\Omega^{f}_{\varepsilon}}\nabla\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\left(\nabla_{\tilde{x}}\theta^{f}+\nabla_{y}\theta_{1}^{f}\right),$
  • iii)
    a $u\in L^{2}(S\times\Sigma;H^{1}_{\#}(\overline{Z}))^{d}$ with
    $\chi_{\Omega^{f}_{\varepsilon}}u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}u\quad\text{ and }\quad\chi_{\Omega^{f}_{\varepsilon}}\varepsilon\nabla u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\nabla_{y}u.$
    It holds $\operatorname{div}_{y}{(u(\tilde{x},y))}=0$ and $\operatorname{div}_{\tilde{x}}{\left(\int_{Z}u(\tilde{x},y)\,\mathrm{d}y\right)}=0$ for almost all $(\tilde{x},y)\in\Sigma\times Z$. Moreover, $u$ has, for almost all $(t,x)\in S\times\Sigma$, the boundary values $u(t,\tilde{x},\tilde{y},0)=u_{\text{motion}}$ for $\tilde{y}\in Y^{d-1}$ and $u(t,\tilde{x},y)=0$ for $y\in\Gamma$ almost everywhere.
  • iv)
    a $p\in L^{2}_{0}(S\times\Sigma\times Z)$ that is periodic in $Y^{d-1}$ and
    $\varepsilon^{2}\chi_{\Omega^{f}_{\varepsilon}}p_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}p.$
Additionally, on the rough interface, it holds
$\lim_{\varepsilon\to 0}\int_{S}\int_{\Gamma_{\varepsilon}}\alpha(\theta^{s}_{\varepsilon}-\theta^{f}_{\varepsilon})\varphi\,\mathrm{d}\sigma_{x}\,\mathrm{d}t=\int_{S}\int_{\Sigma}\alpha|\Gamma|(\theta^{s}-\theta^{f})\varphi\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$
for all admissible test functions $\varphi$.
Proof.
Given the estimates of Lemma 4 the convergence in $i)$ can be deduced with the same arguments as in [19, Theorem 4].
The aspects $ii)$-$iv)$ are obtained by the estimates of Lemma 4 and Lemma 7 and usual two-scale properties, see for example [14, 25, 29]. For the convergence of the time derivative $\partial_{t}\theta^{f}_{\varepsilon}$ as well as the strong two-scale convergence of $\theta^{f}_{\varepsilon}$, we point to the results of [27, Proposition 5], since we have the same $\varepsilon$-dependent solution estimates. There are already several results for fluid and pressure limits in thin layers, see for example [22, 26]. The general idea for the divergence properties is similar to the case of porous domains in non-thin domains [4, Proposition 1.14]. The last integral identity is also a common equality, see for example [46] for similar setups and the general proof idea. ∎
Given the strong two-scale convergence and the properties of $\mu$ we obtain the following convergence of the viscosity function.
Corollary 2.
Given the Assumptions (A2) it holds $\chi_{\Omega^{f}_{\varepsilon}}\mu(\theta^{f}_{\varepsilon})\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\mu(\theta^{f})$ strongly in $L^{2}(S\times\Omega^{f}_{\varepsilon})$.
4.2 Homogenized equations
Given the previous results concerning the limit of the solution, we can now carry out the homogenization process in the weak formulation. The general procedure when determining the effective model consists of three steps:
  1. 1.
    Use smooth test functions, compatible with the two-scale concept, to pass to the limit $\varepsilon\to 0$.
  2. 2.
    Try to average and decouple the effective systems to simplify the equations and extract cell problems.
  3. 3.
    Apply a density argument to show that the limit also holds for test functions from more general spaces.
Most of the steps are standard for our case, therefore we do not explain everything in detail but only comment on the general behavior and explain the important terms more in-depth. The complete homogenized model is written down at the end of this section in Theorem 3.
Limit procedure for the heat system.
We start with temperature limit inside of $\Omega^{s}_{\varepsilon}$. Given the Assumptions (A6) and (A7) for the convergence of the initial condition and heat source, passing to the limit in the weak formulation can be carried out directly. To be exact, we do not require two-scale convergence inside the bulk domain. We point to [19, Section 5] for a similar setup with a rough interface.
Most of the equations for the temperature limit in $\Omega^{f}_{\varepsilon}$ are also quite standard and we point to [14, 46, 29] for similar setups. As usual for these kinds of diffusion problems, one obtains that the limit $\theta^{f}_{1}$ can be represented by
$\theta^{f}_{1}(t,\tilde{x},y)=\sum_{i=1}^{d-1}\omega_{i}(y)\partial_{x_{i}}\theta^{f}(t,\tilde{x})$
where $\omega_{i}\in H^{1}_{\#}(Z)$ is the weak solution of
$\begin{split}\operatorname{div}_{y}{(\nabla_{y}\omega_{i}+e_{i})}&=0\quad\text{ in }Z,\\ (\nabla_{y}\omega_{i}+e_{i})\cdot n&=0\quad\text{ on }\Gamma\cup\{y\in\partial Z:y_{d}=0\}.\end{split}$ (4.4)
Next, we have a closer look at the boundary values of the limit $\theta^{f}$. Take $\varphi\in C^{\infty}(\overline{S\times\Sigma})$ and $\psi\in C_{\#}^{\infty}(\overline{Z})^{d}$, such that $\operatorname{div}_{y}{\psi}=0$, $\psi(y)=0$, if $y\in\Gamma$ or $y_{d}=0$ and $\varphi_{|\Sigma^{\text{out}}}=0$. Testing with $\varphi(t,\tilde{x})\psi(\nicefrac{{x}}{{\varepsilon}})$ yields
$\frac{1}{\varepsilon}\int_{S}\int_{\Omega^{f}_{\varepsilon}}\theta^{f}_{\varepsilon}(t,x)\operatorname{div}_{x}{\left(\varphi(t,\tilde{x})\psi(\nicefrac{{x}}{{\varepsilon}})\right)}\,\mathrm{d}x\,\mathrm{d}t=-\frac{1}{\varepsilon}\int_{S}\int_{\Omega^{f}_{\varepsilon}}\nabla_{x}\theta^{f}_{\varepsilon}(t,x)\cdot\varphi(t,\tilde{x})\psi(\nicefrac{{x}}{{\varepsilon}})\,\mathrm{d}x\,\mathrm{d}t.$
Passing to the limit on both sides and afterward integrating by parts on the right side, we obtain
$\int_{S}\int_{\Sigma}\int_{Z}\theta^{f}(t,\tilde{x})\operatorname{div}_{\tilde{x}}{(\varphi(t,\tilde{x})\psi(y))}\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t=-\int_{S}\int_{\Sigma}\int_{Z}\left(\nabla_{\tilde{x}}\theta^{f}(t,\tilde{x})+\nabla_{y}\theta^{f}_{1}(t,\tilde{x},y)\right)\cdot\varphi(t,\tilde{x})\psi(y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t$
$=\int_{S}\int_{\Sigma}\int_{Z}\theta^{f}(t,\tilde{x})\operatorname{div}_{\tilde{x}}{(\varphi(t,\tilde{x})\psi(y))}\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t-\int_{S}\int_{\Sigma^{\text{in}}}\theta^{f}(t,\tilde{x})\varphi(t,\tilde{x})\int_{Z}\psi(y)\,\mathrm{d}y\cdot n_{\Sigma}\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t$
$+\int_{S}\int_{\Sigma}\int_{Z}\theta^{f}_{1}(t,\tilde{x},y)\varphi(t,\tilde{x})\operatorname{div}_{y}{(\psi(y))}\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t-\int_{S}\int_{\Sigma}\int_{\partial Z}\theta^{f}_{1}(t,\tilde{x},y)\varphi(t,\tilde{x})\psi(y)\cdot n_{Z}\,\mathrm{d}y\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t.$
Given the properties of $\psi$ and periodicity of $\theta^{f}_{1}$ the last two integrals are 0. So we obtain
$\int_{S}\int_{\Sigma^{\text{in}}}\theta^{f}(t,\tilde{x})\varphi(t,\tilde{x})\int_{Z}\psi(y)\,\mathrm{d}y\cdot n_{\Sigma}\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t=0.$
Since the above quantity has to hold for all $\psi$, we conclude that already
$\int_{S}\int_{\Sigma^{\text{in}}}\theta^{f}(t,\tilde{x})\varphi(t,\tilde{x})\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t=0.$
By a density argument we can conclude that this property has to hold for all $\varphi\in L^{2}(S;H^{1}_{\Sigma^{\text{out}}}(\Sigma))$, which yields that $\theta^{f}=0$ almost everywhere on $S\times\Sigma^{\text{in}}$. Note, that similarly the boundary values for $u$ on $\partial Z$ follow.
The other interesting part is the convection term. Here, Assumption (A5) as well as the strong two-scale convergence of $\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightarrow}}\theta^{f}$ are used to pass to the limit. Take $\varphi_{\varepsilon}=\varphi_{0}(t,\tilde{x})+\varepsilon\varphi_{1}\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)$ with $\varphi_{0}\in C^{\infty}(\overline{S\times\Sigma})$ and $\varphi_{1}\in C^{\infty}(\overline{S\times\Sigma},C_{\#}^{\infty}(\overline{Z}))$, such that $\varphi_{0}(t,\tilde{x})=0$ and $\varphi_{1}(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}})=0$ if $\tilde{x}\in\Sigma^{\text{in}}$. In the case of the convection term, we can integrate by parts to first obtain
$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}(u_{\varepsilon}\nabla\theta_{\varepsilon},\varphi_{\varepsilon})_{S\times\Omega^{f}_{\varepsilon}}$ $=\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\Bigl{[}-(u_{\varepsilon}\theta_{\varepsilon},\nabla_{\tilde{x}}\varphi_{0})_{S\times\Omega^{f}_{\varepsilon}}-\varepsilon(u_{\varepsilon}\theta_{\varepsilon},\nabla_{\tilde{x}}\varphi_{1})_{S\times\Omega^{f}_{\varepsilon}}-(u_{\varepsilon}\theta_{\varepsilon},\nabla_{y}\varphi_{1})_{S\times\Omega^{f}_{\varepsilon}}$
$\hskip 42.67912pt+(u_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi_{0})_{S\times\partial\Omega^{f}_{\varepsilon}}+\varepsilon(u_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi_{1})_{S\times\partial\Omega^{f}_{\varepsilon}}\Bigr{]}$
Utilizing the strong two-scale convergence and (4.3) we can pass to the limit on the right side, producing
$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}(u_{\varepsilon}\nabla\theta_{\varepsilon},\varphi_{\varepsilon})_{S\times\Omega^{f}_{\varepsilon}}$ $=-(u\theta^{f},\nabla_{\tilde{x}}\varphi_{0})_{S\times\Sigma\times Z}-(u\theta^{f},\nabla_{y}\varphi_{1})_{S\times\Sigma\times Z}+(u_{BC}(\tilde{x},y_{d})\cdot n_{\Sigma}\theta^{f},\varphi_{0})_{S\times\partial\Sigma\times Y^{d}}.$
Integration by parts in the second term on the right-hand side, that $\operatorname{div}_{y}(u\theta^{f})=0$, the periodicity and boundary values of $u$ yield $(u\theta^{f},\nabla_{y}\varphi_{1})_{S\times\Sigma\times Z}=0$. Define $\bar{u}(t,\tilde{x})=\int_{Z}u(t,\tilde{x},y)\,\mathrm{d}y$. Then integrating the first term by parts yields
$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}(u_{\varepsilon}\nabla\theta_{\varepsilon},\varphi_{\varepsilon})_{S\times\Omega^{f}_{\varepsilon}}$ $=(\bar{u}\nabla_{\tilde{x}}\theta^{f},\varphi_{0})_{S\times\Sigma}+\left(\bar{u}\cdot n_{\Sigma}-\int_{Y^{d}}u_{BC}(\tilde{x},y_{d})\cdot n_{\Sigma}\,\mathrm{d}y,\theta^{f}\varphi_{0}\right)_{S\times\partial\Sigma}.$
To proceed we now investigate the outer boundary conditions of $\bar{u}$. By the two-scale convergence of the Dirichlet condition $u_{\varepsilon,BC}$, see Eq. 4.3, testing with a $\varphi\in C^{\infty}(\overline{\Sigma})$, independent of $y$ and $t$ (which is reasonable since $u_{\varepsilon}$ two-scale converges a.e. in $S$), yields
00 0 $=\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{\Omega^{f}_{\varepsilon}}\operatorname{div}{(u_{\varepsilon}(x))}\varphi(\tilde{x})\,\mathrm{d}x=-\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{\Omega^{f}_{\varepsilon}}u_{\varepsilon}(x)\nabla\varphi(\tilde{x})\,\mathrm{d}x+\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{\partial\Omega^{f}_{\varepsilon}}u_{\varepsilon}\cdot n_{\varepsilon}\varphi(\tilde{x})\,\mathrm{d}\sigma_{x}$
$=-\int_{\Sigma}\int_{Z}u(\tilde{x},y)\nabla_{\tilde{x}}\varphi(\tilde{x})\,\mathrm{d}y\,\mathrm{d}\tilde{x}+\int_{\partial\Sigma}\left(\int_{0}^{1}u_{BC}(\tilde{x},s)\,\mathrm{d}s\right)\cdot n_{\Sigma}\varphi(\tilde{x})\,\mathrm{d}\sigma_{\tilde{x}}.$
Integration by parts in the last line, lets us conclude that
$\bar{u}(t,\tilde{x})\cdot n_{\Sigma}=\int_{0}^{1}u_{BC}(\tilde{x},s)\,\mathrm{d}s\cdot n_{\Sigma}$ (4.5)
on $\partial\Sigma$ and for almost all $t\in S$. We can conclude that the convection term converges to a convection of the effective quantities. After calculating all limits $\varepsilon\to 0$ in Eq. 3.1, one arrives at the weak form
$\begin{split}\langle\partial_{t}\theta^{s},\varphi\rangle_{L^{2}(S;H^{1}(\Omega)^{\prime})}+|Z|\langle\partial_{t}\theta^{f},\psi\rangle_{L^{2}(S;H^{1}_{\Sigma^{\text{in}}}(\Sigma)^{\prime})}+\kappa^{s}(\nabla\theta^{s},\nabla\varphi)_{S\times\Omega}&+(\tilde{\kappa}\nabla_{\tilde{x}}\theta^{f},\nabla\psi)_{S\times\Sigma}+(\bar{u}\cdot\nabla_{\tilde{x}}\theta^{f},\psi)_{S\times\Sigma}\\ &+\alpha|\Gamma|(\theta^{s}-\theta^{f},\phi-\psi)_{S\times\Sigma}=(\bar{f},\theta^{s})_{S\times\Sigma}\end{split}$ (4.6)
which, by a density arguments, holds for all $(\varphi,\psi)\in L^{2}(S;H^{1}(\Omega)\times H^{1}_{\Sigma^{\text{in}}}(\Sigma))$. Here, the entries of $\tilde{\kappa}$ and $\bar{f}$ are given by
$\tilde{\kappa}_{ij}=\int_{Z}\kappa^{f}(\delta_{ij}+\nabla_{y}\omega_{i}\cdot e_{j})\,\mathrm{d}y,\qquad\bar{f}(t,\tilde{x})=\int_{\Gamma}f(t,\tilde{x},y)\,\mathrm{d}\sigma_{y}.$
Limit procedure for the Stokes system.
Shifting to the Stokes system, we observe that our setup is somewhat close to fluid flows in porous media. The following steps are therefore similar to the classical derivation of the Darcy equation, see [3, 8]. There is also quite a lot of literature considering thin porous layers, which is even closer related to the present setup, for example, [22, 37, 26]. Nevertheless, to better understand the final homogenized model we briefly present the procedure and point out any differences compared to deriving a Darcy equation.
Let $\varphi\in C_{c}^{\infty}(\Sigma;C_{\#}^{\infty}(\overline{Z}))^{d}$ be a smooth function with compact support in $\Sigma$ and $\varphi(\tilde{x},y)=0$ if $y\in\Gamma\cup\{y_{d}=0\}$. We test the weak form Eq. 3.2 with $\varphi_{\varepsilon}\in H_{0}^{1}(\Omega^{f}_{\varepsilon})^{d}$ given by $\varphi_{\varepsilon}(x)=\varepsilon^{k}\varphi(\tilde{x},\nicefrac{{x}}{{\varepsilon}})$ ($k=1,2$) to get (for almost all $(t\in S)$)
$\left(\mu(\theta_{\varepsilon}^{f})\nabla u_{\varepsilon},\left[\varepsilon^{k}\nabla_{\tilde{x}}\varphi+\varepsilon^{k-1}\nabla_{y}\varphi\right](\tilde{x},\nicefrac{{x}}{{\varepsilon}})\right)_{\Omega^{f}_{\varepsilon}}-\left(p_{\varepsilon},\left[\varepsilon^{k}\operatorname{div}_{\tilde{x}}{\varphi}+\varepsilon^{k-1}\operatorname{div}_{y}{\varphi}\right](\tilde{x},\nicefrac{{x}}{{\varepsilon}})\right)_{\Omega^{f}_{\varepsilon}}=0.$
If we set $k=2$ and pass to the limit $\varepsilon\to 0$, we see that
$(p(t,\tilde{x},y),\operatorname{div}_{y}{\varphi}(\tilde{x},y))_{\Sigma\times Z}=0.$
It follows that $p$ is constant in $y$. On the other hand, if we take only functions with $\operatorname{div}_{y}{\varphi}=0$ and $\operatorname{div}_{\tilde{x}}{\left(\int_{Z}\varphi\,\mathrm{d}y\right)}=0$ and set $k=1$, the limit yields
$(\mu(\theta^{f}(t,\tilde{x}))\nabla_{y}u(t,\tilde{x},y),\nabla_{y}\varphi(\tilde{x},y))_{\Sigma\times Z}=0.$ (4.7)
Here, we used Corollary 2. Next, let us introduce
$V=\left\{\varphi\in L^{2}(\Sigma;H^{1}_{\#}(Z))^{d}:\,\begin{array}[]{c}\operatorname{div}_{y}{\varphi}=0,\,\operatorname{div}_{\tilde{x}}{\left(\int_{Z}\varphi\,\mathrm{d}y\right)}=0,\\ \varphi=0\text{ on }\Gamma\cup\{y_{d}=0\}\text{ and }\int_{Z}\varphi\,\mathrm{d}y\cdot n_{\Sigma}=0\text{ on }\partial\Sigma\end{array}\right\}.$
By a density argument Eq. 4.7 also holds for all $\varphi\in V$. By the choice of test functions the pressure $p$ does not appear in (4.7). To recover the pressure, we use the properties of the orthogonal of $V$. With the results from [22, Lemma 5.3 and Remark 5.5.$(i)$], one obtains that
$V^{\perp}=\left\{q=\nabla_{\tilde{x}}q_{0}+\nabla_{y}q_{1}:\,q_{0}\in H^{1}(\Sigma)/\mathbb{R},\,q_{1}\in L^{2}_{0}(\Sigma\times Z)\right\}\subset L^{2}(\Sigma;\mathcal{H}(Z)^{\prime}),$
where $\mathcal{H}(Z)\coloneqq\left\{\varphi\in H^{1}_{\#}(Z)^{d}:\varphi(y)=0\text{ if }y\in\Gamma\cup\{y_{d}=0\}\right\}$. The decomposition into $(q_{0},q_{1})$ is also unique for every $q\in V^{\perp}$. Therefore, there exists a $q=(q_{0},q_{1})\in L^{2}(S;V^{\perp})$ such that $\mu(\theta^{f})\Delta_{y}u=\nabla_{\tilde{x}}q_{0}+\nabla_{y}q_{1}$ (for almost all $t\in S$) understood as an element in $L^{2}(\Sigma;\mathcal{H}(Z)^{\prime})$. Collecting the boundary conditions of $u$ derived in Eq. 4.5 and Proposition 1 as well as the above characterization, we see that
$(\bar{u},u,q_{0},q_{1})\in L^{2}(S;H(\operatorname{div},\Sigma)\times L^{2}(\Sigma;H^{1}(Z))\times H^{1}(\Sigma)/\mathbb{R}\times L^{2}_{0}(\Sigma\times Z))$
are weak solutions to the following two-pressure Stokes model:
[…]
Note that both $u_{\text{motion}}\cdot e_{d}=0$ and Assumption (A5) (which guarantees $\int_{\partial\Sigma}n_{\Sigma}\cdot\int_{0}^{1}u_{BC}=0$) are essential to ensure that the boundary conditions are compatible with the incompressibility requirements. Furthermore, we introduced the new pressure-like functions $q_{0},q_{1}$ that still have no clear connection to the pressure limit $p$ of the microscale model. We first show $p=q_{0}$. To this end, we test the original momentum equation of the micro problem with $\varepsilon\varphi$ for $\varphi\in C_{c}^{\infty}(\Sigma;C_{\#}^{\infty}(\overline{Z}))^{d}$ with $\varphi=0$ on $\Gamma\cup\{y_{d}=0\}$ and $\operatorname{div}_{y}{\varphi}=0$:
$(\mu(\theta^{f})\nabla_{y}u,\nabla_{y}\varphi)_{\Sigma\times Z}-(p,\operatorname{div}_{\tilde{x}}\varphi)_{\Sigma\times Z}=0.$ (4.8)
With a density argument, this also holds for all $\varphi\in L^{2}(\Sigma;\mathcal{H}(Z))$ with $\operatorname{div}_{\tilde{x}}{\left(\int_{Z}\varphi\,\mathrm{d}y\right)}\in L^{2}(\Sigma)$ and $\operatorname{div}_{y}{\varphi}=0$. Since the representation $\mu(\theta^{f})\Delta_{y}u=\nabla_{\tilde{x}}q_{0}+\nabla_{y}q_{1}$ is unique, we deduce from Eq. 4.8 that it needs to hold $p=q_{0}$. We recovered the pressure $p$ for the two-pressure model and therefore replace, in the following, $q_{0}$ by $p$.
By the linearity of the above two-pressure Stokes model, we can further simplify the problem by constructing suitable cell problems. The first set of cell problems are quite standard, see [3] where they are used to characterize the permeability in a porous medium: for $i=1,\dots,d-1$ find $(\xi_{i},\eta_{i})\in H^{1}_{\#}(Z)^{d}\times L^{2}_{0,\#}(Z)$ such that
$-\Delta_{y}\xi_{i}-\nabla_{y}\eta_{i}$ $=e_{i}\quad$ in $\text{in }Z,$
$\operatorname{div}_{y}{(\xi_{i})}$ $=0$ in $\text{in }Z,$ (4.9)
$\xi_{i}$ $=0$ on $\text{on }\Gamma\cup\{y_{d}=0\}.$
Additionally, a cell problem originating from the nonhomogeneous Dirichlet condition at the bottom boundary appears: find $(\xi_{0},\eta_{0})\in H^{1}_{\#}(Z)^{d}\times L^{2}_{0,\#}(Z)$ such that
$-\Delta_{y}\xi_{0}-\nabla_{y}\eta_{0}$ $=0\quad$ in $\text{in }Z,$
$\operatorname{div}_{y}{(\xi_{0})}$ $=0$ in $\text{in }Z,$ (4.10)
$\xi_{0}$ $=0$ on $\text{on }\Gamma,$
$\xi_{0}$ $=u_{\text{motion}}\quad$ on $\text{on }\{y_{d}=0\}.$
Note, the cell problem (4.2) is a well-defined equation and admits a unique solution, given that $u_{\text{motion}}\cdot e_{d}=0$. We have the following property of the cell solutions:
Proposition 2.
For the cell solutions $\xi_{i}\in H^{1}_{\#}(Z)^{d}$ ($i=0,\dots,d-1$) of Section 4.2 and Section 4.2, respectively, it holds
$\int_{Z}\xi_{i}\cdot e_{d}\,\mathrm{d}y=0.$
Proof.
We present the arguments for $\xi_{0}$; the proof is the same for the other functions. A direct computation, including an integration by parts, yields
$0=\int_{Z}\operatorname{div}_{y}{(\xi_{0})}y_{d}\,\mathrm{d}y=\int_{Z}\xi_{0}\cdot e_{d}\,\mathrm{d}y+\int_{\partial Z}\xi_{0}\cdot n\,y_{d}\,\mathrm{d}\sigma_{y}.$
By the choice of boundary conditions and periodicity, the above integral over $\partial Z$ vanishes and we obtain the desired result. ∎
Given the solutions of the cell problems (4.2) and (4.2), we obtain for a.a. $(t,\tilde{x},y)\in S\times\Sigma\times Z$ the representations
$u(t,\tilde{x},y)=\xi_{0}(y)-\sum_{i=1}^{d-1}\frac{\xi_{i}(y)}{\mu(\theta^{f}(t,\tilde{x}))}\partial_{x_{i}}p(t,\tilde{x}),\quad q_{1}(t,\tilde{x},y)=\eta_{0}(y)-\sum_{i=1}^{d-1}\frac{\eta_{i}(y)}{\mu(\theta^{f}(t,\tilde{x}))}\partial_{x_{i}}p(t,\tilde{x}).$ (4.11)
This can be checked easily by plugging them into the two-pressure Stokes system. Looking at $\bar{u}=\int_{Z}u\,\mathrm{d}y$, we can also check that with $\operatorname{div}_{\tilde{x}}\bar{u}=0$ and by testing with $\psi\in H^{1}(\Sigma)$ while using Eq. 4.11:
$\begin{split}0&=\left(\operatorname{div}_{\tilde{x}}{\bar{u}},\psi\right)_{\Sigma}=-\left(\bar{u},\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\left(\bar{u}\cdot n_{\Sigma},\psi\right)_{\partial\Sigma}\\ &=-\left(\int_{Z}\xi_{0}\,\mathrm{d}y,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\sum_{i=1}^{d-1}\left(\frac{\partial_{i}p}{\mu(\theta^{f})}\int_{Z}\xi_{i}\,\mathrm{d}y,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\left(\left(\int_{0}^{1}u_{BC}(\cdot,s)\,\mathrm{d}s\right)\cdot n_{\Sigma},\psi\right)_{\partial\Sigma}\\ &=-(\bar{\xi}_{0},\nabla_{\tilde{x}}\psi)_{\Sigma}+\left(\frac{1}{\mu(\theta^{f})}K\nabla_{\tilde{x}}p,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+(\bar{u}_{\text{BC}}\cdot n_{\Sigma},\psi)_{\partial\Sigma}.\end{split}$ (4.12)
Where for the last step we defined
$\bar{\xi}_{0}\coloneqq\int_{Z}\xi_{0}(y)\,\mathrm{d}y,\quad K_{ij}\coloneqq\int_{Z}\xi_{j}(y)\cdot e_{i}\,\mathrm{d}y\quad\text{and}\quad\bar{u}_{\text{BC}}\coloneqq\int_{0}^{1}u_{BC}(\cdot,s)\,\mathrm{d}s.$ (4.13)
Note, Eq. 4.12 is similar to a Darcy equation, but with an additional term induced by the Dirichlet condition at $y_{d}=0$. The following Theorem 3 collects all equations and shows the complete model.
Theorem 3 (Homogenized model).
Let the Assumptions (A1)-(A7) be fulfilled and let $(\theta^{s},\theta^{f},u,p)$ be the limit functions defined in Proposition 1. Set $\bar{u}(t,\tilde{x})=\int_{Z}u(t,\tilde{x},y)\,\mathrm{d}y$. Then $(\theta^{s},\theta^{f},\bar{u},p)$ are weak solutions of the following problem: inside $\Omega$ it holds
[…]
where $\bar{f}(t,\tilde{x})=\int_{\Gamma}f(t,\tilde{x},y)\,\mathrm{d}\sigma_{y}$. The fluid temperature $\theta^{f}$ fulfills
[…]
with the effective conductivity $\tilde{\kappa}\in\mathbb{R}^{d\times d}$ given by the cell solutions $\{\omega_{i}\}_{i=1}^{d-1}$ of Eq. 4.4, $\omega_{d}=0$ and
$\tilde{\kappa}_{ij}=\int_{Z}\kappa^{f}(\delta_{ij}+\nabla_{y}\omega_{i}\cdot e_{j})\,\mathrm{d}y.$
The fluid movement in the layer is governed by
[…]
here the permeability $K\in\mathbb{R}^{d\times d}$ is given by the solutions $\{\xi_{i}\}_{i=1}^{d-1}$ of the cell problems Section 4.2 via
$K_{ij}=\int_{Z}\xi_{j}\cdot e_{i}\,\mathrm{d}y$
where we set $\xi_{d}=0$ and $\bar{\xi}_{0}=\int_{Z}\xi_{0}(y)\,\mathrm{d}y$, with $\xi_{0}$ given by Section 4.2.
Proof.
That $(\theta^{s},\theta^{f},\bar{u},p)$ fulfill Eq. 4.14 can be derived with the concept of two-scale convergence, as explained and partially carried out above. ∎
Remark 4.
Note, that by Proposition 2 and the equality Eq. 4.14i we obtain that $\bar{u}\cdot e_{d}=0$. This is reasonable since we remain with an averaged interface velocity in the effective model and a nonzero flow in direction $d$ would imply a flow into (or out of) the solid domain. We recovered, ignoring the temperature-dependent viscosity, the same effective equation as in [12], where the needed assumptions for convergence are reduced, thanks to the two-scale concept.
Additionally, in 2D the effective problem for the fluid simplifies further. By Eq. 4.14j it holds $\operatorname{div}_{\tilde{x}}(\bar{u})=\partial_{x_{1}}\bar{u}_{1}=0$ and the velocity is a constant solely given by the boundary function $u_{BC}$ in Eq. 4.14l. This linearizes the problem since the temperature equation is independent of the pressure $p$.
The existence of a solution of model (4.14) is ensured by the existence of the original problem and the two-scale limits. Similar to the micro model, for a unique solution, we need additional regularity assumptions.
Theorem 4 (Uniqueness of the homogenized model).
If there is a solution of the homogenized problem where $\nabla_{\tilde{x}}\theta\in L^{\infty}(S;L^{2+\delta}(\Sigma))$ and $\nabla_{\tilde{x}}p\in L^{\infty}(S;L^{2+\delta}(\Sigma))$ for some $\delta>0$, then this solution is the only.
Proof.
Assume that $(\theta^{f}_{i},\theta^{s}_{i},\bar{u}_{i},p_{i})$, $i=1,2$, are two sets of solutions of the homogenization limit given by System (4.14), i.e., they solve the variational problems given via Eqs. 4.6 and 4.12. In the following, we will denote their differences by $(\bar{\theta}^{f},\bar{\theta}^{s},\bar{\bar{u}},\bar{p})$. W.l.o.g. we assume that $(\theta_{2}^{f},p_{2})$ is the solution enjoying the higher gradient integrability.
From Eq. 4.11 we obtain for $\bar{\bar{u}}$, that
$\|\bar{\bar{u}}\|_{L^{2}(S\times\Sigma)}^{2}\leq C\left(\nabla_{\tilde{x}}\bar{p},\bar{\bar{u}}\right)_{S\times\Sigma}+\left(\left[\frac{1}{\mu(\theta^{f}_{1})}-\frac{1}{\mu(\theta^{f}_{2})}\right]\nabla_{\tilde{x}}p_{2},\bar{\bar{u}}\right)_{S\times\Sigma}$
which, using the Lipschitz continuity of $\mu$ as well as the higher regularity of $p$, yields
$\|\bar{\bar{u}}\|_{L^{2}(S\times\Sigma)}\leq C\left(\|\nabla_{\tilde{x}}\bar{p}\|_{L^{2}(S\times\Sigma)}+\|\nabla_{\tilde{x}}p_{2}\|_{L^{\infty}(S;L^{2+\delta}(\Omega))}\|\bar{\theta}^{f}\|_{L^{2}(S;L^{2+\frac{2}{\delta}}(\Sigma))}\right).$
Testing Eq. 4.12 with $\psi=\bar{p}$, we get:
$c\|\nabla_{\tilde{x}}\bar{p}\|_{L^{2}(S\times\Sigma)}\leq C\|\nabla_{\tilde{x}}p_{2}\|_{L^{\infty}(S;L^{2+\delta}(\Omega))}\|\bar{\theta}^{f}\|_{L^{2}(S;L^{2+\frac{2}{\delta}}(\Sigma))}$
As a result, the uniqueness of velocity and pressure follows from the uniqueness of the fluid temperatures via
$\|\bar{\bar{u}}\|_{L^{2}(S\times\Sigma)}+\|\nabla_{\tilde{x}}\bar{p}\|_{L^{2}(S\times\Sigma)}\leq C\|\bar{\theta}^{f}\|_{L^{2}(S;L^{2+\frac{2}{\delta}}(\Sigma))}.$ (4.15)
We now look at the weak forms for the heat system (Eq. 4.6) and take their difference:
$(\partial_{t}\bar{\theta}^{s},\varphi)_{S\times\Omega}+|Z|(\partial_{t}\bar{\theta}^{f},\psi)_{S\times\Sigma}+\kappa^{s}(\nabla\bar{\theta}^{s},\nabla\varphi)_{S\times\Omega}+(\tilde{\kappa}\nabla_{\tilde{x}}\bar{\theta}^{f},\nabla\psi)_{S\times\Sigma}\\ +(\bar{u}_{1}\cdot\nabla_{\tilde{x}}\bar{\theta}^{f},\psi)_{S\times\Sigma}+\alpha|\Gamma|(\bar{\theta}^{s}-\bar{\theta}^{f},\phi-\psi)_{S\times\Sigma}=-(\bar{\bar{u}}\cdot\nabla_{\tilde{x}}\theta^{f}_{2},\psi)_{S\times\Sigma}$
for all $(\varphi,\psi)\in L^{2}(S;H^{1}(\Omega)\times H^{1}(\Sigma))$. For the advection term on the left-hand side, we test with $(\bar{\theta}^{s},\bar{\theta}^{f})\in L^{2}(S;H^{1}(\Omega)\times H^{1}(\Sigma))$ and exploit that $\bar{u}_{1}$ is divergence-free (cf. Eq. 3.5):
$(\bar{u}_{1}\cdot\nabla_{\tilde{x}}\bar{\theta}^{f},\bar{\theta}^{f})_{S\times\Sigma}=\frac{1}{2}(u_{BC}\cdot n,|\bar{\theta}^{f}|^{2})_{S\times\Sigma}\geq 0.$
The advective term on the right-hand side can be handled via the additional regularity of $\nabla_{\tilde{x}}\theta^{f}_{2}$. With the same line of argument as in Theorem 2 and using the Sobolev embedding $H^{1}\subset L^{2+\frac{2}{\delta}}$ for 2D domains, uniqueness now follows. ∎
5 Simulations
In addition to the previous analysis, we investigate the properties of the effective model with the help of numerical simulations and examine what aspects can be captured in the homogenized problem. In particular, we are looking at the effects of the roughness patterns, viscosity function and inflow conditions. We also study the accuracy of the homogenization limits by direct comparison with the $\varepsilon$-model for small values of $\varepsilon$. Note that we focus on the two-dimensional case for this comparison as the computational cost for the $\varepsilon$-problem is prohibitively high in the case $d=3$.
All simulations are carried out with the finite element solver FEniCS [7] and meshes are created with Gmsh [31]. To stabilize the convection term in our model, a SUPG scheme [15] is applied. We use continuous, piecewise linear finite elements to represent temperature fields $(\theta^{f},\theta^{s})$ and pressure $p$, while the velocity $u$ is modeled with continuous piecewise quadratic elements. The time stepping is performed with a semi-implicit Euler scheme, where we linearized the convection term by using the velocity field from the previous time step. Our implementation is freely available online [24].
We start by collecting all the parameters that are fixed in the following simulations: We consider the time interval $S=(0,5)$ and the macroscopicdomain is the unit square $\Omega=(0,1)^{2}$. The conductivities are set to $\kappa^{s}=0.5$ and $\kappa^{f}=0.1$, and the heat exchange parameter to $\alpha=1.0$. For the initial conditions, everything is set to zero. At $\Sigma$, the Dirichlet condition for the velocity is $u_{\text{motion}}=(1,0)^{T}$. The heat source is given by
$f_{\varepsilon}(x)=\frac{1-\nicefrac{{x_{d}}}{{\varepsilon}}}{1-\gamma_{0}},$
so that there is a higher production at the tip of the roughness, which is motivated by the underlying application of grinding. The temperature-dependent viscosity is modeled by a Vogel-Fulcher-Tammann equation:
$\mu(\theta)=\mu_{0}\exp{\left(\frac{\alpha}{\theta-T_{0}}\right)},$
which is a reasonable approach for fluids [30]. To obtain a viscosity profile that varies in our temperature range, we choose $\mu_{0}=0.2$, $\alpha=3.0$, and $T_{0}=0.6$. Unless stated otherwise, the time step $\Delta t=0.05$ and the mesh resolution $h=0.03$ are used, where the mesh is locally refined around the thin layer with resolution $h\varepsilon$. A convergence study was carried out with respect to $\Delta t$ and $h$ and the expected linear or quadratic convergence was observed. To study the influence of different roughness functions, we investigated two different cases:
  • A sine function
    $\gamma_{\text{sine}}(x_{1};\gamma_{0})=\begin{cases}1\quad&\text{, if }x<0.1\text{ or }x>0.9,\\ 1-\frac{1-\gamma_{0}}{2}\left(\sin{\left(2\pi\frac{(x_{1}-0.1)}{0.8}-\frac{\pi}{2}\right)}+1\right)&\text{, if }x\in[0.1,0.9].\end{cases}$
  • A pattern given by smoothed rectangles
    $\gamma_{\text{rect}}(x_{1};\gamma_{0})=\begin{cases}1\quad&\text{, if }x<0.1\text{ or }x>0.9,\\ 1-(1-\gamma_{0})\left(-2\left(\frac{x_{1}-0.1}{0.1}\right)^{3}+3\left(\frac{x_{1}-0.1}{0.1}\right)^{2}\right)\quad&\text{, if }x\in[0.1,0.2),\\ \gamma_{0}\quad&\text{, if }x\in[0.2,0.8],\\ \gamma_{0}+(1-\gamma_{0})\left(-2\left(\frac{x_{1}-0.8}{0.1}\right)^{3}+3\left(\frac{x_{1}-0.8}{0.1}\right)^{2}\right)\quad&\text{, if }x\in(0.8,0.9].\end{cases}$
The reference cell $Z$ as well as the effective parameters for both roughness patterns, and three different $\gamma_{0}$, are stated in Table 1.
$Z$ $\gamma_{0}$ 0.1 0.5 0.9
[Uncaptioned image] $\begin{array}[]{c}\tilde{\kappa}=0.238\,\kappa^{f}\\ K=0.00048\\ \bar{\xi}_{0}=0.0692\end{array}$ $\begin{array}[]{c}\tilde{\kappa}=0.630\,\kappa^{f}\\ K=0.01704\\ \bar{\xi}_{0}=0.2871\end{array}$ $\begin{array}[]{c}\tilde{\kappa}=0.951\,\kappa^{f}\\ K=0.07025\\ \bar{\xi}_{0}=0.4716\end{array}$
[Uncaptioned image] $\begin{array}[]{c}\tilde{\kappa}=0.120\,\kappa^{f}\\ K=0.00011\\ \bar{\xi}_{0}=0.0526\end{array}$ $\begin{array}[]{c}\tilde{\kappa}=0.523\,\kappa^{f}\\ K=0.01124\\ \bar{\xi}_{0}=0.2556\end{array}$ $\begin{array}[]{c}\tilde{\kappa}=0.917\,\kappa^{f}\\ K=0.06292\\ \bar{\xi}_{0}=0.4552\end{array}$
Table 1: Roughness patterns used in the numerical simulations, and the corresponding effective parameters. Since the setup is two-dimensional, all parameters are reduced to scalar values.
Three different inflow patterns are compared, a linear inflow, a quadratic, and a linear inflow that is not zero above $\gamma_{0}$ given by
$(i)\ \ u_{\text{BC, lin}}=\begin{cases}\left(1-\frac{x_{2}}{\gamma_{0}}\right)u_{\text{motion}}&\text{, if }x_{2}\leq\gamma_{0},\\ 0&\text{, if }x_{2}>\gamma_{0},\end{cases}\quad$ $,\quad(ii)\ \ u_{\text{BC, quad}}=\begin{cases}\left(1-\left(\frac{x_{2}}{\gamma_{0}}\right)^{2}\right)u_{\text{motion}}&\text{, if }x_{2}\leq\gamma_{0},\\ 0&\text{, if }x_{2}>\gamma_{0},\end{cases}$
and ($\text{and }(iii)\ \ u_{\text{BC, lin, 2}}$ $=(1-x_{2})u_{\text{motion}}.$
The effective inflow then equals $\bar{u}_{\text{BC, lin}}=\frac{1}{2}\gamma_{0}$, $\bar{u}_{\text{BC, quad}}=\frac{2}{3}\gamma_{0}$ or $\bar{u}_{\text{BC, lin, 2}}=\frac{1}{2}$. With $u_{\text{BC, lin, 2}}$ we investigate how the model behaves if Assumption Item (A5) is not fulfilled. In this case, we can not build an extension of the boundary condition to the whole domain by a simple scaling argument. Instead, the extension has to be carried out in the thin rough layer directly. If not specified otherwise, we use $u_{\text{BC, lin}}$ in the following simulation studies.
Remark 5.
Please note that the overall simulation scenarios are directly motivated by the grinding application in the way we chose the boundary conditions (inflow/outflow), the roughness patterns, and the viscosity functions. However, these simulation experiments are only qualitative, as we are not working with realistic and sensible parameter values for the conductivities, the heat source, etc. We refer to [57] where quantitative simulations using realistic material parameters were performed for a single grain scenario.
Comparison between $\varepsilon$-model and effective model. We start by comparing the solutions to the effective model with solutions of the resolved $\varepsilon$-model (where we generally compare with values $\varepsilon=0.1,0.05.,0.01$) for the sinusoidal roughness pattern and $\gamma_{0}=0.5$. To give a general overview of the solution, we present them in Fig. 2 and Fig. 3(d). In the following, we focus mainly on comparing the solutions inside the thin layer as this is the critical area where the interesting features (rough boundary and thin domain) are present. As the solutions are defined on different domains (the effective solutions are defined on a line and the $\varepsilon$-solutions on a thin domain), we take an average of the microscopic solution with respect to the vertical direction and compare this with the corresponding effective quantities, as done in similar numerical studies [40, Section 5]. To this end, we define the averaged functions
$\begin{split}\bar{\theta}^{f}_{\varepsilon}(t,\tilde{x})\coloneqq\frac{1}{\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}&\int_{0}^{\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}\theta^{f}_{\varepsilon}(t,\tilde{x},s)\,\mathrm{d}s,\quad\bar{p}_{\varepsilon}(t,\tilde{x})\coloneqq\frac{\varepsilon}{\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}\int_{0}^{\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}p_{\varepsilon}(t,\tilde{x},s)\,\mathrm{d}s\\ &\bar{u}_{\varepsilon}(t,\tilde{x})\coloneqq\frac{1}{\varepsilon}\int_{0}^{\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}u_{\varepsilon}(t,\tilde{x},s)\,\mathrm{d}s.\end{split}$ (5.1)
Note that $\bar{p}_{\varepsilon}$ is scaled with $\varepsilon^{2}$ ($\nicefrac{{\varepsilon^{2}}}{{\varepsilon\gamma}}=\nicefrac{{\varepsilon}}{{\gamma}}$) since the effective pressure $p$ is the limit of $\varepsilon^{2}p_{\varepsilon}$. We observe that the effective model captures the resolved $\varepsilon$-model fairly well with regard to both temperature and pressure, as observable in Fig. 2 to 4. We obtain negative pressure at the inflow ($x_{1}=0$) since the amount of fluid pushed into the domain by $u_{BC,lin}$ is less than the fluid movement induced by the bottom boundary. This becomes also apparent, by comparing $\bar{u}_{\text{BC, lin}}$ with the flow induced by the cell problem (4.2), which gives the value $\bar{\xi}_{0}=0.29$ (see Table 1).
Note that for velocity, the average of the microscopic solution is already close to the effective value of $\bar{u}_{\text{BC, lin}}=\frac{1}{2}\gamma_{0}=0.25$ and only oscillates slightly around this value, see Fig. 4. These oscillations originate mainly from numerical inaccuracies, induced both by the FE-scheme and the numeric integration to compute Eq. 5.1.
(a) $\varepsilon=0.1$
(b) $\varepsilon=0.05$
(c) $\varepsilon=0.01$
(d) Effective
Figure 2: Comparison of the temperature profiles, at $t=5$, for the micro and effective model, with sinusoidal roughness. The effective fluid temperature $\theta^{f}$ is only defined on the line $(0,1)\times\{0\}$ but for visualization purposes the line is being shown with a thickness
(a) $\varepsilon=0.1$
(b) $\varepsilon=0.05$
(c) $\varepsilon=0.01$
(d) Effective
Figure 3: Comparison of the fluid (top) and pressure (bottom) profiles, at $t=5$, for the micro and effective model, with sinusoidal roughness. The line in the effective case is again visualized with a thickness
Figure 4: Comparison inside the thin layer of the effective solutions with the averaged micro solutions defined in Eq. 5.1. All solutions are shown at the final time step $t=5$ and plotted over $\Sigma$. The velocity field for $\varepsilon=0.01$ is left out for a clearer visualization.
We also want to study numerically the convergence behavior for $\varepsilon\to 0$. To this end, we compute inside the thin layer the difference of the effective and averaged solution in the $L^{2}(S\times\Sigma)$ norm, and in the solid bulk domain, we compare in the $L^{2}(S\times\Omega^{s}_{\varepsilon})$ norm. The results are shown in Fig. 5. For temperature and pressure, we observe a convergence of order $\mathcal{O}(\varepsilon)$. On the other hand, there is no observable convergence for the velocity, as mentioned before, we assume that this is mainly due to numerical errors. The difference between effective and averaged velocity is also already small compared to temperature and pressure error.
Figure 5: Convergence study for the limit $\varepsilon\to 0$. Both axes are on a logarithmic scale. Inside the thin layer, meaning for $\theta^{f},p,u$, we present the error in the $L^{2}(S\times\Sigma)$ norm. The error for $\theta^{s}$ is computed in $L^{2}(S\times\Omega^{s}_{\varepsilon})$. To estimate the order of convergence, the black line equals the identity.
Varying the inflow function. To study how well the effective model captures the influence of the prescribed inflow, we compare different Dirichlet conditions for the roughness given by $\gamma_{\text{sine}}$ for $\gamma_{0}=0.5$ and fixed $\varepsilon=0.05$. The resulting temperature, pressure, and velocity profiles at time $t=5$ are visualized in Fig. 6. We first observe that, naturally, a higher average fluid flow leads to a cooler system since the heat energy is removed faster from the system. Furthermore, inflows $u_{\text{BC, quad}}$ and $u_{\text{BC, lin, 2}}$ result in a positive pressure at the inflow interface, since the amount of fluid pressed into the domain is larger than the amount moved by the bottom boundary condition. The homogenized model can capture the influence of the different inflow conditions fairly well with regard to temperature and pressure. This holds true even for $u_{\text{BC, lin, 2}}$, which does not satisfy the assumptions made in Section 2.3.
Figure 6: Comparison of the fluid temperature, pressure, and horizontal velocity component for different inflow functions, at $t=5$. In black the micro solution, and in orange the effective results. The velocity is visualized along $\{0.5\}\times(0,\varepsilon)$. For the other functions, we again compare the average (5.1).
Varying the grain height $\gamma_{0}$. We investigate the influence of the minimal height of the interface $\gamma_{0}$. Here, we again focus on the roughness given by $\gamma_{\text{sine}}$ and fixed $\varepsilon=0.05$. We use the inflow function $u_{\text{BC, lin, 2}}$ as it is independent of $\gamma_{0}$; this allows us to focus on the influence of the interface height. Similarly, we modify the heat source to $f_{\varepsilon}=\tfrac{1}{|\Gamma|}$ ensuring that the total amount of heat produced is one and is independent of $\gamma_{0}$. The simulated solutions, again at $t=5$, are presented in Fig. 7. As in the results before, the homogenized model shows good agreement with the exact solutions for $\varepsilon=0.05$. The grain height has a strong influence on the pressure profile; since we have a fixed amount of inflow for all cases, a smaller gap necessarily leads to higher pressures inside the layer. At the same time, the fluid velocity in between the ”humps” created by the roughness is also reduced, leading to slightly higher temperatures.
Figure 7: Comparison of the fluid temperature, pressure, and horizontal velocity component for different roughness heights $\gamma_{0}$, at $t=5$. The solution is visualized as in Fig. 6.
Varying the roughness pattern. In this simulation scenario for the two-dimensional case, we demonstrate the influence of different roughness patterns. We utilize the same setup as in the previous case for varying roughness height but this time we fix $\gamma_{0}=0.5$ and compare the differences between $\gamma_{\text{sine}}$ and $\gamma_{\text{rect}}$. As before, we present the results at the last time step, in Fig. 8. The homogenized model accurately reflects the influence of the roughness pattern on all solution fields. For 2D we can conclude, that the derived effective model yields a good approximation of the resolved micro model and the temperature-dependent viscosity has a, rather large, impact on the pressure profile.
Figure 8: Comparison of the fluid temperature, pressure, and horizontal velocity component for different roughness types, at $t=5$. The solution is visualized as in Fig. 6.
3D simulation study. Lastly, we demonstrate the effects of the temperature-dependent viscosity on the fluid velocity $u$. For this, a three-dimensional setup is required. Unfortunately, the $\varepsilon$-problems become computationally very expensive in this case as they require a lot of memory. Therefore, we only simulate the $\varepsilon$-model for $\varepsilon=0.1$ (as opposed to the 2D case where we went down to $\varepsilon=0.01$). To increase the effect of temperature-dependent viscosity on the flow field, we set $\kappa^{f}=0.01$ and $C=0.4$. Additionally, we consider a local heat source given by
$f_{\varepsilon}(x)=\begin{cases}1\quad&\text{, if }x_{2}>0.5\text{ and }|x_{1}-0.5|<0.1,\\ 0&\text{ else}.\end{cases}$
This right-hand side is chosen in such a way that local variations in the heat production occur, leading to local differences in the viscosity. The inflow is defined by $u_{\text{BC, lin}}$ with $\gamma=0.5$ and $u_{\text{motion}}=(1,0,0)^{T}$. For the roughness we define
$\gamma=1-0.9\left(\cos{(2\pi(x_{2}-0.5)^{2})}\cos(2\pi(x_{1}-0.25))\right)^{2},$
so that the effective quantities behave differently in directions $e_{1}$ and $e_{2}$. The corresponding effective parameters are collected in Table 2.
$Z$ $\tilde{\kappa}$ $K$ $\bar{\xi}_{0}$
[Uncaptioned image] $\left(\begin{array}[]{cc}0.493&0.0\\ 0.0&0.635\end{array}\right)$ $\left(\begin{array}[]{cc}0.004566&0.0\\ 0.0&0.004947\end{array}\right)$ $\left(\begin{array}[]{c}0.134\\ 0.0\end{array}\right)$
Table 2: Effective parameters for the three-dimensional simulation. The effective values corresponding to dimension $d$ are not listed, since they are zero.
The nonlinearity is resolved by splitting up the problem and doing the time stepping for temperature and Stokes problem sequentially. To be precise, we do a time step for the fluid velocity and pressure using the temperature from the previous time step, followed by a time step of the temperature problem with the velocity of the new time step.
The simulation results at three distinct time steps are presented in Fig. 10. The effect of the temperature-dependent viscosity can be observed in the flow profile. At the start ($t=0.05$), the temperature is more or less constant, so there is a uniform velocity field. The viscosity decreases locally ($t=1.85$) due to the localized heat source, and the fluid flows in a diagonal direction towards the area of lower viscosity. As soon as the heat starts to spread through the advection, the observed effect decreases $(t=5.0)$. We conclude, that the temperature-dependent viscosity can lead to effects that a constant viscosity could not depict, both in the microscale and effective model. This effect appears in both models but is easier to recognize, in Fig. 10, for the homogenized case, since the mesh does not need to resolve the roughness and has a uniform structure. Again, the effective model yields a good approximation of the resolved $\varepsilon$-model while requiring much less computational effort.
(a) $t=0.05$
(b) $t=1.85$
(c) $t=5.0$
Figure 10: Temperature and velocity profile for the three-dimensional setup, at different snapshots in time. At the top, is the resolved micro model, and at the bottom, is the effective model. For the micro model, a cross-section at $x_{3}=0.01$ is visualized, such that the flow profile is better recognizable. Note, that the mesh for the micro model needs to resolve the roughness structure and therefore the representation of the velocity profile is rather disorganized, while in the effective case, we can use a uniform mesh.
6 Conclusion
We investigated the effective behavior of a nonlinear Stokes-temperature system in a thin layer, where the nonlinearity arises through the convection term in the heat equation and temperature-dependent viscosity. With the concept of two-scale convergence, an effective model was derived where the effective fluid temperature is given by an interface function. We point out that, for the fluid velocity and pressure, we recovered the results of [12] with an additional coupling to the temperature field through viscosity. Multiple simulation results validate the derived model and show that temperature-dependent viscosity also plays a non-negligible role in the limit $\varepsilon\to 0$.
In reality, other parameters (such as conductivity or heat exchange parameter) may also be temperature dependent. Similarly to the present approach, for example, by decoupling the system and showing the existence of solutions with a fixed-point argument, temperature-dependent $\kappa$ and $\alpha$ could be handled as well. Of course, with fitting assumptions for the parameters. Another aspect that could be investigated further is the current $\varepsilon$-scaling of $\kappa$ in the thin layer. Removing this scaling complicates the analysis and may lead to an effective equation without diffusion on the interface $\Sigma$.
Regarding the motivating application of a grinding process, we could derive a justification for the models used in the literature; see [33]. Currently, we are looking at the idealized scenario of periodic roughness. In future research, we plan to extend this model to real grinding wheels and workpiece surfaces. There, we also plan to investigate the effect of including microstructures in the whole process.
Acknowledgements
TF acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project nr. 281474342/GRK2224/2. Additionally, his research was also partially funded by the DFG – project nr. 439916647.
The research activity of ME is funded by the Horizon 2022 research and innovation program of the European Union under the Marie Skłodowska-Curie fellowship project MATT (project nr. 101061956, https://doi.org/10.3030/101061956).
Conflict of interest
The authors declare that there is no conflict of interest.
Data Availability Statement
The research code associated with this article is available in the public repository [24].

### v3

@@
-Analysis and Simulation of a Fluid-Heat System in a Thin, Rough Layer in Contact With a Solid Bulk Domain
+
 Abstract
@@
 Moreover, the opposing vertical surfaces of the thin layer are in relative motion.
-This setup is particularly relevant to grinding processes, where cooling lubricants interact with the rough surface of a rotating grinding wheel.
+This setup is motivated by grinding processes, where cooling lubricants interact with the rough surface of a rotating grinding wheel.
 The resulting model is nonlinearly coupled through $(i)$ temperature-dependent viscosity and $(ii)$ convective heat transport.
@@
 The corresponding numerical code is freely available on GitHub.
-Keywords: Homogenization, mathematical modeling, dimension reduction, two-scale convergence, thin fluid films, numerical simulations
-MSC 2020: 35B27, 76-10, 76M50, 65N30, 80M40, 80A19 {NoHyper} 1Corresponding author, Email: tomfre@uni-bremen.de
+keywords:
+Homogenization, mathematical modeling, dimension reduction, two-scale convergence, thin fluid films, numerical simulations
 1 Introduction
@@
 Here, the workpiece is the material that is processed or shaped through grinding and the cooling lubricant is a fluid that is used to reduce friction and remove heat during the grinding process [1].
-The grinding wheel itself is usually a composite material consisting of abrasive grains (e.g., diamonds) and a cementing matrix called bond.
 Given the complex granular surface of the grinding wheel, thin gaps and channels emerge in the contact area of the wheel and the workpiece, allowing the coolant to flow through.
@@
 These models are heuristically derived and the microstructures and fluid flow in the gap are not accounted for.
-Additionally, fluid pressure estimates are sometimes also investigated and used as a tool to control the supply of the coolant fluid [45, 58].
 The fluid dynamics in the gap is strongly influenced by the viscosity of the coolant, which in reality is temperature dependent [5, 6]; and given the large temperature variations during the grinding process, this effect should not be ignored as it can influence the quality of the results [7, 8].
 Numerical simulations to predict the fluid pressure in the contact zone are usually done with the help of the Reynolds equation [9, 10], but often without taking into account the temperature variations.
-Closely related to grinding are lubrication problems.
-Here, the goal is not to remove material from an object, but instead to reduce the friction between two moving objects with the help of oils.
+Similar setups appear in lubrication problems and bearings.
+There, the aim is not to remove material, but instead to reduce the friction between two moving objects with the help of oils.
 Often, temperature variations are encountered (either by friction-induced heat or external supply) that influence the viscosity of the fluid [11, 12].
@@
 We also point to [14] for a numerical study on temperature-dependent viscosity on flows through fractures.
-Although there are differences, one typically expects much lower velocities than in our case, the analytical results presented for the temperature could be transferred.
-Starting with an idealized periodic setup with a small parameter $\varepsilon\ll 1$ characterizing both the length scale of the periodicity as well as the average thickness of the grinding gap, we rigorously derive an effective, homogenized model accounting for the nonlinear coupling between heat and fluid flow dynamics in such a setup.
+Although there are differences, one typically expects much lower velocities than in our case, the analytical results presented in this article for the temperature could still be transferred.
+We point out that while our model is motivated by a grinding process, we do not try to capture all or even most of the relevant aspects of such a process; for example, we do not include solid-to-solid contact, time-dependent contact zones, or mechanical deformations.
+Instead, we consider a simplified framework to study some phenomena in a mathematically rigorous setting.
+We acknowledge that our assumptions are strong idealizations and may not directly reflect a real-world application.
+Many of the not included aspects are more challenging to model mathematically, and remain open from a rigorous mathematical standpoint.
+However, we believe that the mathematical results in this article still offer a starting point for further studies and are also relevant to other problems.
+Starting with an idealized periodic setup with a small parameter $\varepsilon\ll 1$ characterizing both the length scale of the periodicity as well as the average thickness of the thin layer, we rigorously derive an effective, homogenized model accounting for the nonlinear coupling between heat and fluid flow dynamics.
 This is accomplished by conducting a limiting procedure $\varepsilon\to 0$ within the framework of two-scale homogenization for thin domains.
-From a mathematical point of view, the problem consists of two main aspects: a large bulk domain where the rough interface acts similar to a rough boundary and a thin rough layer.
-For bulk domains with a rough boundary, homogenization procedures, including diffusion or fluid flow problems, can be found in [10, 16, 44], or in the case of rough interfaces we point to [19].
+From a mathematical point of view, the problem consists of two main aspects: a large bulk domain where the rough interface acts similarly to a rough boundary, and a thin rough layer.
+For bulk domains with a rough boundary, homogenization procedures, including diffusion or fluid flow problems, can be found in [15, 16, 17], or in the case of rough interfaces, we point to [18].
 The case where the amplitude of the roughness does not scale with $\varepsilon$ has also been considered in the literature, see for example [16, 18, 19] and the references therein.
@@
 Both formal as well as rigorous limit procedures have been carried out for such problems and the thin layer reduces to effective interface equations, see [20, 13, 21].
-On the other hand, homogenization of diffusion problems in thin layers with periodic structures has been extensively studied before with the concept of two-scale convergence; see for example [29, 46].
+On the other hand, homogenization of diffusion problems in thin layers with periodic structures has been extensively studied before with the concept of two-scale convergence; see [22, 23].
 Effective fluid flow through thin layers was rigorously investigated in [24] and later extended for a periodic roughness in [25].
-Starting with the stationary Stokes equation, they proved the validity of the Reynolds equation [51] when the layer height is small when compared to the length of the roughness period.
+Starting with the stationary Stokes equation, they proved the validity of the Reynolds equation [26] when the layer height is small compared to the length of the roughness period.
 If the periodicity and the height are of the same order (usually called the Stokes roughness regime), a model similar to a Darcy equation is derived.
 The Reynolds equation plays an important role in lubrication, which has a setup similar to our underlying scenario.
-The equation has been further studied and homogenized in the literature [6, 13, 17, 23]; see also [42, Section 1] for a good overview.
+The equation has been further studied and homogenized in the literature [27, 28, 29, 30]; see also [31, Section 1] for a good overview.
 For layers without roughness, the first derivation of the Reynolds equation with the concept of two-scale convergence was achieved in [32].
 Flow through a thin layer, without oscillations, coupled with a temperature-dependent viscosity was studied in [33], where a Reynolds equation with temperature-dependent viscosity was derived via asymptotic expansion.
-To be more specific, we consider a complex geometry incorporating a rough, thin layer similar to the setup in [12], where the boundaries of the layer are additionally in relative motion to each other (see Fig.
-1).
+Furthermore, fluid flow in thin (porous) layers has been rigorously homogenized with the concept of two-scale convergence [34, 35] and flow through a porous interface [36].
+To be more specific, we consider a geometry incorporating a rough, thin layer, where the boundaries of the layer are additionally in relative motion to each other.
 In this geometry, we are considering the coupling between a quasi-stationary Stokes system for the fluid flow and the heat dynamics.
@@
 We assume that the viscosity is a Lipschitz continuous function with respect to the temperature.
-We show that the resulting nonlinear problem has a weak solution (Theorem 1) via a Schauder fixed point argument where uniqueness follows under additional regularity assumptions at the solution(Theorem 2).
+We show that the resulting nonlinear problem has a weak solution (Theorem 1) via a Schauder fixed point argument.
+A priori estimates of the temperature are obtained with energy estimates.
+The pressure is estimated with the help of Bogovskii operators for thin layers, which can be found in existing literature [34, 35].
+Uniqueness follows under additional regularity assumptions at the solution (Theorem 2).
 In particular, any strong solution is necessarily unique.
-We then perform a limiting procedure $\varepsilon\to 0$ in the framework of two-scale convergence for thin domains and, as our main result, arrive at a homogenized model which captures the effective interaction between heat dynamics and fluid flow (Theorem 3).
-A novel aspect of our work is the derivation of $\varepsilon$-independent $L^{\infty}$ bounds of the temperature field, utilizing results of [41].
-These are used to obtain strong two-scale convergence in the thin layer, which is required to pass to the limit in the nonlinear terms.
+The main novel aspect of our work is the derivation of $\varepsilon$-independent $L^{\infty}$ bounds of the temperature field in the thin layer (Lemma 6), utilizing results of [37].
+These are used to obtain strong two-scale convergence, which is required to pass to the limit in the nonlinear terms.
 To this end, we also show in Section 3.1 multiple ($\varepsilon$-dependent) trace estimates and embedding estimates for both the thin layer and the bulk domain.
-Please note that the fluid part of this effective model (e.g., when the temperature is constant) is the same as the limit presented in [12, Section 3.3].
+The limit procedure itself can then be carried out with arguments already established in prior works [38, 36, 35], with slight modifications of the arguments to obtain a two-pressure representation and the boundary conditions.
+The homogenized model is stated in Theorem 4.
 Similarly, as for the $\varepsilon$-problem, we are able to establish uniqueness for the homogenized model under additional regularity assumptions (Theorem 5).
-A small convergence study $\varepsilon\to 0$ is carried out numerically in two dimensions, investigating the order of convergence of the error between the effective and microscale model averaged over the thin layer.
+In the end, a small convergence study $\varepsilon\to 0$ is carried out numerically, investigating the order of convergence of the error between the effective model and the microscale model.
 Moreover, we conduct several numerical experiments showcasing and investigating the effects of the temperature-dependent viscosity, the roughness pattern, as well as the inflow condition.
 The code of our implementation is uploaded and made freely accessible on GitHub [39].
+The limit model (see Theorem 4) consists of a heat equation for the solid bulk domain coupled to heat and fluid dynamics on a lower-dimensional surface.
+The effective equation for the fluid temperature recovers the models derived for diffusion problems in thin connected layers [22, 36], but with an additional advection term induced by the effective flow field.
+The effective fluid velocity is given by a two-pressure Stokes system, which can be reduced to a Darcy-like equation with a temperature-dependent viscosity.
+Ignoring the viscosity, the same effective equation as in [25] is obtained; with fewer mathematical assumptions thanks to the use of two-scale convergence.
+The effective flow model is also similar to [35] where pressure-driven flows in thin porous layers were analyzed.
+The cell problems (4.2.2) used to compute the effective permeability in the layer are the same as in classical porous media [40, Eq.
+(1.11)] and also appear, with slightly modified boundary conditions, in [35, Eq.
+(6.4)].
+In addition to the permeability, a term induced by the relative motion appears in the Darcy equation, see Section 4.2.2.
+A similar cell problem occurs in [36, Eq.
+(20)] when computing the effective flow through a thin porous interface.
+We conclude that individual parts of the derived limit model have appeared in the literature in different related setups.
+The main novelty of this work is to combine these interesting features, in particular with respect to the coupling through the temperature-dependent viscosity and the advection term.
+Mathematically speaking, this introduces several difficulties.
+For one, the nonlinear coupling necessitates a careful fixed-point argument to ensure existence of solutions where uniqueness is generally only to be expected under higher regularity assumptions (see Theorem 2).
+In addition, uniform (w.r.t.
+the scale parameter $\varepsilon$) $L^{\infty}$-estimates are needed for this fixed-point argument to work.
+Due to the domain interactions (bulk domain connected to a thin fluid layer), this calls for the careful use of $\varepsilon$-controlled embeddings (see lemmas 3 and 6).
+From an application point of view, this is of particular interest whenever large temperature variations are possible as this generally implies changes in viscosity.
+Returning specifically to the application of grinding, a similar model was heuristically introduced in [4].
+There, the coolant is also modeled by a function defined on the interface between the wheel and the workpiece.
+We could therefore give some mathematical justification for the model used in the literature.
+However, the heuristic model is not completely the same; the fluid velocity is predetermined, no effective parameters depending on the microstructure appear, and the heat diffusion is set to zero along the interface.
 This paper is structured as follows: In Section 2, we start by introducing the geometrical setup, the mathematical model, and the assumptions needed to pass to the limit $\varepsilon\to 0$.
@@
 Finally, in Section 5, we investigate and compare the effective model and the resolved micro model with the help of numerical simulations.
-2 Underlying geometry, model and assumptions
-2.1 Geometry setup
+2 Problem formulation and assumptions
+In this section, we provided the setup of the underlying geometry and the concrete differential equations on the micro scale.
+2.1 Description of the geometry
 We start by describing the geometry and the assumptions on the regularity of the interface.
@@
 The considered time interval is denoted by $S=(0,T)$, for $T>0$.
-The general setup considered is presented in Fig.
+The general setup is presented in Fig.
 1.
 We consider spatial domains of dimension $d=2,3$.
-The unit vectors are denoted with $e_{i}\in\mathbb{R}^{d}$ for $i=1,\dots,d$ and the $i$-dimensional unit cube will be denoted by $Y^{i}=(0,1)^{i}$.
-The overall domain $\Omega\subset\mathbb{R}^{d}$ is assumed to be of the form $\Omega=\Sigma\times(0,H)$ where $\Sigma\subset\mathbb{R}^{d-1}$ is a bounded $d-1$ dimensional cuboid $\Sigma\subset\mathbb{R}^{d-1}$ with corner coordinates in $\mathbb{Z}^{d-1}$.
+The unit vectors are denoted with $e_{i}\in\mathbb{R}^{d}$ for $i=1,\dots,d$ and the $i$-dimensional unit cube is denoted by $Y^{i}=(0,1)^{i}$.
+The overall domain $\Omega\subset\mathbb{R}^{d}$ is assumed to be of the form $\Omega=\Sigma\times(0,H)$ where $\Sigma\subset\mathbb{R}^{d-1}$ is a bounded d1d-1 dimensional cuboid $\Sigma\subset\mathbb{R}^{d-1}$ with corner coordinates in $\mathbb{Z}^{d-1}$.
 [^1] So $\Omega$ is also a $d$-dimensional cuboid.
-By a slight abuse of notation, we will understand $\Sigma$ both as a set in $\mathbb{R}^{d-1}$ and the lower boundary part $\Sigma\times\{0\}\subset\partial\Omega$.
-Any point $x\in\Omega$ will also be denoted by $x=(\tilde{x},x_{d})\subset\Sigma\times(0,H)$, the same notation is used for $y\in Y^{d}$.
+By a slight abuse of notation, we understand $\Sigma$ both as a set in $\mathbb{R}^{d-1}$ and the lower boundary part $\Sigma\times\{0\}\subset\partial\Omega$.
+Any point $x\in\Omega$ is also denoted by $x=(\tilde{x},x_{d})\in\Sigma\times(0,H)$, the same notation is used for $y\in Y^{d}$.
 Figure 1: Schematic depiction of the geometric setup.
@@
 Right: The reference cell.
-The rough interface between solid and fluid is given through a $Y^{d-1}$-periodic $C^{1}$ function $\gamma:\overline{Y^{d-1}}\to[\gamma_{0},1]$ for some $10<\gamma_{0}<$ 1.
-This lower bound $\gamma_{0}>0$ ensures that the interface does not touch the bottom boundary; see Remark 2 for a discussion regarding this choice and case $\gamma_{0}=0$ and a connection to our motivating application.
-We introduce the reference cell $Z\subseteq Y^{d}$ and the reference interface $\Gamma$ given by
-$Z=\left\{x\in Y^{d}:x_{d}<\gamma(\tilde{x})\right\}\quad\text{ and }\quad\Gamma=\left\{x\in Y^{d}:x_{d}=\gamma(\tilde{x})\right\}.$
-Since the corners of $\Sigma$ are at integer coordinates, it can be tiled perfectly with $\varepsilon Y^{d-1}$ for $\varepsilon:=\varepsilon_{n}=n^{-1}$ for any $n\in\mathbb{N}$.
+To construct the rough thin layer, we introduce the open periodic reference cell $Z\subseteq Y^{d}$ for the fluid domain. We assume, for technical reasons, that
+  • $Z$ is Lipschitz and connected, $Y^{d}\setminus Z$ is either empty or also Lipschitz and connected,
+  • there is a $\gamma_{0}>0$ such that $y_{d}\geq\gamma_{0}$ for all $y\in Y^{d}\setminus Z$,
+  • the domain is periodic in the lateral directions, meaning
+    $\partial Z\cap\{y_{i}=0\}+e_{i}=\partial Z\cap\{y_{i}=1\},\quad\text{for }i=1,\dots,d-1,$
+  • if $Y^{d}\setminus Z\neq\emptyset$, then $Y^{d}\setminus Z$ touches the top boundary:
+    $\left|\partial(Y^{d}\setminus Z)\cap\{y_{d}=1\}\right|>0.$
+The lower bound $\gamma_{0}>0$ ensures that the solid does not touch the bottom boundary.
+The reference interface $\Gamma$ and the combined upper and lower interface of $Z$, denoted by $\Gamma_{d}$, are given by
+$\Gamma\coloneqq\operatorname{int}(\partial Z\cap\partial(Y\setminus Z)),\quad\Gamma_{d}:=\Gamma\cup\{y_{d}=0\}.$
+Since the corners of $\Sigma$ are at integer coordinates, we can tile $\Sigma$ with $\varepsilon Y^{d-1}$ for $\varepsilon\coloneqq\varepsilon_{n}=n^{-1}$ for any $n\in\mathbb{N}$.
 In the following, we suppress the $n$ and note that any limit $\varepsilon\to 0$ is understood as a limit $n\to\infty$.
-With that in mind, we introduce the rough solid domain $\Omega^{s}_{\varepsilon}$ and the rough, thin fluid layer $\Omega^{f}_{\varepsilon}$ by (assuming that $\varepsilon<H$)
-$\Omega^{s}_{\varepsilon}=\left\{x\in\Omega:x_{d}>\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right\}\quad\text{and}\quad\Omega^{f}_{\varepsilon}=\left\{x\in\Omega:x_{d}<\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right\}.$
-Here, roughness refers to the fast, oscillatory structure (due to $\nicefrac{{x}}{{\varepsilon}}$) of the interface between the fluid and grinding wheel which is given by
-$\Gamma_{\varepsilon}=\left\{x\in\Omega:x_{d}=\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right\}.$
+With that in mind, we define $\mathcal{K}_{\varepsilon}\coloneqq\{k\in\mathbb{Z}^{d-1}\times\{0\}:\varepsilon(Y+k)\subset\Omega\}$.
+We can introduce the rough, thin fluid layer $\Omega^{f}_{\varepsilon}$, the rough solid domain $\Omega^{s}_{\varepsilon}$, and the rough interface $\Gamma_{\varepsilon}$ by (assuming that $\varepsilon<H$)
+$\Omega^{f}_{\varepsilon}\coloneqq\operatorname{int}\left(\bigcup_{k\in\mathcal{K}_{\varepsilon}}\varepsilon(\overline{Z}+k)\right),\quad\Omega^{s}_{\varepsilon}\coloneqq\Omega\setminus\overline{\Omega}_{\varepsilon}^{f},\quad\text{and}\quad\Gamma_{\varepsilon}\coloneqq\operatorname{int}\left(\partial\Omega^{f}_{\varepsilon}\cap\partial\Omega^{s}_{\varepsilon}\right).$
+By construction, $\Omega^{f,s}_{\varepsilon}$ are connected, and we assume they are both also Lipschitz.
 Note that $\Omega=\Omega^{s}_{\varepsilon}\cup\Omega^{f}_{\varepsilon}\cup\Gamma_{\varepsilon}$.
@@
 where the sets $\Sigma^{\text{in}}_{\varepsilon},\Sigma^{\text{out}}_{\varepsilon}$ are given by
-$\Sigma^{\text{in}}_{\varepsilon}=\left\{(\tilde{x},x_{d}):\tilde{x}\in\Sigma^{\text{in}},x_{d}\in\left(0,\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right)\right\}\quad\text{ and }\quad\Sigma^{\text{out}}_{\varepsilon}=\left\{(\tilde{x},x_{d}):\tilde{x}\in\Sigma^{\text{out}},x_{d}\in\left(0,\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right)\right\}.$
+$\Sigma^{\text{in}}_{\varepsilon}\coloneqq(\Sigma^{\text{in}}\times(0,\varepsilon))\cap\partial\Omega^{f}_{\varepsilon}\quad\text{ and }\quad\Sigma^{\text{out}}_{\varepsilon}\coloneqq(\Sigma^{\text{out}}\times(0,\varepsilon))\cap\partial\Omega^{f}_{\varepsilon}.$
 Finally, with $\Omega_{\varepsilon}=\Sigma\times(0,\varepsilon)$, we denote the thin layer without any roughness.
@@
 Furthermore, for the length of the rough interface and the remaining boundary of the solid domain, it holds $|\Gamma_{\varepsilon}|,|\Sigma^{s}_{\varepsilon}|=\mathcal{O}(1)$.
-In our motivating example of a grinding process, the grinding wheel moves over the workpiece, with a given velocity $u_{\text{motion}}$.
-Since we consider only the section of the grinding gap, it is reasonable to assume that $u_{\text{motion}}\in\mathbb{R}^{d}$ and $u_{\text{motion}}\cdot e_{d}=0$.
-Because of the movement both domains $\Omega^{s}_{\varepsilon}$ and $\Omega^{f}_{\varepsilon}$ would be time-dependent.
-To circumvent the homogenization of a moving domain we will view the problem from the reference frame of the grinding wheel and obtain a fixed geometry, see also the model description in the next section.
-Remark 1.
-In the case of grinding, the lower boundary of the fluid layer would also be given by a rough surface describing the workpiece.
-In our setup, this roughness may be characterized by another function $\gamma_{2}$, but with the same period of $\varepsilon$.
-It is also reasonable to assume that the workpiece roughness is constant in the direction of $u_{\text{motion}}$, therefore one can again work in a fixed reference frame.
-To simplify the notation we only consider one oscillating interface, but the following analysis can be transferred to the more general case.
-The smoothness assumption of $\gamma$ stems from technical reasons and is needed to estimate the fluid pressure.
-With the subscript ##\## we will indicate a space of periodic functions in the directions $11,\dots,d-$ 1, for example
-$H^{1}_{\#}(Z)\coloneqq\left\{u\in H^{1}_{\text{loc}}(\mathbb{R}^{d}):u_{|Z}\in H^{1}(Z),u(y+e_{i})=u(y)\text{ for almost all }y\in Z\text{ and }i=1,\dots,d-1\right\}.$
-Additionally, we define $\nabla_{\tilde{x}}u=(\partial_{x_{1}}u,\dots,\partial_{x_{d-1}}u,0)$, where a 00 is appended such that the gradient it compatible with the usual gradient in $\mathbb{R}^{d}$.
-With $\|\cdot\|_{H^{1}(\Omega^{f}_{\varepsilon}\cup\Omega^{s}_{\varepsilon})}$ we will denote the combined $H^{1}$ norm on both subdomains, e.g.,
+With $n$, we denote the outer normal vector on the domains $\Omega,\Omega^{f,s}_{\varepsilon}$ and $Z$.
+On the common interface $\Gamma_{\varepsilon}$ we further write $n^{f,s}$ to denote the outer normal vector corresponding to $\Omega^{f,s}_{\varepsilon}$, but the superscript is suppressed whenever the direction is clear from context.
+Additionally, the normal vector $n_{\Sigma}$ is introduced in Equation 2 for the lateral boundaries of the fluid domain.
+With the subscript #\#, we indicate a space of periodic functions in the directions $1,\dots,d-1$; in particular, we use
+$C_{\#}(Z)$ $\coloneqq\left\{u\in C(\mathbb{R}^{d})\ :u(x+e_{i})=u(x)\text{ for all }x\in\mathbb{R}^{d}\text{ and }i=1,\dots,d-1\right\},$
+$L^{2}_{0,\#}(Z)$ $\coloneqq\left\{u\in L^{2}_{\text{loc}}(\mathbb{R}^{d})\ :\ u_{|Z}\in L^{2}_{0}(Z),\ u(x+e_{i})=u(x)\text{ for a.a. }x\in\mathbb{R}^{d}\text{ and }i=1,\dots,d-1\right\},$
+$H^{1}_{\#}(Z)$ $\coloneqq\left\{u\in H^{1}_{\text{loc}}(\mathbb{R}^{d}):u_{|Z}\in H^{1}(Z),u(x+e_{i})=u(x)\text{ for a.a. }x\in\mathbb{R}^{d}\text{ and }i=1,\dots,d-1\right\},$
+$H^{1}_{\#,\Gamma_{d}}(Z)$ $\coloneqq\left\{u\in H^{1}_{\#}(Z)\ :\ u=0\text{ for a.a. }y\in\Gamma_{d}\right\}.$
+Additionally, we define $\nabla_{\tilde{x}}u=(\partial_{x_{1}}u,\dots,\partial_{x_{d-1}}u,0)$, where a 0 is appended such that the gradient it compatible with the usual gradient in $\mathbb{R}^{d}$.
+With $\|\cdot\|_{H^{1}(\Omega^{f}_{\varepsilon}\cup\Omega^{s}_{\varepsilon})}$ we denote the combined $H^{1}$ norm on both subdomains, e.g.,
 $\|\cdot\|_{H^{1}(\Omega^{f}_{\varepsilon}\cup\Omega^{s}_{\varepsilon})}\coloneqq\sqrt{\|\cdot\|_{H^{1}(\Omega^{f}_{\varepsilon})}^{2}+\|\cdot\|_{H^{1}(\Omega^{s}_{\varepsilon})}^{2}}.$
-2.2 Underlying model equations
-For a function $\varphi\in L^{2}(\Omega)$, we use the notation $\varphi_{\varepsilon}^{k}:=\varphi_{|\Omega_{\varepsilon}^{k}}$ ($k=f,s$) for the restriction of $\varphi$ to the subdomain $\Omega_{\varepsilon}^{k}$.
+2.2 The micro model
+For a function $\varphi\in L^{2}(\Omega)$, we use the notation $\varphi_{\varepsilon}^{k}\coloneqq\varphi_{|\Omega_{\varepsilon}^{k}}$ ($k=f,s$) for the restriction of $\varphi$ to the subdomain $\Omega_{\varepsilon}^{k}$.
 A similar notation holds for the coefficients of the models, which are assumed to be constant in each subdomain.
 We suppress superscripts and the index $\varepsilon$ when possible to improve readability.
-We consider a reference frame of the grinding wheel, where instead of a moving interface $\Gamma_{\varepsilon}$ we have a prescribed velocity at the bottom boundary $\Sigma$.
 To describe the temperature $\theta_{\varepsilon}=(\theta^{s}_{\varepsilon},\theta^{f}_{\varepsilon})$, we utilize heat equations in both domains
-[…]
-[…]
-[…]
-Here, $\kappa=k_{f}$ in $\Omega^{f}_{\varepsilon}$ and $\kappa=\kappa_{s}$ in $\Omega^{s}_{\varepsilon}$ are the heat conductivities in each subdomain and $u_{\varepsilon}$ is the fluid velocity given via a Stokes system below.
+$\partial_{t}\theta_{\varepsilon}-\operatorname{div}\left(\kappa\nabla\theta_{\varepsilon}\right)$ $=0\quad\quad$ $\text{ in }S\times\Omega^{s}_{\varepsilon},$ (1a)
+$\frac{1}{\varepsilon}\partial_{t}\theta_{\varepsilon}-\frac{1}{\varepsilon}\operatorname{div}\left(\kappa\nabla\theta_{\varepsilon}-u_{\varepsilon}\theta_{\varepsilon}\right)$ $=0\quad\quad$ $\text{ in }S\times\Omega^{f}_{\varepsilon},$ (1b)
+$\theta_{\varepsilon}(0,\cdot)$ $=\theta_{\varepsilon,0}$ $\text{ in }\Omega.$ (1c)
+Here, $\kappa=\kappa_{f}$ in $\Omega^{f}_{\varepsilon}$ and $\kappa=\kappa_{s}$ in $\Omega^{s}_{\varepsilon}$ are the heat conductivities in each subdomain and $u_{\varepsilon}$ is the fluid velocity given via a Stokes system below.
 As they have no specific bearing on the mathematical analysis or the homogenization result, we have set the volumetric heat capacity to 1 in both phases.
 At the outer boundary, we apply a Dirichlet condition on $\Sigma^{\text{in}}_{\varepsilon}$ for a prescribed cooling temperature (for simplification set to 0, e.g., the temperature is seen relative to the inflow temperature, which is possible in this case since we always assume $\operatorname{div}{u_{\varepsilon}}=0$), and on the other boundary sections we apply a homogeneous Neumann condition:
-[…]
-[…]
+$\theta_{\varepsilon}$ $=0$ $\text{ on }S\times\Sigma^{\text{in}}_{\varepsilon},$ (1d)
+$-\kappa\nabla\theta_{\varepsilon}\cdot n$ $=0$ $\text{ on }S\times\left(\Sigma^{\text{out}}_{\varepsilon}\cup\Sigma^{s}_{\varepsilon}\cup\Sigma\right).$ (1e)
 The temperature profiles of the two domains are coupled with a Robin exchange condition:
-[…]
-[…]
-where $\llbracket\theta\rrbracket=(\theta^{f}-\theta^{s})$ denotes the usual jump across an interface, $\alpha$ denotes the heat exchange parameter, and $f_{\varepsilon}$ describes a heat source.
-Here, the chosen heat source is motivated by our underlying application problem of grinding, there the system heats up through friction between the workpiece and the grinding wheel.
-Therefore, the heat is produced at the interface.
-To simplify the notation the heat source is only located on the interface side of the solid domain.
-Through the Robin condition heat is then also transferred into the fluid.
-Note, the specific $\varepsilon$-scaling can be reasoned from a non-dimensionalization step and can be found in similar problems [20, 29] and originates from the scaling of the physical parameters (e.g., a fast heat conductivity in the thin layer compared to the height of the layer and the scaling of the density keeps the total amount of heat energy constant with respect to $\varepsilon$).
-Clearly Eq.
-2.1b could be multiplied with $\varepsilon$ and the scaling would then only enter on the right-hand side of (2.1f).
-For the following analysis, the above form is more appropriate.
+$-\frac{1}{\varepsilon}\kappa^{f}\nabla\theta^{f}_{\varepsilon}\cdot n^{f}$ $=\alpha\llbracket\theta_{\varepsilon}\rrbracket+f_{\varepsilon}^{f}\quad$ $\text{ on }S\times\Gamma_{\varepsilon},$ (1f)
+$-\kappa^{s}\nabla\theta^{s}_{\varepsilon}\cdot n^{s}$ $=\alpha\llbracket\theta_{\varepsilon}\rrbracket+f_{\varepsilon}^{s}\quad$ $\text{ on }S\times\Gamma_{\varepsilon},$ (1g)
+where $\llbracket\theta\rrbracket=(\theta^{f}-\theta^{s})$ denotes the usual jump across an interface, $\alpha$ denotes the heat exchange parameter, and $f_{\varepsilon}^{f,s}$ describe heat sources.
+Note, the chosen $\varepsilon$-scaling appears in similar problems [41, 22] and is designed to keep the contributions of the equations in the effect model for $\varepsilon\to 0$.
+From a physical point of view, the scaling originates from assumptions about the physical parameters (e.g., a fast heat conductivity in the thin layer compared to the height of the layer) and can be incorporated by a non-dimensionalization step; see for example [13, Section 2.1], where this is explained for fluid flow through thin cracks.
 Next to the temperature, we also model the fluid flow inside $\Omega^{f}_{\varepsilon}$.
 Fluid velocity $u_{\varepsilon}$ and pressure $p_{\varepsilon}$ are given as the solution of a Stokes system with temperature-dependent viscosity $\mu$,
-[…]
-[…]
-[…]
-[…]
+$-\operatorname{div}\left(\mu\left(\theta^{f}_{\varepsilon}\right)\nabla u_{\varepsilon}\right)+\nabla p_{\varepsilon}$ $=0\quad\quad$ $\text{ in }S\times\Omega^{f}_{\varepsilon},$ (1h)
+$\operatorname{div}\left(u_{\varepsilon}\right)$ $=0\quad\quad$ $\text{ in }S\times\Omega^{f}_{\varepsilon},$ (1i)
+$u_{\varepsilon}$ $=u_{\varepsilon,BC}$ $\text{ on }S\times\partial\Omega^{f}_{\varepsilon},$ (1j)
+$\int_{\Omega^{f}_{\varepsilon}}p_{\varepsilon}\,\mathrm{d}x$ $=0$ $\text{ in }S.$ (1k)
+Here $u_{\varepsilon,BC}\in H^{1/2}(\partial\Omega^{f}_{\varepsilon})\cap L^{\infty}(\partial\Omega^{f}_{\varepsilon})$ describes a given boundary flow.
 2.3 Assumptions on data
@@
   4. (A4)
-    For the heat source it holds $f_{\varepsilon}\in L^{\infty}(S\times\Gamma_{\varepsilon})$ and $C_{f}\coloneqq\sup_{\varepsilon>0}\|f_{\varepsilon}\|_{L^{\infty}(S\times\Gamma_{\varepsilon})}<\infty$.
+    For the heat sources it holds $f_{\varepsilon}^{f,s}\in L^{\infty}(S\times\Gamma_{\varepsilon})$ and $\sup_{\varepsilon>0}\|f_{\varepsilon}^{f,s}\|_{L^{\infty}(S\times\Gamma_{\varepsilon})}<\infty$.
   5. (A5)
-    For the prescribed velocity on the lateral fluid boundaries, we assume there is $u_{BC}\in W^{1,\infty}(\partial\Sigma\times(0,1))^{d}$ with the properties:
+    For the Dirichlet condition on the lateral fluid boundaries, we assume there is $u_{BC}\in H^{1/2}(\partial\Sigma\times(0,1))^{d}\cap L^{\infty}(\partial\Sigma\times(0,1))^{d}$ with the properties:
     • $u_{BC}\cdot e_{d}=0$,
-    • $u_{BC}(\tilde{x},0)=u_{\text{motion}}$, where it still holds that $u_{\text{motion}}\cdot e_{d}=0$,
+    • $u_{BC}(\tilde{x},0)=u_{\text{motion}}$, with $u_{\text{motion}}\in\mathbb{R}^{d}$ and $u_{\text{motion}}\cdot e_{d}=0$,
     • $u_{BC}(\tilde{x},x_{d})=0$ for almost all $\tilde{x}\in\partial\Sigma$ and $x_{d}\geq\gamma_{0}$,
-    • inflow and outflow conditions $\left(u_{BC}\cdot n_{\Sigma}\right)_{|\Sigma^{\text{in}}}<0\text{ and }\left(u_{BC}\cdot n_{\Sigma}\right)_{|\Sigma^{\text{out}}}\geq 0$.
+    • inflow and outflow conditions $\left(u_{BC}\cdot n_{\Sigma}\right)_{|\Sigma^{\text{in}}\times(0,1)}<0\text{ and }\left(u_{BC}\cdot n_{\Sigma}\right)_{|\Sigma^{\text{out}}\times(0,1)}\geq 0$.
     In addition, we require
-    $\int_{\partial\Sigma\times(0,1)}u_{BC}\cdot n_{\Sigma}\,\mathrm{d}\sigma_{x}=0$ (2.2)
-    to ensure the existence of solutions.
-Here, $n_{\Sigma}$ is the normal vector that points outward from $\Sigma\times(0,1)$.
-The boundary condition on the thin layer is given by $u_{\varepsilon,BC}(x)=u_{BC}(\tilde{x},x_{d}/\varepsilon)$ for $x\in\Sigma^{\text{in}}_{\varepsilon}\cup\Sigma^{\text{out}}_{\varepsilon}$.
+    $\int_{\partial\Sigma\times(0,1)}u_{BC}\cdot n_{\Sigma}\,\mathrm{d}\sigma_{x}=0$ (2)
+    to ensure the existence of solutions. Here, $n_{\Sigma}$ is the normal vector that points outward from $\Sigma\times(0,1)$.
   6. (A6)
     For the initial conditions there are limit functions $\theta^{s}_{0}\in L^{2}(\Omega)$ and $\theta^{f}_{0}\in L^{2}(\Sigma)$ such that $\chi_{\Omega^{s}_{\varepsilon}}\theta_{\varepsilon,0}\to\theta^{s}_{0}$ in $L^{2}(\Omega^{s}_{\varepsilon})$ and $\chi_{\Omega^{f}_{\varepsilon}}\theta^{f}_{\varepsilon,0}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\theta^{f}_{0}$ in $L^{2}(\Omega_{\varepsilon})$, for $\varepsilon\to 0$.
-Here, $\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}$ denotes two-scale convergence in the thin layer which will be introduced in Definition 1.
+Here, $\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}$ denotes two-scale convergence in the thin layer which is introduced in Definition 1.
   7. (A7)
-    For the heat source there is a $f\in L^{2}(S\times\Sigma\times\Gamma)$ such that $f_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}f$, for $\varepsilon\to 0$.
+    For the heat sources there are $f^{f,s}\in L^{2}(S\times\Sigma\times\Gamma)$ such that $f_{\varepsilon}^{f,s}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}f^{f,s}$, for $\varepsilon\to 0$.
 The Assumptions (A1) and (A2) are often encountered in similar problems, especially the properties of $\mu$ are needed to ensure the existence of a solution.
-Given the assumption (A5), we can always construct a continuation $U\in H^{1}(\Sigma\times(0,1))$ of $u_{BC}$, with $U(\tilde{x},y)=0$ for $y\geq\gamma_{0}$ and $\operatorname{div}{U}=0$, [32, Theorem.
+The assumption (A5) characterizes the boundary condition on the scaled boundary $\partial\Sigma\times(0,1)$, the microscopic boundary condition is defined via
+$u_{\varepsilon,BC}(x)\coloneqq\begin{cases}u_{BC}(\tilde{x},x_{d}/\varepsilon)\quad&\text{for }x\in\Sigma^{\text{in}}_{\varepsilon}\cup\Sigma^{\text{out}}_{\varepsilon},\\ u_{\text{motion}}&\text{for }x\in\Sigma,\\ 0&\text{for }x\in\Gamma_{\varepsilon}.\end{cases}$
+The inflow and outlfow conditions of $u_{BC}$ are transferred to $u_{\varepsilon,BC}$, to be precise
+$\left(u_{\varepsilon,BC}\cdot n\right)_{|\Sigma^{\text{in}}_{\varepsilon}}<0\quad\text{ and }\quad\left(u_{\varepsilon,BC}\cdot n\right)_{|\Sigma^{\text{out}}_{\varepsilon}}\geq 0.$ (3)
+Examples of functions that fulfill the assumptions can be found in Section 5.
+Furthermore, by assumption (A5), we can construct a continuation $U\in H^{1}(\Sigma\times(0,1))$ of $u_{BC}$, with $U(\tilde{x},y)=0$ for $y\geq\gamma_{0}$ and $\operatorname{div}{U}=0$, [42, Theorem.
 3.5].
-By setting $U_{\varepsilon}(x)=(U_{1},\dots,\varepsilon U_{d})(\tilde{x},\frac{x_{d}}{\varepsilon})$ the extension can also be transferred to the thin rough layer $\Omega^{f}_{\varepsilon}$, this construction was already derived in [12].
-We expect that the condition $u_{BC}=0$ for $x_{d}\geq\gamma_{0}$ could be removed and the extension still be carried out by first building an extension around the boundary falling back to a rectangle of height $\varepsilon\gamma_{0}$ and then proceeding as above, but this is not further investigated here.
-Examples of functions that fulfill the assumptions can be found in Section 5.
-We could also utilize a time-dependent boundary function $u_{BC}\in L^{\infty}(S;W^{1,\infty}(\partial\Sigma\times(0,1)))^{d}$, this would not change the following results.
-To simplify the notation we keep a time-independent $u_{BC}$.
-Additionally, given the prescribed boundary conditions of $u_{BC}$, it holds $u_{\varepsilon,BC}=u_{\text{motion}}$ on $\Sigma$ and we and we have inflow at $\Sigma^{\text{in}}_{\varepsilon}$ and outflow at $\Sigma^{\text{out}}_{\varepsilon}$, to be precise
-$\left(u_{\varepsilon,BC}\cdot n_{\varepsilon}\right)_{|\Sigma^{\text{in}}_{\varepsilon}}<0\quad\text{ and }\quad\left(u_{\varepsilon,BC}\cdot n_{\varepsilon}\right)_{|\Sigma^{\text{out}}_{\varepsilon}}\geq 0.$ (2.3)
+By setting $U_{\varepsilon}(x)=(U_{1},\dots,\varepsilon U_{d})(\tilde{x},\frac{x_{d}}{\varepsilon})$ the extension can also be transferred to the thin rough layer $\Omega^{f}_{\varepsilon}$ (this requires $u_{BC}\cdot e_{d}=0$ such that $U_{\varepsilon}$ has the correct boundary condition), this construction was already derived in [25].
 Remark 1.
-We provide some additional comments regarding our model and our mathematical assumptions specifically in regards to the application case of grinding:
+We provide some additional comments regarding our model and our mathematical assumptions:
   • $(i)$
-    During the actual physical process there is, of course, contact between the grinding wheel and the workpiece.
-In our model, that would imply that $\Omega^{s}_{\varepsilon}$ and $\Sigma$ are touching (i.e., $\gamma_{0}=0$).
-Note that for dimension $d=2$, the model would no longer be reasonable as any contact would necessarily disrupt fluid flow; while for $d=3$ it would still be possible for the fluid to flow around the contact points.
-In other works where the behavior of the fluid pressure inside the grinding gap is studied, an artificial gap is introduced such that the pressure can be modeled by the Reynolds equation [53, 55].
-Their findings demonstrate that this is a reasonable approximation to the real dynamics.
-Therefore, we keep this simplification of no direct contact and focus on how far a temperature-dependent viscosity influences the model.
+    $u_{\text{motion}}$ describes a prescribed velocity at the bottom boundary (for example, the movement induced by a rotating grinding wheel in our motivation).
+It is not mandatory that $u_{\text{motion}}$ is constant but could also be function in $H^{1/2}(\Sigma)^{d}\cap L^{\infty}(\Sigma)^{d}$ (as long as $u_{\text{motion}}\cdot e_{d}=0$).
+To simplify the notation, we use a constant $u_{\text{motion}}$.
+Similarly the boundary condition could also be time-dependent, $u_{\varepsilon,BC}\in L^{\infty}(S;H^{1/2}(\Omega^{f}_{\varepsilon})^{d}\cap L^{\infty}(\Omega^{f}_{\varepsilon}))$, as this would not influence the following results.
   • $(ii)$
-    The heat which is released during grinding is caused by the friction between the grains of the grinding wheel and the workpiece and only part of this heat is released into the system directly (e.g., part of it is also transported with the formation of small chips [48]).
-Here, we consider a simplified model where we assume a prescribed heat source at the interface but where we do not account for any mechanical effects in the solid part (in particular: no changing geometries, no chips, and no strains).
-  • $(iii)$
-    The relationship between temperature and viscosity of fluids is complex and often modeled heuristically. A typical example is the Vogel-Fulcher-Tammann equation given as
-    $\mu(\theta)=\mu_{0}\exp\left(\frac{\alpha}{\theta-T_{0}}\right)$
-    where $\mu_{0},\alpha,T_{0}$ are fitting parameters, we refer to [30] and references therein.
-This, of course, is only valid in a certain temperature range (usually a proper subset of the liquid phase range of the substance).
-However, it suggests that the relationship tends to be very smooth and that the Lipschitz continuity is a reasonable assumption.
-  • $(iv)$
     The specific assumptions regarding $u_{BC}$ and, as a consequence, $u_{\varepsilon,BC}$ are a bit technical, but they ensure that there is an inflow and an outflow region and that $\Sigma$ and $\Gamma_{\varepsilon}$ are in relative motion to each other.
 The regularity assumption ensures compatibility across the different boundary parts.
-The inflow condition models the continuous injection of cooling liquid where the corresponding Dirichlet condition Eq.
-2.1d accounts for the cooling effect.
-  • $(v)$
-    We do not include inertia, turbulence, and buoyancy effects by using the quasi-stationary Stokes system.
-This is somewhat justified by the thinness of the contact area where the fluid flow does not have enough space to create strong turbulence and is mainly pulled along by the moving wheel as demonstrated in the experiments in [21].
-We also neglect the real and important phenomenon of air entrainment into the fluid layer [34, 36].
-Additionally, the fluid movement adjusts on a shorter time interval than the temperature.
-3 The microscale model
+  • $(iii)$
+    The relationship between temperature and the viscosity of fluids is complex and often modeled heuristically. A commonly used example is the Vogel-Fulcher-Tammann equation, given by
+    $\mu(\theta)=\mu_{0}\exp\left(\frac{\alpha}{\theta-T_{0}}\right)$
+    where $\mu_{0},\alpha,T_{0}$ are empirical fitting parameters [5].
+The model suggests that the relationship between viscosity and temperature tends to be smooth which means that the Lipschitz continuity is a reasonable assumption.
+However, this function $\mu$ is not bounded from above and therefore does not satisfy Assumption (A2).
+In practice, this law is only considered valid within a specific temperature range—namely, the interval where the material remains in its liquid state.
+Therefore, it makes sense to restrict this law to such a temperature range and extend it to $\mathbb{R}$ by its boundary values, which also preserves Lipschitz regularity.
+Moreover, the $L^{\infty}$-estimates we establish for the temperature field (see Lemma 6) provide a posteriori a qualitative justification for this restriction.
+3 Analysis of the microscale model
 We start by analyzing the model on the $\varepsilon$-scale to ensure that it is well-posed and to derive solution bounds needed for the limiting procedure.
@@
 Lemma 1 (Extension operators and transformation).
+For all $\varepsilon>0$ and $p\in[1,\infty)$ there exists
   • $(i)$
-    For all $\varepsilon>0$, there exist a linear extension operator $\mathbb{E}_{\varepsilon}:W^{1,p}(\Omega^{f}_{\varepsilon})\to W^{1,p}(\Omega_{\varepsilon})$ such that
-    $\|\mathbb{E}_{\varepsilon}\psi\|_{W^{1,p}(\Omega_{\varepsilon})}\leq C\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}\quad\text{ for all }\psi\in W^{1,p}(\Omega^{f}_{\varepsilon})$
-    and for all $p\in[1,\infty)$. Here, $C>0$ may depend on $p$ but is independent of $\varepsilon$.
+    a linear extension operator $\mathbb{E}_{\varepsilon}^{f}:W^{1,p}(\Omega^{f}_{\varepsilon})\to W^{1,p}(\Omega_{\varepsilon})$ such that
+    $\|\mathbb{E}_{\varepsilon}^{f}\psi\|_{W^{1,p}(\Omega_{\varepsilon})}\leq C\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}\quad\text{ for all }\psi\in W^{1,p}(\Omega^{f}_{\varepsilon}).$
+    Here, $C>0$ may depend on $p$ but is independent of $\varepsilon$.
   • $(ii)$
-    There exists a diffeomorphism $\Phi:\Omega\to\Omega^{s}_{\varepsilon}$ such that
-    $\frac{1}{C}\|\varphi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}\leq\|\varphi\circ\Phi\|_{W^{1,p}(\Omega)}\leq C\|\varphi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}\quad\text{ for all }\varphi\in W^{1,p}(\Omega^{s}_{\varepsilon})$
-    and for all $p\in[1,\infty)$. Here, $C>0$ may depend on $p$ but is independent of $\varepsilon$.
+    a linear extension operator $\mathbb{E}_{\varepsilon}^{s}:W^{1,p}(\Omega^{s}_{\varepsilon})\to W^{1,p}(\Omega)$ such that
+    $\|\mathbb{E}_{\varepsilon}^{s}\varphi\|_{W^{1,p}(\Omega)}\leq C\|\varphi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}\quad\text{ for all }\varphi\in W^{1,p}(\Omega^{s}_{\varepsilon}),$
+    with $C>0$ depended on $p$ but independent of $\varepsilon$.
 Proof.
-$(i)$.
-The extension operators can be constructed using a classical extension procedure [1].
-For thin layers with a more general geometry, this was carried out in detail and further studied in [28, Theorem 1].
-The general idea is to first construct extension operators for a single cell, in our case from $W^{1,p}(Z)$ to $W^{1,p}(Y^{d})$.
-Those exist since $\gamma$ is Lipschitz-continuous, and they are naturally $\varepsilon$-independent (as there is no $\varepsilon$ involved yet).
-The statement itself then follows by gluing these extension operators together via a periodic partition of unity in conjunction with a scaling argument.
-$(ii)$. For the transformation of $\Phi$ to the complete domain $\Omega$, we define the functions
-$\phi_{1}(x)\coloneqq\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})+x_{d}\frac{H-\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}{H}\quad\text{ and }\quad\phi_{2}(x)\coloneqq H\frac{x_{d}-\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}{H-\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})},$
-where for a fixed $\tilde{x}\in\Sigma$ it holds
-$\phi_{1}(\tilde{x},\cdot):(0,H)\to(\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}}),H)\quad\text{ and }\quad\phi_{2}(\tilde{x},\cdot):(\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}}),H)\to(0,H).$
-With $\phi_{1}$ we obtain the transformation
-$\Phi:\Omega\to\Omega^{s}_{\varepsilon},\quad x\mapsto(\tilde{x},\phi_{1}(\tilde{x},x_{d}))$
-Since $\gamma$ is in $C^{1}(\overline{Y^{d-1}};[\gamma_{0},1])$ and $H>\varepsilon$ it holds that $G$ is bijective and continuously differentiable.
-The inverse is given by $\Phi^{-1}(x)=(\tilde{x},\phi_{2}(\tilde{x},x_{d}))$.
-Given the properties of $\gamma$, it holds
-$|\phi_{1}(x)|,|\phi_{2}(y)|\leq C,\quad|\nabla\phi_{1}(x)|,|\nabla\phi_{2}(y)|\leq C\quad\text{ and }\quad c\leq\det(\nabla\Phi),\det(\nabla\Phi^{-1})\leq C$
-for all $x\in\Omega,y\in\Omega^{s}_{\varepsilon}$, where $C,c>0$ can be chosen independent of $\varepsilon$.
-From these bounds we also obtain $\varepsilon$ independent bounds of $|\nabla\Phi|$ and $|\nabla\Phi^{-1}|$.
-The estimates in the $W^{1,p}$ norm for $\varphi$ and $\varphi\circ\Phi$ then follow via integration by substitution.
+Both extension operators can be constructed using a classical extension procedure [43].
+For thin layers, this was carried out in detail and further studied in [44, Theorem 1].
+See also [36, Lemma A.6] for a generalization to other boundary conditions in the case $p=2$.
+The general idea is to first construct an extension operator for a single cell, in the case of the fluid from $W^{1,p}(Z)$ to $W^{1,p}(Y^{d})$.
+Which exists since $\Gamma$ is Lipschitz.
+The extension is $\varepsilon$-independent (as there is no $\varepsilon$ involved yet).
+The statement is then obtained by gluing these extension operators together via a periodic partition of unity in conjunction with a scaling argument.
 ∎
@@
     For every $p\in[1,\infty)$, there is a constant $C>0$ (independent of $\varepsilon$) such that
-    $\|\psi\|_{L^{p}(\Gamma_{\varepsilon})}\leq\frac{C}{\varepsilon^{1/p}}\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}\quad\text{for all}\quad\psi\in W^{1,p}(\Omega^{f}_{\varepsilon})$
+    $\|\psi\|_{L^{p}(\Gamma_{\varepsilon})}^{p}\leq\frac{C}{\varepsilon}\left(\|\psi\|_{L^{p}(\Omega^{f}_{\varepsilon})}^{p}+\varepsilon^{p}\|\nabla\psi\|_{L^{p}(\Omega^{f}_{\varepsilon})}^{p}\right)\quad\text{for all}\quad\psi\in W^{1,p}(\Omega^{f}_{\varepsilon}).$
   • $(ii)$
     For every $p\in[1,\infty)$, there is $C>0$ (independent of $\varepsilon$) such that
-    $\|\phi\|_{L^{p}(\Gamma_{\varepsilon})}\leq C\|\phi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}\quad\text{for all}\quad\varphi\in W^{1,p}(\Omega^{s}_{\varepsilon})$
+    $\|\phi\|_{L^{p}(\Gamma_{\varepsilon})}\leq C\|\phi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}\quad\text{for all}\quad\varphi\in W^{1,p}(\Omega^{s}_{\varepsilon}).$
   • $(iii)$
@@
 $(i)$.
-This estimate for $\psi\in W^{1,p}(\Omega^{f}_{\varepsilon})$ is quite standard and follows by a splitting argument, which we shortly sketch.
-Split the integral over $\Gamma_{\varepsilon}$ into integrals over the boundary of the cell $\varepsilon Z$, applying a scaling argument followed by a trace estimate in the rescaled domain and then transforming back to the $\varepsilon$ domain.
-Note, that the estimate could be improved by a $\varepsilon^{1/p}$ in front of the gradient terms, but this is not needed in the following analysis.
-$(ii)$.
-A similar estimate was derived in [19, Proposition 2], albeit with homogeneous boundary conditions at the outer boundary $\partial\Omega^{s}_{\varepsilon}\setminus\Gamma_{\varepsilon}$ for $\varphi\in H^{1}(\Omega^{s}_{\varepsilon})$.
-By construction of the transformation $\Phi:\Omega\to\Omega^{s}_{\varepsilon}$ from Lemma 1, it holds $(\varphi\circ\Phi)_{|\Sigma}=\varphi_{|\Gamma_{\varepsilon}}$ and
-$\int_{\Gamma_{\varepsilon}}|\varphi(x)|^{p}\,\mathrm{d}\sigma_{x}$ $=\int_{\Sigma}\left|\varphi(\tilde{x},\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}}))\right|^{p}\sqrt{1+|\varepsilon\nabla_{\tilde{x}}\left[\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})\right]|^{2}}\,\mathrm{d}\sigma_{\tilde{x}}$
-$=\int_{\Sigma}\left|(\varphi\circ\Phi)(\tilde{x},0)\right|^{p}\sqrt{1+|\left[\nabla_{\tilde{x}}\gamma\right](\nicefrac{{\tilde{x}}}{{\varepsilon}})|^{2}}\,\mathrm{d}\sigma_{\tilde{x}}$
-$\leq C\int_{\Sigma}\left|(\varphi\circ\Phi)(\tilde{x},0)\right|^{p}\,\mathrm{d}\sigma_{\tilde{x}}.$
-On $\Omega$ we can use a trace theorem independent of $\varepsilon$ to obtain with Lemma 1.($ii$)
-$\|\varphi\|_{L^{p}(\Gamma_{\varepsilon})}\leq C\|\varphi\circ\Phi\|_{L^{p}(\Sigma)}\leq C\|\varphi\circ\Phi\|_{W^{1,p}(\Omega)}\leq C\|\varphi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}.$
-$(iii).$ This follows by a scaling argument. With the extension operator from Lemma 1, it holds
-$\|\psi\|_{L^{p}(\Sigma^{f}_{\varepsilon})}$ $=\|\mathbb{E}\psi\|_{L^{p}(\Sigma^{f}_{\varepsilon})}\leq\|\mathbb{E}\psi\|_{L^{p}(\partial\Sigma\times(0,\varepsilon))}=\varepsilon^{\frac{1}{p}}\|\mathbb{E}\psi(\tilde{x},\varepsilon x_{d})\|_{L^{p}(\partial\Sigma\times(0,1))}$
-$\leq C\varepsilon^{\frac{1}{p}}\|\mathbb{E}\psi(\tilde{x},\varepsilon x_{d})\|_{W^{1,p}(\Omega)}\leq C\|\mathbb{E}\psi\|_{W^{1,p}(\Omega_{\varepsilon})}\leq C\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}.$
+This estimate for $\psi\in W^{1,p}(\Omega^{f}_{\varepsilon})$ is quite standard in homogenization and follows by a splitting argument and a change of variables, see for example [45, Lemma 5] for a more detailed explanation.
+$(ii)$. This was proven in [36, Lemma A.5].
+$(iii).$ The estimate can be derived by a scaling argument. With the extension operator from Lemma 1, it holds
+$\|\psi\|_{L^{p}(\Sigma^{f}_{\varepsilon})}$ $=\|\mathbb{E}_{\varepsilon}^{f}\psi\|_{L^{p}(\Sigma^{f}_{\varepsilon})}\leq\|\mathbb{E}_{\varepsilon}^{f}\psi\|_{L^{p}(\partial\Sigma\times(0,\varepsilon))}=\varepsilon^{\frac{1}{p}}\|\mathbb{E}_{\varepsilon}^{f}\psi(\tilde{x},\varepsilon x_{d})\|_{L^{p}(\partial\Sigma\times(0,1))}$
+$\leq C\varepsilon^{\frac{1}{p}}\|\mathbb{E}_{\varepsilon}^{f}\psi(\tilde{x},\varepsilon x_{d})\|_{W^{1,p}(\Omega)}\leq C\|\mathbb{E}_{\varepsilon}^{f}\psi\|_{W^{1,p}(\Omega_{\varepsilon})}\leq C\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}.$
 Where we used the trace theorem on $\Omega$.
-Note, as before, the estimate could be improved by a $\varepsilon^{1/p}$ in front of the term with the $\partial_{x_{d}}$ norm, but this is not needed in the following analysis.
+Note, as before, the estimate could be improved by a $\varepsilon$ in front of the term with the $\partial_{x_{d}}$ norm, but this is not needed in the following analysis.
 ∎
@@
 For any $\psi\in L^{\infty}(S;L^{2}(\Omega_{\varepsilon})\cap L^{2}(S;H^{1}(\Omega_{\varepsilon}))$, let $\bar{\psi}\in L^{\infty}(S;L^{2}(\Omega))\cap L^{2}(S;H^{1}(\Omega))$ be the corresponding upscaled function given by $\bar{\psi}(t,\tilde{x},x_{d})=\psi(t,\tilde{x},\varepsilon x_{d})$.
-Using standard interpolation embeddings (see, e.g., [41, Chapter 2, Inequality 3.8]), it holds
+Using standard interpolation embeddings (see, e.g., [37, Chapter 2, Inequality 3.8]), it holds
 $\|\bar{\psi}\|_{L^{r_{d}}(S\times\Omega)}\leq C\left(\|\bar{\psi}\|_{L^{\infty}(S;L^{2}(\Omega))}+\|\bar{\psi}\|_{L^{2}(S;H^{1}(\Omega))}\right)$
@@
 Since $\nicefrac{{1}}{{r_{d}}}-\nicefrac{{1}}{{2}}=\nicefrac{{-1}}{{d+2}}$, the statement follows. ∎
-3.2 Analysis of the micro model
+To show the $L^{\infty}$ estimate of the temperature field, we use a truncation argument.
+In the following lemma, we state some properties of the time derivative of the truncated function in the case that the original time derivative is only an element in the dual space.
+We denote by $\langle\cdot,\cdot\rangle_{X^{\prime}}$ the dual pairing between elements from the space $X^{\prime}$ and $X$.
+Lemma 4.
+Let $\psi\in L^{2}(S;H^{1}(\Omega))\cap H^{1}(S;H^{1}(\Omega)^{\prime})$. For $k>0$ define the truncated function
+$\psi_{k}(t,x)\coloneqq\begin{cases}\psi(t,x)-k&\text{if }\psi(t,x)>k,\\ 0&\text{otherwise}.\end{cases}$
+Then $\psi_{k}\in L^{2}(S;H^{1}(\Omega))\cap H^{1}(S;H^{1}(\Omega)^{\prime})$.
+Furthermore, for all $\phi\in L^{2}(S,H^{1}(\Omega))$ with $\phi(t,x)=0$ if $\psi(t,x)\geq k$, it holds $\langle\partial_{t}\psi_{k},\phi\rangle_{L^{2}(S;H^{1}(\Omega)^{\prime})}=0$.
+If additionally $\psi(0)\in L^{\infty}(\Omega)$.
+Then for $k\geq\|\psi(0)\|_{L^{\infty}(\Omega)}$ and almost all $t\in S$ it holds
+$\int_{0}^{t}\langle\partial_{t}\psi(\tau),\psi_{k}(\tau)\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau=\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),\psi_{k}(\tau)\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau=\frac{1}{2}(\psi_{k}(t),\psi_{k}(t))_{L^{2}(\Omega)}.$ (4)
+Proof.
+That $\psi_{k}\in L^{2}(S;H^{1}(\Omega))\cap H^{1}(S;H^{1}(\Omega)^{\prime})$ follows via an approximation with a smooth mollification of $\psi$ and that the cut-off function is Lipschitz continuous.
+This mollification also directly yields the property $\langle\partial_{t}\psi_{k},\phi\rangle_{L^{2}(S;H^{1}(\Omega)^{\prime})}=0$ for $\phi=0$ if $\psi\geq k$.
+For the integral equality (4), we use the fundamental theorem of calculus (with $k\geq\|\psi(0)\|_{L^{\infty}(\Omega)}$) to obtain
+$(\psi_{k}(t),\psi_{k}(t))_{L^{2}(\Omega)}=2\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),\psi_{k}(\tau)\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau.$ (5)
+Furthermore, it holds
+$\begin{split}(\psi_{k}(t),\psi_{k}(t))_{L^{2}(\Omega)}&=(\chi_{\{\psi(t)>k\}}(\psi(t)-k),\psi_{k}(t))_{L^{2}(\Omega)}\\ &=(\psi(t)-k,\psi_{k}(t))_{L^{2}(\Omega)}\\ &=\int_{0}^{t}\langle\partial_{t}\psi(\tau),\psi_{k}(\tau)\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau+\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),\psi(\tau)-k\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau.\end{split}$ (6)
+In the last integral, we can split $\psi-k$ into positive and negative parts. Since $(\psi-k)_{-}=0$ if $\psi\geq k$, this leads to
+$\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),\psi(\tau)-k\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau$ $=\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),(\psi(\tau)-k)_{+}\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau+\underbrace{\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),(\psi(\tau)-k)_{-}\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau}_{=0}$
+$=\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),\psi_{k}(\tau)\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau.$
+Plugging the last equation into Equation 6 and comparing with Equation 5 yields the desired result. ∎
+3.2 Well-posedness and a prior estimates
 Next, we investigate the underlying model on the microscale. First, we define
@@
 For the solutions of the temperature space we define $\Theta_{\varepsilon}\coloneqq H^{1}(\Omega^{s}_{\varepsilon})\times H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})$.
-For the fluid velocity and pressure we consider
-$V_{\varepsilon}\coloneqq\left\{v\in L^{\infty}(S;H^{1}(\Omega^{f}_{\varepsilon})^{d}):v=u_{\varepsilon,BC}\text{ on }\partial\Omega^{f}_{\varepsilon}\text{ and }\operatorname{div}{v}=0\text{ a.e.
-in }\Omega^{f}_{\varepsilon}\right\}\quad\text{ and }\quad Q_{\varepsilon}\coloneqq L^{\infty}(S;L^{2}_{0}(\Omega^{f}_{\varepsilon})).$
-The weak formulation of the heat equation, for a fixed $v\in V_{\varepsilon}$, is given by: Find $\theta_{\varepsilon}=(\theta^{s}_{\varepsilon},\theta^{f}_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})$ such that $\partial_{t}\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$ and for all $\varphi\in L^{2}(S;\Theta_{\varepsilon})$ and almost all $t\in S$
-$\begin{split}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{s}_{\varepsilon}}&+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{f}_{\varepsilon}}\\ &+\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}+\alpha(\llbracket\theta_{\varepsilon}\rrbracket,\llbracket\varphi\rrbracket)_{\Gamma_{\varepsilon}}=(f_{\varepsilon},\varphi^{s})_{\Gamma_{\varepsilon}}.\end{split}$ (3.1)
+The dual of a finite product space is isometrically isomorph to the product of the dual spaces.
+For any element $\psi\in\Theta_{\varepsilon}^{\prime}$, there are unique $\psi^{s}\in H^{1}(\Omega^{s}_{\varepsilon})^{\prime}$, $\psi^{f}\in H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}$ such that for all $\varphi\in\Theta_{\varepsilon}$
+$\langle\psi,\varphi\rangle_{\Theta_{\varepsilon}^{\prime}}=\langle\psi^{s},\varphi^{s}\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+\langle\psi^{f},\varphi^{f}\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}.$
+In the following, we identify $\psi=(\psi^{s},\psi^{f})\in\Theta_{\varepsilon}^{\prime}$ and suppress the superscript when the context is clear from the dual pairing.
+For the fluid velocity and pressure, we define
 […]
-$\begin{split}&(\mu(\psi_{\varepsilon})\nabla u_{\varepsilon},\nabla v)_{\Omega^{f}_{\varepsilon}}-(p_{\varepsilon},\operatorname{div}{(v)})_{\Omega^{f}_{\varepsilon}}=0.\end{split}$ (3.2)
+$\hskip 85.35826ptQ_{\varepsilon}\coloneqq L^{\infty}(S;L^{2}_{0}(\Omega^{f}_{\varepsilon})).$
+The weak formulation of the heat equation, for a fixed $v\in V_{\varepsilon}$, is given by: Find $\theta_{\varepsilon}=(\theta^{s}_{\varepsilon},\theta^{f}_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})$ such that $\partial_{t}\theta_{\varepsilon}=(\partial_{t}\theta_{\varepsilon}^{s},\partial_{t}\theta_{\varepsilon}^{f})\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$ and for all $\varphi\in L^{2}(S;\Theta_{\varepsilon})$ and almost all $t\in S$
+$\begin{split}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}&+(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{s}_{\varepsilon}}+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{f}_{\varepsilon}}\\ &+\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}+\alpha(\llbracket\theta_{\varepsilon}\rrbracket,\llbracket\varphi\rrbracket)_{\Gamma_{\varepsilon}}=(f_{\varepsilon}^{s},\varphi^{s})_{\Gamma_{\varepsilon}}+(f_{\varepsilon}^{f},\varphi^{f})_{\Gamma_{\varepsilon}}.\end{split}$ (7)
+Here, $(\cdot,\cdot)_{D}$ denotes the usual $L^{2}$ inner product over the domain $D$.
+Given a $\psi_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$, the weak form of the Stokes equation is: Find $(u_{\varepsilon},p_{\varepsilon})\in V_{\varepsilon}\times Q_{\varepsilon}$ such that for all $v\in H^{1}_{0}(\Omega^{f}_{\varepsilon})^{d}$ and almost all $t\in S$
+$\begin{split}&(\mu(\psi_{\varepsilon})\nabla u_{\varepsilon},\nabla v)_{\Omega^{f}_{\varepsilon}}-(p_{\varepsilon},\operatorname{div}{(v)})_{\Omega^{f}_{\varepsilon}}=0.\end{split}$ (8)
 For a fixed $\varepsilon>0$ we now prove that there exists a solution of the coupled nonlinear system given by Eq.
-3.1 and (3.2).
+7 and (8).
 To show this, we follow a similar strategy as in [6], where we split up the problem, first consider both weak formulations on their own, and then show a solution of the coupled system by a fix-point argument of an iterative scheme.
@@
 and Eq.
-3.1.
+7.
 In addition, it holds
-$\begin{split}\|\theta_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{s}_{\varepsilon}))}+\frac{1}{\sqrt{\varepsilon}}\|\theta_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\|\nabla\theta_{\varepsilon}\|_{L^{2}(S;L^{2}(\Omega^{s}_{\varepsilon}))}+\frac{1}{\sqrt{\varepsilon}}\|\nabla\theta_{\varepsilon}\|_{L^{2}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\|\llbracket\theta_{\varepsilon}\rrbracket\|_{L^{2}(S\times\Gamma_{\varepsilon})}\leq C,\end{split}$ (3.3)
+$\begin{split}\|\theta_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{s}_{\varepsilon}))}&+\frac{1}{\sqrt{\varepsilon}}\|\theta_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\|\nabla\theta_{\varepsilon}\|_{L^{2}(S;L^{2}(\Omega^{s}_{\varepsilon}))}\\ &+\frac{1}{\sqrt{\varepsilon}}\|\nabla\theta_{\varepsilon}\|_{L^{2}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\|\llbracket\theta_{\varepsilon}\rrbracket\|_{L^{2}(S\times\Gamma_{\varepsilon})}\leq C,\end{split}$ (9)
 where $C<\infty$ can be chosen independent of $\varepsilon$ and $v$.
 Proof.
-We start by showing the existence and uniqueness of a solution, for a fixed $v\in V_{\varepsilon}$. First, note that our problem is of the form
-$\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+A_{\varepsilon}(\theta_{\varepsilon},\varphi;t)=(f_{\varepsilon},\varphi^{s})_{\Gamma_{\varepsilon}},$ (3.4)
+We start by showing the existence and uniqueness of a solution for a fixed $v\in V_{\varepsilon}$. First, note that our problem is of the form
+$\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+A_{\varepsilon}(\theta_{\varepsilon},\varphi;t)=(f_{\varepsilon}^{s},\varphi^{s})_{\Gamma_{\varepsilon}}+(f_{\varepsilon}^{f},\varphi^{f})_{\Gamma_{\varepsilon}},$ (10)
 with the time-dependent bilinear form
@@
 $|A_{\varepsilon}(\psi,\varphi;t)|\leq C\left(1+\frac{1}{\varepsilon}+\|v\|_{L^{\infty}(S;H^{1}(\Omega^{f}_{\varepsilon}))}\right)\|\psi(t)\|_{H^{1}(\Omega^{f}_{\varepsilon}\cup\Omega^{s}_{\varepsilon})}\|\phi(t)\|_{H^{1}(\Omega^{f}_{\varepsilon}\cup\Omega^{s}_{\varepsilon})},$
-for a $C<\infty$. Additionally, via integration by parts we have (suppressing the time variable $t$)
-$(v\cdot\nabla\psi,\varphi)_{\Omega^{f}_{\varepsilon}}=-(v\cdot\nabla\varphi,\psi)_{\Omega^{f}_{\varepsilon}}-(\operatorname{div}{(v)}\psi,\varphi)_{\Omega^{f}_{\varepsilon}}+(v\cdot n_{\varepsilon},\psi\varphi)_{\partial\Omega^{f}_{\varepsilon}}.$ (3.5)
-Plugging in $\varphi=\psi$, that $\operatorname{div}{(v)}=0$ and the boundary conditions on $\partial\Omega^{f}_{\varepsilon}$ into Eq. 3.5 leads to
-$(v\cdot\nabla\psi,\psi)_{\Omega^{f}_{\varepsilon}}=\frac{1}{2}\int_{\Sigma^{\text{out}}_{\varepsilon}}u_{\varepsilon,BC}\cdot n_{\varepsilon}\,\psi^{2}\,\mathrm{d}\sigma_{x}\geq 0$
+for a $C<\infty$. Additionally, via integration by parts, we have (suppressing the time variable $t$)
+$(v\cdot\nabla\psi,\varphi)_{\Omega^{f}_{\varepsilon}}=-(v\cdot\nabla\varphi,\psi)_{\Omega^{f}_{\varepsilon}}-(\operatorname{div}{(v)}\psi,\varphi)_{\Omega^{f}_{\varepsilon}}+(v\cdot n,\psi\varphi)_{\partial\Omega^{f}_{\varepsilon}}.$ (11)
+Plugging in $\varphi=\psi$, that $\operatorname{div}{(v)}=0$ and the boundary conditions on $\partial\Omega^{f}_{\varepsilon}$ into Eq. 11 leads to
+$(v\cdot\nabla\psi,\psi)_{\Omega^{f}_{\varepsilon}}=\frac{1}{2}\int_{\Sigma^{\text{out}}_{\varepsilon}}u_{\varepsilon,BC}\cdot n\,\psi^{2}\,\mathrm{d}\sigma_{x}\geq 0$
 for almost all $t\in S$. Therefore, there exists a $\lambda>0$ such that
@@
 for all $\psi\in L^{2}(S;\Theta_{\varepsilon})$ and almost all $t\in S$.
-Given the representation (3.4) and the above estimates for the bilinear form $A_{\varepsilon}$, there exists a unique solution $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ of the problem [52, Chapter 3.2].
-For the norm estimate, we utilize the above computations of the convection term, use $\varphi=\theta_{\varepsilon}$ in the weak formulation and integrate afterward over the time interval $(0,t)$, for $t>0$, to obtain
-$\begin{split}\frac{1}{2}\|\theta_{\varepsilon}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}&+\frac{1}{2\varepsilon}\|\theta_{\varepsilon}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon}\|_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}^{2}+\alpha\|\llbracket\theta_{\varepsilon}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\\ &\leq C\|f_{\varepsilon}\|_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\|\theta_{\varepsilon}\|_{L^{2}((0,t);H^{1}(\Omega^{s}_{\varepsilon}))}+\frac{1}{2}\|\theta_{\varepsilon,0}\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}+\frac{1}{2\varepsilon}\|\theta_{\varepsilon,0}\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}.\end{split}$
-Applying Young’s inequality followed by Gronwall’s inequality leads to the desired estimate (3.3), where $C$ is independent of $\varepsilon$. ∎
-Remark 3.
-Note, the condition in (2.3) is used in the previous Lemma to derive that $(v\cdot\nabla\psi,\psi)_{\Omega^{f}_{\varepsilon}}\geq 0$, this is not a necessary condition to derive coercivity and carry out the energy estimates.
-Given the trace inequality and the compact embedding of $H^{1}(\Omega^{f}_{\varepsilon})$ into $L^{2}(\partial\Omega^{f}_{\varepsilon})$, the terms stemming from the convection terms (after integration by part) could be included in the already present gradient terms and $H^{1}$-norm.
-But (2.3) is a reasonable assumption in our case and let us simplify the notation.
+Given the representation (10) and the above estimates for the bilinear form $A_{\varepsilon}$, there exists a unique solution $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ [46, Chapter 3, Proposition 2.3].
+For the norm estimate, we utilize the above computations of the convection term, use $\varphi=\theta_{\varepsilon}$ in the weak formulation, and integrate afterward over the time interval $(0,t)$ to obtain (with the trace estimates from Lemma 2)
+$\begin{split}\frac{1}{2}\|\theta_{\varepsilon}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}&+\frac{1}{2\varepsilon}\|\theta_{\varepsilon}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon}\|_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}^{2}+\alpha\|\llbracket\theta_{\varepsilon}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\\ &\leq C\|f^{s}_{\varepsilon}\|_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\|\theta_{\varepsilon}\|_{L^{2}((0,t);H^{1}(\Omega^{s}_{\varepsilon}))}+\frac{C}{\sqrt{\varepsilon}}\|f^{f}_{\varepsilon}\|_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\|\theta_{\varepsilon}\|_{L^{2}((0,t);H^{1}(\Omega^{f}_{\varepsilon}))}\\ &\hskip 10.0pt+\frac{1}{2}\|\theta_{\varepsilon,0}\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}+\frac{1}{2\varepsilon}\|\theta_{\varepsilon,0}\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}.\end{split}$
+Applying Young’s inequality followed by Gronwall’s inequality leads to the desired estimate (9), where $C$ is independent of $\varepsilon$. ∎
 Lemma 6.
-Let the same assumptions be fulfilled as in Lemma 4 and let $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ be the solution of Eq.
-3.1.
+Let the same assumptions as in Lemma 5 be fulfilled and let $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ be the solution of Eq.
+7.
 Then there exists a $C>0$ independent of $\varepsilon$ such that $\|\theta_{\varepsilon}\|_{L^{\infty}(S\times\Omega)}\leq C.$
 Proof.
-For any $k>0$ construct the functions
+To show the estimate, we observe that the solution is linear in the source terms $f_{\varepsilon}^{s,f}$.
+Therefore, we can show the estimates for the two cases $f_{\varepsilon}^{s}=0,f_{\varepsilon}^{f}\neq 0$ and $f_{\varepsilon}^{s}\neq 0,f_{\varepsilon}^{f}=0$ separately.
+Both cases follow via similar arguments, where we present the case $f_{\varepsilon}^{s}=0,f_{\varepsilon}^{f}\neq 0$ and in the end shortly comment on the other case.
+To avoid cluttering, we do not introduce an additional index to denote this case.
+First, for any $k>0$ construct the functions
 $\theta_{\varepsilon,k}^{s}\coloneqq\begin{cases}\theta_{\varepsilon}^{s}-k\quad&\text{if }\theta_{\varepsilon}^{s}>k,\\ 0\quad&\text{if }|\theta_{\varepsilon}^{s}|\leq k,\\ \theta_{\varepsilon}^{s}+k\quad&\text{if }\theta_{\varepsilon}^{s}<-k,\end{cases}\quad\text{and}\quad\theta_{\varepsilon,k}^{f}\coloneqq\begin{cases}\theta_{\varepsilon}^{f}-k\quad&\text{if }\theta_{\varepsilon}^{f}>k,\\ 0\quad&\text{if }|\theta_{\varepsilon}^{f}|\leq k,\\ \theta_{\varepsilon}^{f}+k\quad&\text{if }\theta_{\varepsilon}^{f}<-k,\end{cases}$
@@
 $A_{\varepsilon,k}^{s}(\tau)\coloneqq\left\{x\in\Omega_{\varepsilon}^{s}:|\theta_{\varepsilon}^{s}(\tau,x)|>k\right\}\quad\text{and}\quad A_{\varepsilon,k}^{f}(\tau)\coloneqq\left\{x\in\Omega_{\varepsilon}^{f}:|\theta_{\varepsilon}^{f}(\tau,x)|>k\right\}.$
-Note that it holds $\partial_{t}\theta_{\varepsilon,k}^{s}=\chi_{A_{\varepsilon,k}^{s}}\partial_{t}\theta_{\varepsilon}^{s}$ and $\nabla\theta_{\varepsilon,k}^{s}=\chi_{A_{\varepsilon,k}^{s}}\nabla\theta_{\varepsilon}^{s}$ (similiar for $\theta_{\varepsilon,k}^{f}$).
+Note $\nabla\theta_{\varepsilon,k}^{s,f}=\chi_{A_{\varepsilon,k}^{s,f}}\nabla\theta_{\varepsilon,k}^{s,f}$.
+Since $\theta_{\varepsilon,k}^{s,f}=(\theta_{\varepsilon}^{s,f}-k)_{+}+(\theta_{\varepsilon}^{s,f}+k)_{-}$, applying Lemma 4 to both the positive and negative part yields, $\partial_{t}\theta_{\varepsilon,k}^{s,f}\in L^{2}(S;H^{1}(\Omega^{s,f}_{\varepsilon})^{\prime})$.
 By Assumption (A3) we set $k>C_{0}=\sup_{\varepsilon>0}\|\theta_{\varepsilon,0}\|_{L^{\infty}(\Omega)}$.
 Next, we take $\theta_{\varepsilon,k}=(\theta_{\varepsilon,k}^{s},\theta_{\varepsilon,k}^{f})$ as a test function in Eq.
-3.1.
-Integrating over the time interval $(0,t)\subset S$, we obtain:
-$\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}$ $+\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}$
-$+\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}+\alpha\|\llbracket\theta_{\varepsilon,k}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}=(f_{\varepsilon},\theta_{\varepsilon,k}^{s})_{(0,t)\times\Gamma_{\varepsilon}}$
-Since $\nabla\theta_{\varepsilon,k}^{f}=\chi_{A_{\varepsilon,k}^{f}}\nabla\theta_{\varepsilon}^{f}$, the convection is
+7, leading, for almost all $\tau\in S$, to
+$\begin{split}\langle\partial_{t}\theta^{s}_{\varepsilon}(\tau),\theta_{\varepsilon,k}^{s}(\tau)\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}&+\frac{1}{\varepsilon}\langle\partial_{t}\theta^{f}_{\varepsilon}(\tau),\theta_{\varepsilon,k}^{f}(\tau)\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+\kappa^{s}\|\nabla\theta_{\varepsilon,k}^{s}(\tau)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}(\tau)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}\\ &+\frac{1}{\varepsilon}(v(\tau)\cdot\nabla\theta_{\varepsilon}(\tau),\theta_{\varepsilon,k}^{f}(\tau))_{\Omega^{f}_{\varepsilon}}+\alpha\|\llbracket\theta_{\varepsilon,k}(\tau)\rrbracket\|^{2}_{L^{2}(\Gamma_{\varepsilon})}=(f^{f}_{\varepsilon}(\tau),\theta_{\varepsilon,k}^{f}(\tau))_{\Gamma_{\varepsilon}}.\end{split}$
+Integrating over the time interval $(0,t)\subset S$ and using Equation 4 from Lemma 4 we obtain
+$\frac{1}{2}\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}$ $+\frac{1}{2\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}$
+$+\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}+\alpha\|\llbracket\theta_{\varepsilon,k}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}=(f_{\varepsilon}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Gamma_{\varepsilon}}$
+Since $\nabla\theta_{\varepsilon,k}^{f}=\chi_{A_{\varepsilon,k}^{f}}\nabla\theta_{\varepsilon}^{f}$, for the the convection it holds
 $\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}=\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}.$
-Eq. 3.5 and $\operatorname{div}{v}=0$ yield
-$\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}=\frac{1}{2\varepsilon}(v\cdot\nu\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\partial\Omega^{f}_{\varepsilon}}=\frac{1}{2\varepsilon}(v\cdot\nu\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\partial\Sigma^{\text{out}}_{\varepsilon}}\geq 0.$
+Eq. 11 and $\operatorname{div}{v}=0$ yield
+$\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}=\frac{1}{2\varepsilon}(v\cdot n\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\partial\Omega^{f}_{\varepsilon}}=\frac{1}{2\varepsilon}(v\cdot n\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\partial\Sigma^{\text{out}}_{\varepsilon}}\geq 0.$
 Therefore we obtain
-$\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}$ $+\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}$
-$+\alpha\|\llbracket\theta_{\varepsilon,k}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\leq(f_{\varepsilon},\theta_{\varepsilon,k}^{s})_{(0,t)\times\Gamma_{\varepsilon}}.$
+$\begin{split}\frac{1}{2}\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}&+\frac{1}{2\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}\\ &+\alpha\|\llbracket\theta_{\varepsilon,k}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\leq(f_{\varepsilon}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Gamma_{\varepsilon}}.\end{split}$ (12)
 By Assumption (A4) and the trace estimate in Lemma 2 the source term on the right-hand side can be estimated by
-$\begin{split}\left|(f_{\varepsilon},\theta_{\varepsilon,k}^{s})_{(0,t)\times\Gamma_{\varepsilon}}\right|&\leq\|f_{\varepsilon}\|_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}\int_{0}^{t}\int_{\Gamma_{\varepsilon}}|\theta_{\varepsilon,k}^{s}|\,\mathrm{d}x\,\mathrm{d}\tau\\ &\leq C\|f_{\varepsilon}\|_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}\left(\|\theta_{\varepsilon,k}^{s}\|_{L^{1}((0,t)\times\Omega^{s}_{\varepsilon})}+\|\nabla\theta_{\varepsilon,k}^{s}\|_{L^{1}((0,t)\times\Omega^{s}_{\varepsilon})}\right)\\ &\leq\|f_{\varepsilon}\|^{2}_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}C_{\delta}\|1\|_{L^{2}((0,t)\times A_{\varepsilon,k}^{s})}^{2}+C\|\theta_{\varepsilon,k}^{s}\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}+\delta\|\nabla\theta_{\varepsilon,k}^{s}\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}.\end{split}$ (3.6)
+$\begin{split}\left|(f_{\varepsilon}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Gamma_{\varepsilon}}\right|&\leq\|f_{\varepsilon}^{f}\|_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}\int_{0}^{t}\int_{\Gamma_{\varepsilon}}|\theta_{\varepsilon,k}^{f}|\,\mathrm{d}x\,\mathrm{d}\tau\\ &\leq\frac{C}{\varepsilon}\|f_{\varepsilon}^{f}\|_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}\left(\|\theta_{\varepsilon,k}^{f}\|_{L^{1}((0,t)\times\Omega^{f}_{\varepsilon})}+\|\nabla\theta_{\varepsilon,k}^{f}\|_{L^{1}((0,t)\times\Omega^{f}_{\varepsilon})}\right)\\ &\leq\frac{C_{\delta}}{\varepsilon}\|f_{\varepsilon}^{f}\|^{2}_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}\|1\|_{L^{2}((0,t)\times A_{\varepsilon,k}^{f})}^{2}+\frac{C}{\varepsilon}\|\theta_{\varepsilon,k}^{f}\|_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}^{2}+\frac{\delta}{\varepsilon}\|\nabla\theta_{\varepsilon,k}^{f}\|_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}^{2}.\end{split}$ (13)
 In the last step, Young’s inequality with a $\delta>0$ was used.
 We choose $\delta$ such that we can include the gradient term on the left side and additionally assume $k\geq\|f_{\varepsilon}\|_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}$, to arrive at
-$\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}$ $+\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}$
-$\leq C\left[\|\theta_{\varepsilon,k}^{s}\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|A_{\varepsilon,k}^{s}(\tau)|\,\mathrm{d}\tau\right]$
-Isolating the terms that are connected to the temperature of the solid, we have
-$\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}+\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}$ $\leq C\left[\|\theta_{\varepsilon,k}^{s}\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|A_{\varepsilon,k}^{s}(\tau)|\,\mathrm{d}\tau\right].$ (3.7)
-With the transformation $\Phi$ from Lemma 1 we can switch to the fixed domain $\Omega$.
-Note that $x\in\Omega\setminus\Phi(A_{\varepsilon,k}^{s}(\tau))$ iff $\theta_{\varepsilon,k}^{s}\circ\Phi(x,\tau)=0$ for a.a.
-$\tau\in(0,t)$.
-And since $\det{(\nabla\Phi)}(x)=({H-\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})})/{H}>0$, we conclude from Eq.
-3.7, that also
-$\|(\theta_{\varepsilon,k}^{s}\circ\Phi)(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}+\|\nabla(\theta_{\varepsilon,k}^{s}\circ\Phi)\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}$ $\leq C\left[\|\theta_{\varepsilon,k}^{s}\circ\Phi\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|\Phi(A_{\varepsilon,k}^{s}(\tau))|\,\mathrm{d}\tau\right].$
-By Gronwall’s inequality, we arrive at
-$\|\theta_{\varepsilon,k}^{s}\circ\Phi\|_{L^{\infty}(S;L^{2}(\Omega))\cap L^{2}(S;H^{1}(\Omega))}\leq Ck\left(\int_{S}|\Phi(A_{\varepsilon,k}^{s}(\tau))|\,\mathrm{d}\tau\right)^{\frac{1}{2}}$
-The domain $\Omega$ is independent of $\varepsilon$, therefore applying the results of [41, Chapter 2, Theorem 6.1] we obtain $\|\theta^{s}_{\varepsilon}\circ\Phi\|_{L^{\infty}(S\times\Omega)}\leq C$, where $C>0$ can be chosen independent of $\varepsilon$.
-By the properties of $\Phi$ this immediately leads to $\|\theta^{s}_{\varepsilon}\|_{L^{\infty}(S\times\Omega^{s}_{\varepsilon})}\leq C$ and by the properties of the trace also to $|\theta^{s}_{\varepsilon|\Gamma_{\varepsilon}}|=|(\theta^{s}_{\varepsilon}\circ\Phi)_{|\Sigma}|\leq C$.
-For the estimate in the fluid domain, test with $(0,\theta^{f}_{\varepsilon,k})$, with similar computations as above we arrive at
-$\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}+\alpha\|\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}=(\theta^{s}_{\varepsilon},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Gamma_{\varepsilon}}$
-The right-hand side can bounded by the same steps as in Eq.
-3.6.
-Note, that the trace estimate from Lemma 2 introduces an additional $\nicefrac{{1}}{{\varepsilon}}$.
-Setting $k\geq\|\theta^{s}_{\varepsilon}\|_{L^{\infty}(S\times\Omega^{s}_{\varepsilon})}$, we have
 $\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\frac{1}{\varepsilon}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}\leq\frac{C}{\varepsilon}\left[\|\theta_{\varepsilon,k}^{f}\|_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|A_{\varepsilon,k}^{f}(\tau)|\,\mathrm{d}\tau\right].$
-With the extension operator $\mathbb{E}_{\varepsilon}$ from Lemma 1, we transform to the non-rough thin layer.
-Define $\mathbb{E}_{\varepsilon}{A}^{f}_{\varepsilon,k}(\tau)=\{x\in\Omega_{\varepsilon}:\mathbb{E}_{\varepsilon}\theta^{f}_{e,k}(t,x)\neq 0\}$, clearly $|A^{f}_{\varepsilon,k}(\tau)|\leq|\mathbb{E}_{\varepsilon}A^{f}_{\varepsilon,k}(\tau)|$.
+With the extension operator $\mathbb{E}_{\varepsilon}^{f}$ from Lemma 1, we transform to the non-rough thin layer.
+Define $B^{f}_{\varepsilon,k}(\tau)\coloneqq\{x\in\Omega_{\varepsilon}:\mathbb{E}_{\varepsilon}^{f}\theta^{f}_{e,k}(\tau,x)\neq 0\}$, then clearly $|A^{f}_{\varepsilon,k}(\tau)|\leq|B^{f}_{\varepsilon,k}(\tau)|$.
 Multiplying by $\varepsilon$, we obtain that on the non-rough thin layer, it holds
-$\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega_{\varepsilon})}+\kappa^{f}\|\nabla\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega_{\varepsilon})}\leq C\left[\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{2}((0,t)\times\Omega_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|\mathbb{E}_{\varepsilon}A_{\varepsilon,k}^{f}(\tau)|\,\mathrm{d}\tau\right],$
+$\|\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega_{\varepsilon})}+\|\nabla\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega_{\varepsilon})}\leq C\left[\|\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon,k}^{f}\|_{L^{2}((0,t)\times\Omega_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|B_{\varepsilon,k}^{f}(\tau)|\,\mathrm{d}\tau\right],$
 and consequently by Gronwall’s inequality
-$\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S;L^{2}(\Omega_{\varepsilon}))\cap L^{2}(S;H^{1}(\Omega_{\varepsilon}))}\leq Ck\left(\int_{S}|\mathbb{E}_{\varepsilon}A_{\varepsilon,k}^{f}(\tau)|\,\mathrm{d}\tau\right)^{\frac{1}{2}}.$
-Applying again [41, Chapter 2, Theorem 6.1], we now have to be careful regarding the $\varepsilon$ dependency of the estimate. First we obtain
-$\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega_{\varepsilon})}\leq C\left(1+C\beta_{\varepsilon}^{\frac{d+2}{2}}|S|^{\frac{1}{2}}|\Omega_{\varepsilon}|^{\frac{1}{2}}\right)\leq C\left(1+C\varepsilon^{\frac{1}{2}}\beta_{\varepsilon}^{\frac{d+2}{2}}\right),$
-where $\beta_{\varepsilon}$ is the embedding constant from Lemma 3.
-By Lemma 3, $\beta_{\varepsilon}=\mathcal{O}(\varepsilon^{-\frac{1}{d+2}})$.
-Therefore $\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega_{\varepsilon})}\leq C$ for $C>0$ independent of $\varepsilon$.
-Lastly, since $\|\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega^{f}_{\varepsilon})}=\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega^{f}_{\varepsilon})}\leq C$, we have the desired estimate.
+$\|\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S;L^{2}(\Omega_{\varepsilon}))\cap L^{2}(S;H^{1}(\Omega_{\varepsilon}))}\leq Ck\left(\int_{S}|B_{\varepsilon,k}^{f}(\tau)|\,\mathrm{d}\tau\right)^{\frac{1}{2}}.$
+Applying [37, Chapter 2, Theorem 6.1], we now have to be careful regarding the $\varepsilon$ dependency of the estimate. First we obtain
+$\|\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega_{\varepsilon})}\leq C\left(1+C\beta_{\varepsilon}^{\frac{d+2}{2}}|S|^{\frac{1}{2}}|\Omega_{\varepsilon}|^{\frac{1}{2}}\right)\leq C\left(1+C\varepsilon^{\frac{1}{2}}\beta_{\varepsilon}^{\frac{d+2}{2}}\right),$
+where $\beta_{\varepsilon}=\mathcal{O}(\varepsilon^{-\frac{1}{d+2}})$ is the embedding constant from Lemma 3.
+Therefore $\|\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon}^{f}\|_{L^{\infty}(S\times\Omega_{\varepsilon})}\leq C$ independent of $\varepsilon$.
+Consequently $\|\theta_{\varepsilon}^{f}\|_{L^{\infty}(S\times\Omega^{f}_{\varepsilon})}=\|\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon}^{f}\|_{L^{\infty}(S\times\Omega^{f}_{\varepsilon})}\leq C$, and $\|\theta_{\varepsilon}^{f}\|_{L^{\infty}(S\times\Gamma_{\varepsilon})}\leq C$.
+Plugging this result back into Equation 12 and choosing $k$ large enough, we obtain
+$\|\theta_{\varepsilon,k}^{s}(t)\|_{L^{2}(\Omega^{s}_{\varepsilon})}=0\quad\text{ for almost all }t\in S,$
+resulting in $\theta_{\varepsilon,k}^{s}=0$ in $L^{2}(S\times\Omega^{s}_{\varepsilon})$ and consequently $\theta^{s}_{\varepsilon}\in L^{\infty}(S\times\Omega^{s}_{\varepsilon})$.
+As mentioned at the beginning, the estimate for $f_{\varepsilon}^{s}\neq 0,f_{\varepsilon}^{f}=0$ follows by the same steps only with the roles of $\theta^{s}_{\varepsilon}$ and $\theta^{f}_{\varepsilon}$ switched, the other case is also slightly simpler since the estimates in the solid domain do not depend on $\varepsilon$.
+Adding the solutions for both cases yields the solution for the general case $f_{\varepsilon}^{s}\neq 0,f_{\varepsilon}^{f}\neq 0$ and consequently the general $L^{\infty}$ estimate.
 ∎
@@
 Let the same assumptions be fulfilled as in Lemma 5 and let $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ be the solution of Eq.
-3.1, with $\partial_{t}\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$.
+7, with $\partial_{t}\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$.
 If $\|v_{\varepsilon}\|_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\leq C\sqrt{\varepsilon}$, then there exists a $C>0$ independent of $\varepsilon$ such that
-$\begin{split}\|\partial_{t}\theta_{\varepsilon}\|_{L^{2}(S;H^{1}(\Omega^{s}_{\varepsilon})^{\prime})}\leq C\quad\text{ and }\quad\|\partial_{t}\theta_{\varepsilon}\|_{L^{2}(S;H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime})}\leq C\sqrt{\varepsilon}.\end{split}$ (3.8)
+$\begin{split}\|\partial_{t}\theta_{\varepsilon}\|_{L^{2}(S;H^{1}(\Omega^{s}_{\varepsilon})^{\prime})}\leq C\quad\text{ and }\quad\|\partial_{t}\theta_{\varepsilon}\|_{L^{2}(S;H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime})}\leq C\sqrt{\varepsilon}.\end{split}$ (14)
 Proof.
-By testing Eq. 3.1 with arbitrary $\varphi\in L^{2}(S;\Theta_{\varepsilon})$ one obtains for almost all $t\in S$
-$\begin{split}\bigl{|}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}&+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}\bigr{|}\\ &\leq|(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{s}_{\varepsilon}}|+\frac{1}{\varepsilon}|(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{f}_{\varepsilon}}|+\alpha|(\llbracket\theta_{\varepsilon}\rrbracket,\llbracket\varphi\rrbracket)_{\Gamma_{\varepsilon}}|+|(f_{\varepsilon},\varphi^{s})_{\Gamma_{\varepsilon}}|+|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}|\\ &\leq\kappa^{s}\|\nabla\theta_{\varepsilon}\|_{L^{2}(\Omega^{s}_{\varepsilon})}\|\nabla\varphi^{g}\|_{L^{2}(\Omega^{s}_{\varepsilon})}+\frac{\kappa^{f}}{\varepsilon}\|\nabla\theta_{\varepsilon}\|_{L^{2}(\Omega^{f}_{\varepsilon})}\|\nabla\varphi^{f}\|_{L^{2}(\Omega^{f}_{\varepsilon})}+\alpha\|\llbracket\theta_{\varepsilon}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}(\|\varphi^{f}\|_{L^{2}(\Gamma_{\varepsilon})}+\|\varphi^{s}\|_{L^{2}(\Gamma_{\varepsilon})})\\ &\hskip 20.0pt+C\|f_{\varepsilon}\|_{L^{2}(\Gamma_{\varepsilon})}\|\varphi^{s}\|_{H^{1}(\Omega^{s}_{\varepsilon})}+|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}|.\end{split}$
-Choosing $\varphi^{f}=0$, integrating over the time interval $S$, followed by Hölder’s inequality and applying (3.3) gives the desired estimate for $\partial_{t}\theta_{\varepsilon}^{s}$.
+By testing Eq. 7 with arbitrary $\varphi\in L^{2}(S;\Theta_{\varepsilon})$ one obtains for almost all $t\in S$
+$\begin{split}&\bigl|\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}\bigr|\\ &\leq|(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{s}_{\varepsilon}}|+\frac{1}{\varepsilon}|(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{f}_{\varepsilon}}|+\alpha|(\llbracket\theta_{\varepsilon}\rrbracket,\llbracket\varphi\rrbracket)_{\Gamma_{\varepsilon}}|+|(f_{\varepsilon}^{s},\varphi^{s})_{\Gamma_{\varepsilon}}|+|(f_{\varepsilon}^{f},\varphi^{f})_{\Gamma_{\varepsilon}}|+|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}|\\ &\leq\kappa^{s}\|\nabla\theta_{\varepsilon}\|_{L^{2}(\Omega^{s}_{\varepsilon})}\|\nabla\varphi^{g}\|_{L^{2}(\Omega^{s}_{\varepsilon})}+\frac{\kappa^{f}}{\varepsilon}\|\nabla\theta_{\varepsilon}\|_{L^{2}(\Omega^{f}_{\varepsilon})}\|\nabla\varphi^{f}\|_{L^{2}(\Omega^{f}_{\varepsilon})}+|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}|\\ &\hskip 10.0pt+\alpha\|\llbracket\theta_{\varepsilon}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}(\|\varphi^{f}\|_{L^{2}(\Gamma_{\varepsilon})}+\|\varphi^{s}\|_{L^{2}(\Gamma_{\varepsilon})})+C\|f_{\varepsilon}^{s}\|_{L^{2}(\Gamma_{\varepsilon})}\|\varphi^{s}\|_{H^{1}(\Omega^{s}_{\varepsilon})}+\|f_{\varepsilon}^{f}\|_{L^{2}(\Gamma_{\varepsilon})}\|\varphi^{f}\|_{L^{2}(\Gamma_{\varepsilon})}.\end{split}$
+Choosing $\varphi^{f}=0$, integrating over the time interval $S$, followed by Hölder’s inequality and applying (9) gives the desired estimate for $\partial_{t}\theta_{\varepsilon}^{s}$.
 For estimating $\partial_{t}\theta^{f}_{\varepsilon}$ we set $\varphi^{s}=0$ then the only problematic part is the convection term (the trace over $\Gamma_{\varepsilon}$ can be estimated by Lemma 2).
 By Eq.
-3.5 and $\operatorname{div}{v_{\varepsilon}}=0$
-$\begin{split}\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi^{f})_{S\times\Omega^{f}_{\varepsilon}}=-\frac{1}{\varepsilon}(v_{\varepsilon}\theta_{\varepsilon},\nabla\varphi^{f})_{S\times\Omega^{f}_{\varepsilon}}+\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}.\end{split}$
+11 and $\operatorname{div}{v_{\varepsilon}}=0$
+$\begin{split}\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi^{f})_{S\times\Omega^{f}_{\varepsilon}}=-\frac{1}{\varepsilon}(v_{\varepsilon}\theta_{\varepsilon},\nabla\varphi^{f})_{S\times\Omega^{f}_{\varepsilon}}+\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}.\end{split}$
 For the first term on the right-hand side, we use the results from Lemma 6 and the assumption on $v_{\varepsilon}$ to obtain
@@
 For the second term, we again use our choice of boundary conditions, leading to
-$\begin{split}\left|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}\right|\leq\frac{1}{\varepsilon}\|v_{\varepsilon}\|_{L^{\infty}(S\times\partial\Omega^{f}_{\varepsilon})}\|\theta_{\varepsilon}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\|\phi^{f}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}.\end{split}$
+$\begin{split}\left|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}\right|\leq\frac{1}{\varepsilon}\|v_{\varepsilon}\|_{L^{\infty}(S\times\partial\Omega^{f}_{\varepsilon})}\|\theta_{\varepsilon}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\|\varphi^{f}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}.\end{split}$
 With Lemma 2. $(iii)$ and the estimates of $\theta^{f}_{\varepsilon}$ from Lemma 5, we obtain
-$\begin{split}\left|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}\right|\leq\frac{C}{\varepsilon}\|\theta_{\varepsilon}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\|\varphi^{f}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\leq\frac{C}{\sqrt{\varepsilon}}\|\varphi^{f}\|_{L^{2}(S;H^{1}(\Omega^{f}_{\varepsilon}))}.\end{split}$
+$\begin{split}\left|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}\right|\leq\frac{C}{\varepsilon}\|\theta_{\varepsilon}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\|\varphi^{f}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\leq\frac{C}{\sqrt{\varepsilon}}\|\varphi^{f}\|_{L^{2}(S;H^{1}(\Omega^{f}_{\varepsilon}))}.\end{split}$
 Multiplying the complete inequality by $\varepsilon$ we get the desired result. ∎
-The additional assumption $\|v_{\varepsilon}\|_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\leq C\sqrt{\varepsilon}$ is not a new requirement to the problem, as the solution of the Stokes problem will fulfill this estimate naturally.
+The additional assumption $\|v_{\varepsilon}\|_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\leq C\sqrt{\varepsilon}$ is not a new requirement to the problem, as the solution of the Stokes problem satisfies this estimate naturally.
 Lemma 8.
@@
 Then for all $\varepsilon>0$ and $\psi_{\varepsilon}\in L^{2}(S\times\Omega^{f}_{\varepsilon})$ there exists a unique solution $(u_{\varepsilon},p_{\varepsilon})\in V_{\varepsilon}\times Q_{\varepsilon}$ of Eq.
-3.2.
+8.
 In addition, it holds
-$\frac{1}{\sqrt{\varepsilon}}\|u_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\sqrt{\varepsilon}\|\nabla u_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}\leq C\quad\text{ and }\quad\sqrt{\varepsilon}\|p_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}\leq\frac{C}{\varepsilon}$ (3.9)
+$\frac{1}{\sqrt{\varepsilon}}\|u_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\sqrt{\varepsilon}\|\nabla u_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}\leq C\quad\text{ and }\quad\sqrt{\varepsilon}\|p_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}\leq\frac{C}{\varepsilon}$ (15)
 for an $C<\infty$ independent of $\varepsilon$ and $\psi_{\varepsilon}$.
 Proof.
-Given the bounds of the viscosity (A2) and Assumption (A5), this is analog to the steps in [12, Section 3.2 and 3.4]. ∎
-We introduce the operators $\mathbb{T}_{\varepsilon}:V_{\varepsilon}\to L^{2}(S\times\Omega)$, which maps a prescribed velocity field to the corresponding temperature field via Eq.
-3.1, and the Stokes operator $\mathbb{S}_{\varepsilon}:L^{2}(S\times\Omega)\to V_{\varepsilon}$ of Eq.
-3.2, which returns the velocity field for a given temperature distribution.
+Existence and uniqueness follow from [42, Theorem 5.1], where the proof for constant viscosity is stated, which is transferable to our setup given the bounds assumed in (A2).
+Furthermore, by Assumption (A5), and the extension of the boundary function $U_{e}$, the velocity estimates follow by testing with $u_{\varepsilon}-U_{\varepsilon}$ and the Poincaré inequality, as carried out in [25, Section 3.2].
+The pressure estimate follows by combining the estimate of the velocity gradient with the results in [35, Lemma 4.7] for the Bogovskii operator in thin porous layers (which can be directly transferred to our setup).
+See also [34] for further studies of the Bogovskii operator in thin non-porous layers.
+∎
+Next, we introduce the operators $\mathbb{T}_{\varepsilon}:V_{\varepsilon}\to L^{2}(S\times\Omega)$, which maps a prescribed velocity field to the corresponding temperature field via Eq.
+7, and the Stokes operator $\mathbb{S}_{\varepsilon}:L^{2}(S\times\Omega)\to V_{\varepsilon}$ of Eq.
+8, which returns the velocity field for a given temperature distribution.
 We are looking for fixed points of the operator
-$\mathbb{F}_{\varepsilon}:L^{2}(S\times\Omega)\to L^{2}(S\times\Omega),\quad\psi_{\varepsilon}\mapsto\mathbb{T}_{\varepsilon}\left(\mathbb{S}_{\varepsilon}(\psi_{\varepsilon})\right).$ (3.10)
+$\mathbb{F}_{\varepsilon}:L^{2}(S\times\Omega)\to L^{2}(S\times\Omega),\quad\psi_{\varepsilon}\mapsto\mathbb{T}_{\varepsilon}\left(\mathbb{S}_{\varepsilon}(\psi_{\varepsilon})\right).$ (16)
 and point out that any such fixed point is automatically a weak solution to the full problem.
@@
 for a subsequence. We are now able to pass to the limit in the weak formulation and obtain
-00 0 $=\lim_{k\to\infty}\left[(\mu(\psi_{k})\nabla u_{k},\nabla v)_{S\times\Omega^{f}_{\varepsilon}}-(p_{k},\operatorname{div}{(v)})_{S\times\Omega^{f}_{\varepsilon}}\right]=(\mu(\psi)\nabla u,\nabla v)_{S\times\Omega^{f}_{\varepsilon}}-(p,\operatorname{div}{(v)})_{S\times\Omega^{f}_{\varepsilon}},$
+$0$ $=\lim_{k\to\infty}\left[(\mu(\psi_{k})\nabla u_{k},\nabla v)_{S\times\Omega^{f}_{\varepsilon}}-(p_{k},\operatorname{div}{(v)})_{S\times\Omega^{f}_{\varepsilon}}\right]=(\mu(\psi)\nabla u,\nabla v)_{S\times\Omega^{f}_{\varepsilon}}-(p,\operatorname{div}{(v)})_{S\times\Omega^{f}_{\varepsilon}},$
 for all $v\in H^{1}_{0}(\Omega^{f}_{\varepsilon})^{d}$.
-By localization in time, we see that, for almost all $t\in S$, $(u,p)$ is the solution to the Stokes problem (3.2) in $V_{\varepsilon}\times Q_{\varepsilon}$ corresponding to $\psi$.
+By localization in time, we see that, for almost all $t\in S$, $(u,p)$ is the solution to the Stokes problem (8) in $V_{\varepsilon}\times Q_{\varepsilon}$ corresponding to $\psi$.
 Because the solution is unique, the whole sequence weakly converges to the same limit.
@@
 Proof.
-Given the assumptions, the sequences $\{\theta_{k}\}_{k\in\mathbb{N}}:=\{\mathbb{T}_{\varepsilon}(u_{k})\}_{k\in\mathbb{N}}\subset L^{2}(S;\Theta_{\varepsilon})$ and $\{\partial_{t}\theta_{k}\}_{k\in\mathbb{N}}\subset L^{2}(S;\Theta_{\varepsilon}^{\prime})$ are bounded uniformly in $k\in\mathbb{N}$ (see Lemma 4).
+Given the assumptions, the sequences $\{\theta_{k}\}_{k\in\mathbb{N}}\coloneqq\{\mathbb{T}_{\varepsilon}(u_{k})\}_{k\in\mathbb{N}}\subset L^{2}(S;\Theta_{\varepsilon})$ and $\{\partial_{t}\theta_{k}\}_{k\in\mathbb{N}}\subset L^{2}(S;\Theta_{\varepsilon}^{\prime})$ are bounded uniformly in $k\in\mathbb{N}$ (see Lemma 5).
 Therefore, there is a subsequence, still indexed with $k$, and a weak limit $\theta\in L^{2}(S;\Theta_{\varepsilon})$ with $\partial_{t}\theta\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$ such that $\theta_{k}\rightharpoonup\theta$ in $L^{2}(S;\Theta_{\varepsilon})$ and $\partial_{t}\theta_{k}\rightharpoonup\partial_{t}\theta$ in $L^{2}(S;\Theta_{\varepsilon}^{\prime})$.
@@
 Looking at the weak form Eq.
-3.1, we can immediately pass to the limit in most terms with only the convection term as a potential issue.
+7, we can immediately pass to the limit in most terms with only the convection term as a potential issue.
 Now, since $\theta_{k}\to\theta$ strongly in $L^{2}(S\times\Omega_{\varepsilon}^{f})$ and $u_{k}\rightharpoonup u$ weakly in $L^{2}(S\times\Omega_{\varepsilon}^{f})^{d}$, we have $u_{k}\theta_{k}\rightharpoonup u\theta$ weakly in $L^{1}(S\times\Omega_{\varepsilon}^{f})^{d}$.
@@
 Corollary 1.
-The operator $\mathbb{F}_{\varepsilon}$ defined in Eq. 3.10 is continuous from $L^{2}(S\times\Omega)$ to $L^{2}(S\times\Omega)$.
+The operator $\mathbb{F}_{\varepsilon}$ defined in Eq. 16 is continuous from $L^{2}(S\times\Omega)$ to $L^{2}(S\times\Omega)$.
 With the continuity, we can now tackle the existence of a fixed point.
 Theorem 1 (Existence).
-Let Assumptions (A1)-(A5) be satisfied. Then the operator $\mathbb{F}_{\varepsilon}$ defined in Eq. 3.10 has a fixed point in $L^{2}(S;\Theta_{\varepsilon})$.
+Let Assumptions (A1)-(A5) be satisfied. Then, for all $\varepsilon>0$, the operator $\mathbb{F}_{\varepsilon}$ defined in Eq. 16 has a fixed point in $L^{2}(S;\Theta_{\varepsilon})$.
 Proof.
-We observe that $L^{2}(S;\Theta_{\varepsilon})\subset L^{2}(S\times\Omega)$ and $\mathbb{F}_{\varepsilon}[L^{2}(S\times\Omega)]\subset L^{2}(S;\Theta_{\varepsilon})$ by Lemma 4 and 7.
-Furthermore, $\partial_{t}\psi\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$ for all $\psi\in\mathbb{F}_{\varepsilon}[L^{2}(S\times\Omega)]$ with uniform bounds from Lemma 4 and 6 which are independent of $\psi$.
-As in the proof of Lemma 9, we can use the Aubin-Lions lemma to conclude that $\mathbb{F}_{\varepsilon}[L^{2}(S\times\Omega)]$ is compact in $L^{2}(S\times\Omega)$.
-Finally, $\mathbb{F}_{\varepsilon}$ is $L^{2}(S\times\Omega)$-continuous (Corollary 1).
-Therefore, we apply Schauder’s fixed point theorem and obtain the existence of at least one fixed point of the operator $\mathbb{F}_{\varepsilon}$ in $L^{2}(S;\Theta)$.
+By Lemma 5 and 7 there exists a $\varepsilon$-dependent $C_{\varepsilon}>0$ such that
+$\|\mathbb{F}_{\varepsilon}(\psi)\|_{L^{2}(S;H^{1}(\Omega^{s}_{\varepsilon}\cup\Omega^{f}_{\varepsilon}))}+\|\partial_{t}(\mathbb{F}_{\varepsilon}(\psi))\|_{L^{2}(S;\Theta_{\varepsilon}^{\prime})}\leq C_{\varepsilon},\quad\text{ for all }\psi\in L^{2}(S\times\Omega).$
+Define the set
+$M\coloneqq\left\{\psi\in L^{2}(S;\Theta_{\varepsilon}):\partial_{t}\psi\in L^{2}(S;\Theta_{\varepsilon}^{\prime}),\|\psi\|_{L^{2}(S;H^{1}(\Omega^{s}_{\varepsilon}\cup\Omega^{f}_{\varepsilon}))}+\|\partial_{t}\psi\|_{L^{2}(S;\Theta_{\varepsilon}^{\prime})}\leq C_{\varepsilon}\right\}.$
+By construction, $M$ is closed, bounded, convex, and nonempty.
+Furthermore, by Lemma 5 and 8 it holds $\mathbb{F}_{\varepsilon}:M\to M$.
+We know that $\mathbb{F}_{\varepsilon}$ is $L^{2}(S\times\Omega)$-continuous (Corollary 1).
+As in the proof of Lemma 10, we can use the Aubin-Lions lemma to conclude that $M$, and also $\mathbb{F}_{\varepsilon}(M)$, are compact in $L^{2}(S\times\Omega)$.
+Therefore, we apply Schauder’s fixed point theorem [47, Theorem 1.C] and obtain the existence of at least one fixed point of the operator $\mathbb{F}_{\varepsilon}$ in $M$, and subsequently in $L^{2}(S;\Theta_{\varepsilon})$.
 ∎
-The previous theorem ensures that, under the given assumptions, there exists a weak solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})$ of the problem (2.1) for all $\varepsilon>0$.
+The previous theorem ensures that, under the given assumptions, there exists a weak solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})$ of the problem (1) for all $\varepsilon>0$.
 By construction of the operator $\mathbb{F}_{\varepsilon}$ the fixed point also fulfills the estimates of Lemma 5 and Lemma 8.
@@
 If there is a solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})\times V_{\varepsilon}\times Q_{\varepsilon}$ of the coupled system Eq.
-3.1 and (3.2), such that
-  • if $d=2:$ $\nabla\theta_{\varepsilon}\in L^{\infty}(S;L^{1+\delta}(\Omega_{\varepsilon}^{f})^{2})$ and $\nabla u_{\varepsilon}\in L^{\infty}(S;L^{2+\delta}(\Omega_{\varepsilon}^{f})^{2\times 2})$,
-  • if $d=3:$ $\nabla\theta_{\varepsilon}\in L^{\infty}(S;L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f})^{3})$ and $\nabla u_{\varepsilon}\in L^{\infty}(S;L^{3+\delta}(\Omega_{\varepsilon}^{f})^{3\times 3})$,
+7 and (8), such that
+  • if $d=2:$ $\nabla\theta_{\varepsilon}\in L^{\infty}(S;L^{1+\delta}(\Omega_{\varepsilon}^{f}))^{2}$ and $\nabla u_{\varepsilon}\in L^{\infty}(S;L^{2+\delta}(\Omega_{\varepsilon}^{f}))^{2\times 2}$,
+  • if $d=3:$ $\nabla\theta_{\varepsilon}\in L^{\infty}(S;L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f}))^{3}$ and $\nabla u_{\varepsilon}\in L^{\infty}(S;L^{3+\delta}(\Omega_{\varepsilon}^{f}))^{3\times 3}$,
 for arbitrary $\delta>0$. Then, this is the only solution to the problem.
 Proof.
-We provide the proof in the case $d=3$, the other case follows by the same steps but with a better estimate in the Gagliardo-Nierenberg inequality.
+We provide the proof in the case $d=3$, the other case follows by the same steps but with a better estimate in the Gagliardo-Nirenberg inequality.
 Suppose there are two sets of solutions $(\theta_{\varepsilon}^{i},u_{\varepsilon}^{i},p_{\varepsilon}^{i})\in L^{2}(S;\Theta_{\varepsilon})\times V_{\varepsilon}\times Q_{\varepsilon}$ ($i=1,2$) with their differences denoted by $(\overline{\theta_{\varepsilon}},\overline{u_{\varepsilon}},\overline{p_{\varepsilon}})$.
@@
 Now, choosing $(\varphi,v)=(\overline{\theta_{\varepsilon}},\overline{u_{\varepsilon}})$ as test functions, we obtain
-$\frac{1}{2}\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+c_{g}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\frac{1}{2\varepsilon}\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\frac{c_{f}}{\varepsilon}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\alpha\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Gamma_{\varepsilon})}^{2}+\frac{1}{\varepsilon}(u_{\varepsilon}^{2}\cdot\nabla\overline{\theta_{\varepsilon}},\overline{\theta_{\varepsilon}})_{\Omega^{f}_{\varepsilon}}\\ \leq-\frac{1}{\varepsilon}(\overline{u_{\varepsilon}}\cdot\nabla\theta_{\varepsilon}^{1},\overline{\theta_{\varepsilon}})_{\Omega^{f}_{\varepsilon}},$
-$\begin{split}&\underline{\mu}\|\nabla\overline{u_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}\leq((\mu({\theta_{\varepsilon}}^{1})-\mu({\theta_{\varepsilon}}^{2}))\nabla u_{\varepsilon}^{1},\nabla\overline{u_{\varepsilon}})_{\Omega^{f}_{\varepsilon}}.\end{split}$
+$\frac{1}{2}\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\kappa_{g}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\frac{1}{2\varepsilon}\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\frac{\kappa_{f}}{\varepsilon}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\alpha\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Gamma_{\varepsilon})}^{2}+\frac{1}{\varepsilon}(u_{\varepsilon}^{2}\cdot\nabla\overline{\theta_{\varepsilon}},\overline{\theta_{\varepsilon}})_{\Omega^{f}_{\varepsilon}}\\ =-\frac{1}{\varepsilon}(\overline{u_{\varepsilon}}\cdot\nabla\theta_{\varepsilon}^{1},\overline{\theta_{\varepsilon}})_{\Omega^{f}_{\varepsilon}},$
+$\begin{split}&\underline{\mu}\|\nabla\overline{u_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}=((\mu({\theta_{\varepsilon}}^{1})-\mu({\theta_{\varepsilon}}^{2}))\nabla u_{\varepsilon}^{1},\nabla\overline{u_{\varepsilon}})_{\Omega^{f}_{\varepsilon}}.\end{split}$
 Take $r=\nicefrac{{2(3+\delta)}}{{1+\delta}}$, then $1=\nicefrac{{1}}{{2}}+\nicefrac{{1}}{{r}}+\nicefrac{{1}}{{3+\delta}}$.
-Using the Lipschitz property of the viscosity and Hölder’s and the Poincaré inequality, we are led to
-$\|\overline{u_{\varepsilon}}\|_{H^{1}(\Omega_{\varepsilon}^{f})}^{2}\leq C\|\overline{\theta_{\varepsilon}}\|_{L^{r}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}.$ (3.11)
+Using the Lipschitz property of the viscosity and the Hölder and the Poincaré inequalities, we are led to
+$\|\overline{u_{\varepsilon}}\|_{H^{1}(\Omega_{\varepsilon}^{f})}^{2}\leq C\|\overline{\theta_{\varepsilon}}\|_{L^{r}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}.$ (17)
 For the heat problem, we first note that
 $(u_{\varepsilon}^{2}\cdot\nabla\overline{\theta_{\varepsilon}},\overline{\theta_{\varepsilon}})_{\Omega^{f}_{\varepsilon}}\geq 0$
-via Eq. 3.5 resulting in (for $q=\nicefrac{{18+12\delta}}{{3+10\delta}}$)
-$\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\llbracket\overline{\theta_{\varepsilon}}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}^{2}\\ \leq C_{\varepsilon}\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{L^{q}(\Omega_{\varepsilon}^{f})}\|\nabla{\theta_{\varepsilon}}^{1}\|_{{L^{\nicefrac{{3}}{{2}}+\delta}}(\Omega_{\varepsilon}^{f})}.$ (3.12)
-Summing Eqs. 3.11 and 3.12 and integrating over time, we get
-$\|\overline{\theta_{\varepsilon}}(t)\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\overline{\theta_{\varepsilon}}(t)\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\int_{0}^{t}\left(\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\llbracket\overline{\theta_{\varepsilon}}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}^{2}+\|\overline{u_{\varepsilon}}\|_{H^{1}(\Omega_{\varepsilon}^{f})}^{2}\right)\,\mathrm{d}\tau\\ \leq C_{\varepsilon}\int_{0}^{t}\left(\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{L^{q}(\Omega_{\varepsilon}^{f})}\|\nabla{\theta_{\varepsilon}}^{1}\|_{{L^{\nicefrac{{3}}{{2}}+\delta}}(\Omega_{\varepsilon}^{f})}+\|\overline{\theta_{\varepsilon}}\|_{{L^{r}}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}\right)\,\mathrm{d}\tau.$
-Focusing on the second term on the right-hand side, we have
+via Eq. 11, resulting in (for $q=\nicefrac{{18+12\delta}}{{3+10\delta}}$)
+$\begin{split}\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\llbracket\overline{\theta_{\varepsilon}}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}^{2}\\ \leq C_{\varepsilon}\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{L^{q}(\Omega_{\varepsilon}^{f})}\|\nabla{\theta_{\varepsilon}}^{1}\|_{{L^{\nicefrac{{3}}{{2}}+\delta}}(\Omega_{\varepsilon}^{f})}.\end{split}$ (18)
+Summing Eqs. 17 and 18 and integrating over time, we get
+$\begin{split}\|\overline{\theta_{\varepsilon}}(t)\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\overline{\theta_{\varepsilon}}(t)\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\int_{0}^{t}\left(\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\llbracket\overline{\theta_{\varepsilon}}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}^{2}+\|\overline{u_{\varepsilon}}\|_{H^{1}(\Omega_{\varepsilon}^{f})}^{2}\right)\,\mathrm{d}\tau\\ \leq C_{\varepsilon}\int_{0}^{t}\left(\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{L^{q}(\Omega_{\varepsilon}^{f})}\|\nabla{\theta_{\varepsilon}}^{1}\|_{{L^{\nicefrac{{3}}{{2}}+\delta}}(\Omega_{\varepsilon}^{f})}+\|\overline{\theta_{\varepsilon}}\|_{{L^{r}}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}\right)\,\mathrm{d}\tau.\end{split}$ (19)
+Focusing on the second term on the right-hand side of Equation 19, we have
 $\int_{0}^{t}\|\overline{\theta_{\varepsilon}}\|_{{L^{r}}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}\,\mathrm{d}\tau\leq\|\nabla u_{\varepsilon}^{1}\|_{L^{\infty}(S;{L^{3+\delta}}(\Omega_{\varepsilon}^{f}))}^{2}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(S;{L^{r}}(\Omega_{\varepsilon}^{f}))}^{2}.$
-The Gagliardo-Nierenberg interpolation inequality with $\lambda=\nicefrac{{3}}{{3+\delta}}$, yields
+The Gagliardo-Nirenberg interpolation inequality with $\lambda=\nicefrac{{3}}{{3+\delta}}$, yields
 $\|\overline{\theta_{\varepsilon}}\|_{L^{2}(S;{L^{r}}(\Omega_{\varepsilon}^{f}))}\leq C\|\nabla\overline{\theta_{\varepsilon}}\|^{\lambda}_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\|\overline{\theta_{\varepsilon}}\|^{1-\lambda}_{L^{2}(S\times\Omega^{f}_{\varepsilon})}.$
-Using $\nabla u_{\varepsilon}^{1}\in L^{\infty}(S;L^{3}(\Omega_{\varepsilon}^{f}))^{d}$ and the scaled Young’s inequality leads to
+Using $\nabla u_{\varepsilon}^{1}\in L^{\infty}(S;L^{3+\delta}(\Omega_{\varepsilon}^{f}))^{d}$ and the scaled Young’s inequality leads to
 $\int_{0}^{t}\|\overline{\theta_{\varepsilon}}\|_{{L^{r}}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}\,\mathrm{d}\tau\leq\frac{1}{4}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(S\times\Omega_{\varepsilon}^{f})}^{2}+C_{\varepsilon}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(S\times\Omega_{\varepsilon}^{f})}^{2}.$
-For the first term on the right-hand side, we use $\nabla\theta_{\varepsilon}^{1}\in L^{\infty}(S;{L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f})})$ to estimate
+For the first term on the right-hand side of Equation 19, we use $\nabla\theta_{\varepsilon}^{1}\in L^{\infty}(S;{L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f})})$ to estimate
 $\int_{0}^{t}\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{{L^{q}}(\Omega_{\varepsilon}^{f})}\|\nabla\theta_{\varepsilon}^{1}\|_{L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f})}\,\mathrm{d}\tau\leq C_{\varepsilon}\|\overline{u_{\varepsilon}}\|_{L^{2}(S;H^{1}(\Omega_{\varepsilon}^{f}))}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(S;L^{6}(\Omega_{\varepsilon}^{f}))}$
-which (again using Young’s inequality and the Gagliardo-Nierenberg interpolation) gives
+which (again using Young’s inequality and the Gagliardo-Nirenberg interpolation) gives
 $C_{\varepsilon}\int_{0}^{t}\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{{L^{q}}(\Omega_{\varepsilon}^{f})}\|\nabla\theta_{\varepsilon}^{1}\|_{L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f})}\,\mathrm{d}\tau\leq\frac{1}{2}\|\overline{u_{\varepsilon}}\|_{L^{2}(S;H^{1}(\Omega_{\varepsilon}^{f}))}^{2}+\frac{1}{4}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(S\times\Omega_{\varepsilon}^{f})}^{2}+C_{\varepsilon}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(S\times\Omega_{\varepsilon}^{f})}^{2}.$
@@
 4 Homogenization
-To pass to the limit $\varepsilon\to 0$ we utilize the concept of two-scale convergence for thin layers first introduced in [46, Definition 4.1] and further developed in following articles, see for example [14, 29] and the references therein.
+In this section, we investigate the limit behavior for $\varepsilon\to 0$ of the microscopic solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})$ and identify the macroscopic limit problem which is solved by this limit (see Theorem 4).
+For this limit analysis, we utilize the concept of two-scale convergence for thin layers first introduced in [23, Definition 4.1] and further developed in following articles, see for example [38, 22] and the references therein.
 We start by introducing the definition, then consider the existence of limit functions, and then pass to the limit in the weak formulation.
@@
 4.1 Two scale convergence
-Definiton 1 (Two-scale convergence on thin domains).
+Definition 1 (Two-scale convergence on thin domains).
   • i)
     A sequence $u_{\varepsilon}\in L^{2}(S\times\Omega_{\varepsilon})$ is said to weakly two-scale converge to a function $u\in L^{2}(S\times\Sigma\times Y^{d})$ (notation $u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}u$) if
-    $\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{S}\int_{\Omega_{\varepsilon}}u_{\varepsilon}(t,x)\varphi\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}x\,\mathrm{d}t=\int_{S}\int_{\Sigma}\int_{Y^{d}}u(t,\tilde{x},y)\varphi(t,\tilde{x},y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$ (4.1)
-    for all $\varphi\in C(\overline{S\times\Sigma};C_{\#}(\overline{Y^{d}}))$. If in addition
+    $\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{S}\int_{\Omega_{\varepsilon}}u_{\varepsilon}(t,x)\varphi\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}x\,\mathrm{d}t=\int_{S}\int_{\Sigma}\int_{Y^{d}}u(t,\tilde{x},y)\varphi(t,\tilde{x},y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$ (20)
+    for all $\varphi\in C(\overline{S\times\Sigma};C_{\#}(Y^{d}))$. If in addition
     $\lim_{\varepsilon\to 0}\frac{1}{\sqrt{\varepsilon}}\|u_{\varepsilon}\|_{L^{2}(S\times\Omega_{\varepsilon})}=\|u\|_{L^{2}(S\times\Sigma\times Y^{d})}$
@@
     A sequence $u_{\varepsilon}\in L^{2}(S\times\Gamma_{\varepsilon})$ is said to weakly two-scale converge to a function $u\in L^{2}(S\times\Sigma\times\Gamma)$ if
-    $\lim_{\varepsilon\to 0}\int_{S}\int_{\Gamma_{\varepsilon}}u_{\varepsilon}(t,x)\varphi\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}\sigma_{x}\,\mathrm{d}t=\int_{S}\int_{\Sigma}\int_{\Gamma}u(t,\tilde{x},y)\varphi(t,\tilde{x},y)\,\mathrm{d}\sigma_{y}\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$ (4.2)
+    $\lim_{\varepsilon\to 0}\int_{S}\int_{\Gamma_{\varepsilon}}u_{\varepsilon}(t,x)\varphi\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}\sigma_{x}\,\mathrm{d}t=\int_{S}\int_{\Sigma}\int_{\Gamma}u(t,\tilde{x},y)\varphi(t,\tilde{x},y)\,\mathrm{d}\sigma_{y}\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$ (21)
     for all $\varphi\in C(\overline{S\times\Sigma};C_{\#}(\Gamma))$.
-Note, that due to the Assumption (A5) it holds for the boundary condition of the velocity
-$\frac{1}{\sqrt{\varepsilon}}\|u_{\varepsilon,BC}\|_{L^{2}(\partial\Sigma\times(0,\varepsilon))}=\|u_{BC}\|_{L^{2}(\partial\Sigma\times(0,1))},$
-as well as
-$\begin{split}\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{\partial\Sigma\times(0,\varepsilon)}u_{\varepsilon,BC}\varphi\left(\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}x&=\lim_{\varepsilon\to 0}\int_{\partial\Sigma}\int_{0}^{1}u_{BC}(\tilde{x},x_{d})\varphi\left(\tilde{x},\nicefrac{{\tilde{x}}}{{\varepsilon}},x_{d}\right)\,\mathrm{d}x_{d}\,\mathrm{d}\tilde{x}\\ &=\int_{\partial\Sigma}\int_{Y^{d}}u_{BC}(\tilde{x},y_{d})\varphi(\tilde{x},y)\,\mathrm{d}y\,\mathrm{d}\tilde{x},\end{split}$ (4.3)
-for all $\varphi\in C(\partial\Sigma;C_{\#}(\overline{Y^{d}}))$.
-The two above properties imply the strong two-scale convergence of the boundary function, $u_{\varepsilon,BC}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}u_{BC}$.
-Where (given that the underlying domain is rectangular with integer corners) the strong two-scale convergence on the lateral boundary is defined as Definition 1 $i)$, with $\Omega_{\varepsilon}$ and $\Sigma$ replaced by $\partial\Sigma\times(0,\varepsilon)$ and $\partial\Sigma$ respectively.
+We remark that Definition 1 is given for the thin layer $\Omega_{\varepsilon}=\Sigma\times(0,\varepsilon)$, since there the limit behavior is more straightforward to define.
+However, the definition can, with the extension operator from Lemma 1, be directly transferred to the rough thin layer $\Omega^{f}_{\varepsilon}$.
+A function $u_{\varepsilon}\in L^{2}(S\times\Omega^{f}_{\varepsilon})$ is said to two-scale converge to a function $u\in L^{2}(S\times\Sigma\times Y^{d})$ if
+$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{S}\int_{\Omega^{f}_{\varepsilon}}u_{\varepsilon}(t,x)\varphi\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}x\,\mathrm{d}t=\int_{S}\int_{\Sigma}\int_{Z}u(t,\tilde{x},y)\varphi(t,\tilde{x},y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$
+for all admissible test functions $\varphi$. In this case we write $u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}u$.
 Given the assumptions on the data and the estimate derived in Section 3.2 we obtain the following limit functions.
 Proposition 3 (Existence of limit functions).
-Under the Assumption (A1)-(A5), it holds for the weak solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})\times V_{\varepsilon}\times Q_{\varepsilon}$ of Eq.
-2.1 that there exists a subsequence still denoted by $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})$ and:
+Under Assumptions (A1)-(A5), weak solutions $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})\times V_{\varepsilon}\times Q_{\varepsilon}$ of Eq.
+1 have convergent subsequences.
+More specifically, there are limit functions (without relabeling for the subsequences)
   • i)
-    a $\theta^{s}\in L^{2}(S;H^{1}(\Omega))\cap H^{1}(S;H^{1}(\Omega)^{\prime})$ with
-    $\chi_{\Omega^{s}_{\varepsilon}}\theta_{\varepsilon}\to\theta^{s}\text{ in }L^{2}(S\times\Omega)\quad\text{ and }\quad\chi_{\Omega^{s}_{\varepsilon}}\nabla\theta_{\varepsilon}\rightharpoonup\nabla\theta^{s}\text{ in }L^{2}(S\times\Omega)^{d},$
+    $\theta^{s}\in L^{2}(S;H^{1}(\Omega))\cap H^{1}(S;H^{1}(\Omega)^{\prime})$ such that
+    $\theta_{\varepsilon}^{s}\to\theta^{s}\text{ in }L^{2}(S\times\Omega),\quad\partial_{t}\theta_{\varepsilon}^{s}\rightharpoonup\partial_{t}\theta^{s}\text{ in }L^{2}(S;H^{1}(\Omega)^{\prime}),\quad\nabla\theta_{\varepsilon}^{s}\rightharpoonup\nabla\theta^{s}\text{ in }L^{2}(S\times\Omega)^{d},$
   • ii)
-    a $\theta^{f}\in L^{2}(S;H^{1}(\Sigma))\cap H^{1}(S;H^{1}(\Sigma)^{\prime})$ and a $\theta_{1}^{f}\in L^{2}(S\times\Sigma;H^{1}_{\#}(Y^{d}))$ with
-    $\chi_{\Omega^{f}_{\varepsilon}}\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightarrow}}\chi_{Z}\theta^{f}\quad\text{ and }\quad\chi_{\Omega^{f}_{\varepsilon}}\nabla\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\left(\nabla_{\tilde{x}}\theta^{f}+\nabla_{y}\theta_{1}^{f}\right),$
+    $\theta^{f}\in L^{2}(S;H^{1}(\Sigma))\cap H^{1}(S;H^{1}(\Sigma)^{\prime})$ and $\theta_{1}^{f}\in L^{2}(S\times\Sigma;H^{1}_{\#}(Y^{d}))$ such that
+    $\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightarrow}}\chi_{Z}\theta^{f}\quad\text{ and }\quad\nabla\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\left(\nabla_{\tilde{x}}\theta^{f}+\nabla_{y}\theta_{1}^{f}\right).$
+    Further, it holds
+    $\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{S}\langle\partial_{t}\theta_{\varepsilon}^{f},\psi+\varphi_{\varepsilon}\rangle_{H^{1}_{\#}(\Omega^{f}_{\varepsilon})}\,\mathrm{d}t=\int_{S}\int_{Z}\langle\partial_{t}\theta^{f},\psi+\varphi\rangle_{H^{1}_{\#}(\Sigma)}\,\mathrm{d}y\,\mathrm{d}t$
+    for all $\eta\in L^{2}(S;H^{1}_{\#}(\Sigma))$ and $\varphi_{\varepsilon}(t,x)=\varphi(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}})$ with $\varphi\in C_{0}^{\infty}(S;C_{\#}^{\infty}(\Sigma;C_{\#}(Z))).$
   • iii)
-    a $u\in L^{2}(S\times\Sigma;H^{1}_{\#}(\overline{Z}))^{d}$ with
-    $\chi_{\Omega^{f}_{\varepsilon}}u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}u\quad\text{ and }\quad\chi_{\Omega^{f}_{\varepsilon}}\varepsilon\nabla u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\nabla_{y}u.$
-    It holds $\operatorname{div}_{y}{(u(\tilde{x},y))}=0$ and $\operatorname{div}_{\tilde{x}}{\left(\int_{Z}u(\tilde{x},y)\,\mathrm{d}y\right)}=0$ for almost all $(\tilde{x},y)\in\Sigma\times Z$.
-Moreover, $u$ has, for almost all $(t,x)\in S\times\Sigma$, the boundary values $u(t,\tilde{x},\tilde{y},0)=u_{\text{motion}}$ for $\tilde{y}\in Y^{d-1}$ and $u(t,\tilde{x},y)=0$ for $y\in\Gamma$ almost everywhere.
+    $u\in L^{2}(S\times\Sigma;H^{1}_{\#}(Y^{d}))^{d}$ with
+    $u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}u\quad\text{ and }\quad\varepsilon\nabla u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\nabla_{y}u.$
+    It holds $\operatorname{div}_{y}{u(t,\tilde{x},y)}=0$ and $\operatorname{div}_{\tilde{x}}{\left(\int_{Z}u(t,\tilde{x},y)\,\mathrm{d}y\right)}=0$ for almost all $(t,\tilde{x},y)\in S\times\Sigma\times Z$.
+Moreover, $u=u_{\text{motion}}$ on $S\times\Sigma\times\{y_{d}=0\}$ and $u=0$ for $S\times\Sigma\times\Gamma$ almost everywhere.
   • iv)
-    a $p\in L^{2}_{0}(S\times\Sigma\times Z)$ that is periodic in $Y^{d-1}$ and
+    $p\in L^{2}(S\times\Sigma;L^{2}_{0,\#}(Z))$ with
     $\varepsilon^{2}\chi_{\Omega^{f}_{\varepsilon}}p_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}p.$
 Additionally, on the rough interface, it holds
-$\lim_{\varepsilon\to 0}\int_{S}\int_{\Gamma_{\varepsilon}}\alpha(\theta^{s}_{\varepsilon}-\theta^{f}_{\varepsilon})\varphi\,\mathrm{d}\sigma_{x}\,\mathrm{d}t=\int_{S}\int_{\Sigma}\alpha|\Gamma|(\theta^{s}-\theta^{f})\varphi\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$
-for all admissible test functions $\varphi$.
+$\lim_{\varepsilon\to 0}\int_{S}\int_{\Gamma_{\varepsilon}}\alpha(\theta^{s}_{\varepsilon}-\theta^{f}_{\varepsilon})\varphi(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}})\,\mathrm{d}\sigma_{x}\,\mathrm{d}t=\int_{S}\int_{\Sigma}\alpha|\Gamma|(\theta^{s}-\theta^{f})\varphi(t,\tilde{x},y)\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$ (22)
+for all $\varphi\in C(\overline{S\times\Sigma};C_{\#}(\Gamma))$.
 Proof.
-Given the estimates of Lemma 4 the convergence in $i)$ can be deduced with the same arguments as in [19, Theorem 4].
-The aspects $ii)$-$iv)$ are obtained by the estimates of Lemma 4 and Lemma 7 and usual two-scale properties, see for example [14, 25, 29].
-For the convergence of the time derivative $\partial_{t}\theta^{f}_{\varepsilon}$ as well as the strong two-scale convergence of $\theta^{f}_{\varepsilon}$, we point to the results of [27, Proposition 5], since we have the same $\varepsilon$-dependent solution estimates.
-There are already several results for fluid and pressure limits in thin layers, see for example [22, 26].
-The general idea for the divergence properties is similar to the case of porous domains in non-thin domains [4, Proposition 1.14].
-The last integral identity is also a common equality, see for example [46] for similar setups and the general proof idea.
-∎
+$i).$ $\theta_{\varepsilon}^{s}$ is uniformly bounded in $L^{2}(S;H^{1}(\Omega^{s}_{\varepsilon}))\cap H^{1}(S;H^{1}(\Omega^{s}_{\varepsilon})^{\prime})$ as a result of lemmas 5 and 7.
+From here, these convergence results follow via [18, Theorem 4].
+ii).ii).
+These convergence results are due to lemmas 5 and 7 together with the properties of two-scale convergence.
+For $\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\theta^{f}$ and $\nabla\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\left(\nabla_{\tilde{x}}\theta^{f}+\nabla_{y}\theta_{1}^{f}\right)$, we refer to [22, Proposition 4.3]; the same arguments can also be found in [38, 48].
+For the convergence of the time derivative $\partial_{t}\theta^{f}_{\varepsilon}$ as well as the strong two-scale convergence of $\theta^{f}_{\varepsilon}$, we point to [36, Proposition 5].
+$iii)$-$iv.)$ There are already several very similar results for fluid and pressure limits in thin layers available in the literature, see for example [35, 49].
+Here, we rely on the estimates given via Lemma 8.
+The specific convergence results in a stationary setting can be found in [35, Lemma 4.8] and can be easily transferred to our quasi-stationary setup due to the uniform $L^{2}$-bounds.
+The general proof idea for the divergence properties is the same as for porous media in non-thin domains [50, Proposition 1.14].
+Finally, the last integral identity over interfaces (see Eq.
+22) is also a common limit in similar setups, e.g., [38, 23, 36].
+We observe that $\theta_{\varepsilon}^{f}$ and $\theta_{\varepsilon}^{s}$ are both uniformly bounded in $L^{2}(S\times\Gamma_{\varepsilon})$ via Lemma 5 together with Lemma 2.
+As a consequence, there are converging subsequences for which the limit Eq.
+22 holds ([38, Theorem 4.4.(ii)(ii)] and [36, Lemma B.1, Corollary B.2]).
 Given the strong two-scale convergence and the properties of $\mu$ we obtain the following convergence of the viscosity function.
 Corollary 2.
-Given the Assumptions (A2) it holds $\chi_{\Omega^{f}_{\varepsilon}}\mu(\theta^{f}_{\varepsilon})\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\mu(\theta^{f})$ strongly in $L^{2}(S\times\Omega^{f}_{\varepsilon})$.
-4.2 Homogenized equations
-Given the previous results concerning the limit of the solution, we can now carry out the homogenization process in the weak formulation.
-The general procedure when determining the effective model consists of three steps:
-  1. 1.
-    Use smooth test functions, compatible with the two-scale concept, to pass to the limit $\varepsilon\to 0$.
-  2. 2.
-    Try to average and decouple the effective systems to simplify the equations and extract cell problems.
-  3. 3.
-    Apply a density argument to show that the limit also holds for test functions from more general spaces.
-Most of the steps are standard for our case, therefore we do not explain everything in detail but only comment on the general behavior and explain the important terms more in-depth.
-The complete homogenized model is written down at the end of this section in Theorem 3.
-Limit procedure for the heat system.
-We start with temperature limit inside of $\Omega^{s}_{\varepsilon}$.
-Given the Assumptions (A6) and (A7) for the convergence of the initial condition and heat source, passing to the limit in the weak formulation can be carried out directly.
-To be exact, we do not require two-scale convergence inside the bulk domain.
-We point to [19, Section 5] for a similar setup with a rough interface.
-Most of the equations for the temperature limit in $\Omega^{f}_{\varepsilon}$ are also quite standard and we point to [14, 46, 29] for similar setups.
-As usual for these kinds of diffusion problems, one obtains that the limit $\theta^{f}_{1}$ can be represented by
-$\theta^{f}_{1}(t,\tilde{x},y)=\sum_{i=1}^{d-1}\omega_{i}(y)\partial_{x_{i}}\theta^{f}(t,\tilde{x})$
-where $\omega_{i}\in H^{1}_{\#}(Z)$ is the weak solution of
-$\begin{split}\operatorname{div}_{y}{(\nabla_{y}\omega_{i}+e_{i})}&=0\quad\text{ in }Z,\\ (\nabla_{y}\omega_{i}+e_{i})\cdot n&=0\quad\text{ on }\Gamma\cup\{y\in\partial Z:y_{d}=0\}.\end{split}$ (4.4)
-Next, we have a closer look at the boundary values of the limit $\theta^{f}$.
-Take $\varphi\in C^{\infty}(\overline{S\times\Sigma})$ and $\psi\in C_{\#}^{\infty}(\overline{Z})^{d}$, such that $\operatorname{div}_{y}{\psi}=0$, $\psi(y)=0$, if $y\in\Gamma$ or $y_{d}=0$ and $\varphi_{|\Sigma^{\text{out}}}=0$.
-Testing with $\varphi(t,\tilde{x})\psi(\nicefrac{{x}}{{\varepsilon}})$ yields
-$\frac{1}{\varepsilon}\int_{S}\int_{\Omega^{f}_{\varepsilon}}\theta^{f}_{\varepsilon}(t,x)\operatorname{div}_{x}{\left(\varphi(t,\tilde{x})\psi(\nicefrac{{x}}{{\varepsilon}})\right)}\,\mathrm{d}x\,\mathrm{d}t=-\frac{1}{\varepsilon}\int_{S}\int_{\Omega^{f}_{\varepsilon}}\nabla_{x}\theta^{f}_{\varepsilon}(t,x)\cdot\varphi(t,\tilde{x})\psi(\nicefrac{{x}}{{\varepsilon}})\,\mathrm{d}x\,\mathrm{d}t.$
-Passing to the limit on both sides and afterward integrating by parts on the right side, we obtain
-$\int_{S}\int_{\Sigma}\int_{Z}\theta^{f}(t,\tilde{x})\operatorname{div}_{\tilde{x}}{(\varphi(t,\tilde{x})\psi(y))}\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t=-\int_{S}\int_{\Sigma}\int_{Z}\left(\nabla_{\tilde{x}}\theta^{f}(t,\tilde{x})+\nabla_{y}\theta^{f}_{1}(t,\tilde{x},y)\right)\cdot\varphi(t,\tilde{x})\psi(y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t$
-$=\int_{S}\int_{\Sigma}\int_{Z}\theta^{f}(t,\tilde{x})\operatorname{div}_{\tilde{x}}{(\varphi(t,\tilde{x})\psi(y))}\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t-\int_{S}\int_{\Sigma^{\text{in}}}\theta^{f}(t,\tilde{x})\varphi(t,\tilde{x})\int_{Z}\psi(y)\,\mathrm{d}y\cdot n_{\Sigma}\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t$
-$+\int_{S}\int_{\Sigma}\int_{Z}\theta^{f}_{1}(t,\tilde{x},y)\varphi(t,\tilde{x})\operatorname{div}_{y}{(\psi(y))}\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t-\int_{S}\int_{\Sigma}\int_{\partial Z}\theta^{f}_{1}(t,\tilde{x},y)\varphi(t,\tilde{x})\psi(y)\cdot n_{Z}\,\mathrm{d}y\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t.$
-Given the properties of $\psi$ and periodicity of $\theta^{f}_{1}$ the last two integrals are 0. So we obtain
-$\int_{S}\int_{\Sigma^{\text{in}}}\theta^{f}(t,\tilde{x})\varphi(t,\tilde{x})\int_{Z}\psi(y)\,\mathrm{d}y\cdot n_{\Sigma}\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t=0.$
-Since the above quantity has to hold for all $\psi$, we conclude that already
-$\int_{S}\int_{\Sigma^{\text{in}}}\theta^{f}(t,\tilde{x})\varphi(t,\tilde{x})\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t=0.$
-By a density argument we can conclude that this property has to hold for all $\varphi\in L^{2}(S;H^{1}_{\Sigma^{\text{out}}}(\Sigma))$, which yields that $\theta^{f}=0$ almost everywhere on $S\times\Sigma^{\text{in}}$.
-Note, that similarly the boundary values for $u$ on $\partial Z$ follow.
-The other interesting part is the convection term.
+Given the Assumption (A2) it holds $\chi_{\Omega^{f}_{\varepsilon}}\mu(\theta^{f}_{\varepsilon})\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\mu(\theta^{f})$ strongly in $L^{2}(S\times\Omega^{f}_{\varepsilon})$.
+4.2 Homogenization limit
+We now carry out the homogenization process in the weak formulation.
+Some of the steps in this limit process are standard in this case; therefore we do not explain everything in detail but only comment on the general behavior and explain the important terms more in-depth.
+Theorem 4 (Homogenized model).
+Let the Assumptions (A1)-(A7) be fulfilled and let $(\theta^{s},\theta^{f},u,p)$ be the limit functions given by Proposition 3.
+Set $\bar{u}(t,\tilde{x})=\int_{Z}u(t,\tilde{x},y)\,\mathrm{d}y$.
+Then $(\theta^{s},\theta^{f},\bar{u},p)$ are weak solutions of the following problem:
+$\partial_{t}\theta^{s}-\kappa^{s}\Delta\theta^{s}$ $=0\quad\quad$ $\text{ in }S\times\Omega,$ (23a)
+$\theta^{s}(0,\cdot)$ $=\theta^{s}_{0}$ $\text{ in }\Omega,$ (23b)
+$-\kappa^{s}\nabla\theta^{s}\cdot\nu$ $=0$ $\text{ on }S\times\partial\Omega\setminus\Sigma,$ (23c)
+$-\kappa^{s}\nabla\theta^{s}\cdot\nu$ $=\alpha|\Gamma|\left(\theta^{f}-\theta^{s}\right)+\bar{f}^{s}\quad$ $\text{ on }S\times\Sigma,$ (23d)
+where $\bar{f}^{f,s}(t,\tilde{x})\coloneqq\int_{\Gamma}f^{f,s}(t,\tilde{x},y)\,\mathrm{d}\sigma_{y}$. The fluid temperature $\theta^{f}$ satisfies
+$|Z|\partial_{t}\theta^{f}-\operatorname{div}_{\tilde{x}}{\left(\tilde{\kappa}\nabla_{\tilde{x}}\theta^{f}+\bar{u}\theta^{f}\right)}$ $=\alpha|\Gamma|\left(\theta^{f}-\theta^{s}\right)+\bar{f}^{f}\quad\quad$ $\text{ in }S\times\Sigma,$ (23e)
+$\theta^{f}(0,\cdot)$ $=\theta^{f}_{0}$ $\text{ in }\Sigma,$ (23f)
+$-\tilde{\kappa}\nabla_{\tilde{x}}\theta^{f}\cdot n_{\Sigma}$ $=0$ $\text{ on }S\times\Sigma^{\text{out}},$ (23g)
+$\theta^{f}$ $=0\quad$ $\text{ on }S\times\Sigma^{\text{in}}.$ (23h)
+The fluid movement in the layer is governed by
+$\mu(\theta^{f})\bar{u}$ $=-K\nabla_{\tilde{x}}p+\mu(\theta^{f})\bar{\xi}_{0}\quad$ $\text{ in }S\times\Sigma,$ (23i)
+$\operatorname{div}_{\tilde{x}}(\bar{u})$ $=0\quad$ $\text{ in }S\times\Sigma,$ (23j)
+$\int_{\Sigma}p\,\mathrm{d}\tilde{x}$ $=0\quad$ $\text{ in }S,$ (23k)
+$\bar{u}(t,\tilde{x})\cdot n_{\Sigma}$ $=\int_{0}^{1}u_{BC}(\tilde{x},y)\,\mathrm{d}y\cdot n_{\Sigma}\quad$ $\text{ on }S\times\partial\Sigma.$ (23l)
+The constant effective conductivity $\tilde{\kappa}\in\mathbb{R}^{(d-1)\times(d-1)}$, effective permeability $K\in\mathbb{R}^{(d-1)\times(d-1)}$ and induced flow $\bar{\xi}_{0}\in\mathbb{R}^{d-1}$ are given as
+$\tilde{\kappa}_{ij}=\int_{Z}\kappa^{f}(\delta_{ij}+\nabla_{y}\omega_{i}\cdot e_{j})\,\mathrm{d}y,\quad K_{ij}=\int_{Z}\xi_{j}\cdot e_{i}\,\mathrm{d}y,\quad(\bar{\xi}_{0})_{i}=\int_{Z}(\xi_{0})_{i}(y)\,\mathrm{d}y,$
+where the cell solutions $w_{i}\in H^{1}_{\#}(Z)$, $\xi_{i}\in H^{1}_{\#}(Z)^{d}$, and $\xi_{0}\in H^{1}_{\#}(Z)^{d}$ (i=1,,d1i=1,...,d-1) are the unique solutions to the local cell problems given by Eqs.
+28, 4.2.2 and 4.2.2.
+Proof.
+The existence of limit functions $(\theta^{s},\theta^{f},\bar{u},p)$ is guaranteed by Proposition 3.
+The precise structure of the limit model Eq.
+23 is derived with the concept of two-scale convergence as outlined in Sections 4.2.1 and 4.2.2 below.
+∎
+Remark 2.
+Note, that by Proposition 6 and the equality Eq.
+23i we obtain that $\bar{u}\cdot e_{d}=0$.
+This is reasonable since we remain with an averaged interface velocity in the effective model and a nonzero flow in direction $d$ would imply a flow into (or out of) the solid domain.
+We recovered, ignoring the temperature-dependent viscosity, the same effective equation as in [25], where the needed assumptions for convergence are reduced, thanks to the two-scale concept.
+Additionally, in 2D the effective problem for the fluid simplifies further.
+By Eq.
+23j it holds $\operatorname{div}_{\tilde{x}}\bar{u}=\partial_{x_{1}}\bar{u}_{1}=0$ and the velocity is a constant solely given by the boundary function $u_{BC}$ in Eq.
+23l.
+This linearizes the problem since the temperature equation is independent of the pressure pp.
+The existence of a solution of model (23) is ensured by the existence of the original problem and the two-scale limits.
+Similar to the micro model (see Theorem 2), we need additional regularity assumptions to ensure uniqueness.
+These are, however, slightly lower as we have better embeddings to work with due to the dimension reduction in the fluid layer Please note that, in this case, all convergences in Proposition 3 hold for the full sequences and not just for subsequences.
+Theorem 5 (Uniqueness of the homogenized model).
+If there is a weak solution of the homogenized problem where $\nabla_{\tilde{x}}\theta^{f}\in L^{\infty}(S;L^{2+\delta}(\Sigma))$ and $\nabla_{\tilde{x}}p\in L^{\infty}(S;L^{2+\delta}(\Sigma))$ for some $\delta>0$, then this solution is the only one.
+Proof.
+Assume that $(\theta^{f}_{i},\theta^{s}_{i},\bar{u}_{i},p_{i})$, $i=1,2$, are two sets of weak solutions of the homogenization limit given by System (23) (the variational formulations are also stated in Eqs.
+33, 39 and 40).
+In the following, we denote their differences by $(\bar{\theta}^{f},\bar{\theta}^{s},\bar{\bar{u}},\bar{p})$.
+W.l.o.g.
+we assume that $(\theta_{2}^{f},p_{2})$ is the solution enjoying the higher gradient integrability.
+From Eq. 39 we obtain for $\bar{\bar{u}}$, that
+$\|\bar{\bar{u}}\|_{L^{2}(S\times\Sigma)}^{2}\leq C\left|\left(\nabla_{\tilde{x}}\bar{p},\bar{\bar{u}}\right)_{S\times\Sigma}\right|+\left|\left(\left[\frac{1}{\mu(\theta^{f}_{1})}-\frac{1}{\mu(\theta^{f}_{2})}\right]\nabla_{\tilde{x}}p_{2},\bar{\bar{u}}\right)_{S\times\Sigma}\right|$
+which, using the Lipschitz continuity of $\mu$ as well as the higher regularity of $p$, yields
+$\|\bar{\bar{u}}\|_{L^{2}(S\times\Sigma)}\leq C\left(\|\nabla_{\tilde{x}}\bar{p}\|_{L^{2}(S\times\Sigma)}+\|\nabla_{\tilde{x}}p_{2}\|_{L^{\infty}(S;L^{2+\delta}(\Omega))}\|\bar{\theta}^{f}\|_{L^{2}(S;L^{2+\frac{4}{\delta}}(\Sigma))}\right).$
+Testing Eq. 40 with $\psi=\bar{p}$, we get:
+$c\|\nabla_{\tilde{x}}\bar{p}\|_{L^{2}(S\times\Sigma)}\leq C\|\nabla_{\tilde{x}}p_{2}\|_{L^{\infty}(S;L^{2+\delta}(\Omega))}\|\bar{\theta}^{f}\|_{L^{2}(S;L^{2+\frac{4}{\delta}}(\Sigma))}$
+As a result, the uniqueness of velocity and pressure follows from the uniqueness of the fluid temperatures via
+$\|\bar{\bar{u}}\|_{L^{2}(S\times\Sigma)}+\|\nabla_{\tilde{x}}\bar{p}\|_{L^{2}(S\times\Sigma)}\leq C\|\bar{\theta}^{f}\|_{L^{2}(S;L^{2+\frac{4}{\delta}}(\Sigma))}.$ (24)
+We now look at the weak forms for the heat system (Eq. 33), take their difference and test with $(\bar{\theta}^{s},\bar{\theta}^{f})\in L^{2}(S;H^{1}(\Omega)\times H^{1}(\Sigma))$ to get
+$(\partial_{t}\bar{\theta}^{s},\bar{\theta}^{s})_{S\times\Omega}+|Z|(\partial_{t}\bar{\theta}^{f},\bar{\theta}^{f})_{S\times\Sigma}+\kappa^{s}(\nabla\bar{\theta}^{s},\nabla\bar{\theta}^{s})_{S\times\Omega}+(\tilde{\kappa}\nabla_{\tilde{x}}\bar{\theta}^{f},\nabla\bar{\theta}^{f})_{S\times\Sigma}\\ +(\bar{u}_{1}\cdot\nabla_{\tilde{x}}\bar{\theta}^{f},\bar{\theta}^{f})_{S\times\Sigma}+\alpha|\Gamma|(\bar{\theta}^{s}-\bar{\theta}^{f},\bar{\theta}^{s}-\bar{\theta}^{f})_{S\times\Sigma}=-(\bar{\bar{u}}\cdot\nabla_{\tilde{x}}\theta^{f}_{2},\bar{\theta}^{f})_{S\times\Sigma}$
+for all $(\varphi,\psi)\in L^{2}(S;H^{1}(\Omega)\times H^{1}(\Sigma))$. For the advection term on the left-hand side, we exploit that $\bar{u}_{1}$ is divergence-free (cf. Eq. 11):
+$(\bar{u}_{1}\cdot\nabla_{\tilde{x}}\bar{\theta}^{f},\bar{\theta}^{f})_{S\times\Sigma}=\frac{1}{2}(u_{BC}\cdot n,|\bar{\theta}^{f}|^{2})_{S\times\Sigma}\geq 0.$
+The advective term on the right-hand side can be handled via the additional regularity of $\nabla_{\tilde{x}}\theta^{f}_{2}$; in particular we have
+$|(\bar{\bar{u}}\cdot\nabla_{\tilde{x}}\theta^{f}_{2},\bar{\theta}^{f})_{S\times\Sigma}|\leq\|\bar{u}\|_{L^{2}(S\times\Sigma)}\|\bar{\theta}^{f}\|_{L^{2}(S;L^{2+\frac{4}{\delta}}(\Sigma))}\|\nabla\theta_{2}^{f}\|_{L^{2}(S;L^{2+\delta}(\Sigma))}.$
+Please note that $\Sigma$ is at most two-dimensional.
+For that reason, the Sobolev embedding $H^{1}\subset L^{2+\frac{4}{\delta}}$ holds for all $\delta>0$.
+With the same line of argument as in Theorem 2, uniqueness now follows.
+∎
+4.2.1 Limit procedure for the heat system.
+The $\varepsilon$-solutions $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})\cap H^{1}(S;\Theta_{\varepsilon}^{\prime})$ satisfy
+$\langle\partial_{t}\theta_{\varepsilon},\phi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+(\kappa\nabla\theta_{\varepsilon},\nabla\phi)_{\Omega^{s}_{\varepsilon}}+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\phi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}(\kappa\nabla\theta_{\varepsilon},\nabla\phi)_{\Omega^{f}_{\varepsilon}}\\ +\frac{1}{\varepsilon}(u_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\phi)_{\Omega^{f}_{\varepsilon}}+\alpha(\llbracket\theta_{\varepsilon}\rrbracket,\llbracket\phi\rrbracket)_{\Gamma_{\varepsilon}}=(f_{\varepsilon}^{s},\phi^{s})_{\Gamma_{\varepsilon}}+(f_{\varepsilon}^{f},\phi^{f})_{\Gamma_{\varepsilon}}$ (25)
+for all test functions $\phi=(\phi^{s},\phi^{f})\in L^{2}(S;\Theta_{\varepsilon})$. As a reminder, $\Theta_{\varepsilon}=H^{1}(\Omega^{s}_{\varepsilon})\times H^{1}(\Omega^{f}_{\varepsilon})$.
+We start with the limit inside $\Omega^{s}_{\varepsilon}$ by choosing test functions $\phi=(\phi^{s},0)$.
+Given the Assumptions (A6) and (A7) for the convergence of the initial condition and the heat source as well as the convergence results for $\theta_{\varepsilon}^{s}$ (LABEL:{pro:existence_limit}.(i)(i)), passing to the limit in the weak formulation can be carried out directly without relying on two-scale convergence.
+We point to [18, Section 5] for a similar setup with a rough interface.
+In the limit, we get
+$\langle\partial_{t}\theta^{s},\phi^{s}\rangle_{H^{1}(\Omega)^{\prime}}+(\kappa^{s}\nabla\theta^{s},\nabla\phi^{s})_{\Omega}+\alpha|\Gamma|(\theta^{s}-\theta^{f},\phi^{s})_{\Sigma}=(\bar{f}^{s},\phi^{s})_{\Sigma}$ (26)
+which holds for all test functions $\phi^{s}\in H^{1}(\Omega)$ and almost all $t\in S$.
+Most aspects of the temperature limit in $\Omega^{f}_{\varepsilon}$ are also quite standard and we point to [38, 23, 22] for similar setups.
+As usual for such diffusion problems, the limit $\theta^{f}_{1}$ can be represented as
+[…]
+where $\omega_{i}\in H^{1}_{\#}(Z)$, i=1,..,d1i=1,..,d-1, are weak solutions (unique up to a constant) of
+$\begin{split}-\Delta w_{i}&=0\quad\text{ in }Z,\\ (\nabla_{y}\omega_{i}+e_{i})\cdot n&=0\quad\text{ on }\Gamma_{d},\\ w_{i}(\cdot+e_{i})&=w(\cdot)\quad\text{in $Z$, $i=1,...,d-1$.}\end{split}$ (28)
+Now, take $\varphi_{\varepsilon}=\varphi_{0}(t,\tilde{x})+\varepsilon\varphi_{1}\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)$ with $\varphi_{0}\in C^{\infty}(\overline{S\times\Sigma})$ and $\varphi_{1}\in C^{\infty}(\overline{S\times\Sigma},C_{\#}^{\infty}(Z))$ such that $\varphi_{0}(t,\tilde{x})=0$ and $\varphi_{1}(t,\tilde{x},y)=0$ for $\tilde{x}\in\Sigma^{\text{in}}$.
+When testing Eq.
+25 with $\phi=(0,\varphi_{\varepsilon})$ and passing to the limit $\varepsilon\to 0$, the only problematic term is the convection term.
 Here, Assumption (A5) as well as the strong two-scale convergence of $\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightarrow}}\theta^{f}$ are used to pass to the limit.
-Take $\varphi_{\varepsilon}=\varphi_{0}(t,\tilde{x})+\varepsilon\varphi_{1}\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)$ with $\varphi_{0}\in C^{\infty}(\overline{S\times\Sigma})$ and $\varphi_{1}\in C^{\infty}(\overline{S\times\Sigma},C_{\#}^{\infty}(\overline{Z}))$, such that $\varphi_{0}(t,\tilde{x})=0$ and $\varphi_{1}(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}})=0$ if $\tilde{x}\in\Sigma^{\text{in}}$.
-In the case of the convection term, we can integrate by parts to first obtain
-$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}(u_{\varepsilon}\nabla\theta_{\varepsilon},\varphi_{\varepsilon})_{S\times\Omega^{f}_{\varepsilon}}$ $=\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\Bigl{[}-(u_{\varepsilon}\theta_{\varepsilon},\nabla_{\tilde{x}}\varphi_{0})_{S\times\Omega^{f}_{\varepsilon}}-\varepsilon(u_{\varepsilon}\theta_{\varepsilon},\nabla_{\tilde{x}}\varphi_{1})_{S\times\Omega^{f}_{\varepsilon}}-(u_{\varepsilon}\theta_{\varepsilon},\nabla_{y}\varphi_{1})_{S\times\Omega^{f}_{\varepsilon}}$
-$\hskip 42.67912pt+(u_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi_{0})_{S\times\partial\Omega^{f}_{\varepsilon}}+\varepsilon(u_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi_{1})_{S\times\partial\Omega^{f}_{\varepsilon}}\Bigr{]}$
-Utilizing the strong two-scale convergence and (4.3) we can pass to the limit on the right side, producing
-$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}(u_{\varepsilon}\nabla\theta_{\varepsilon},\varphi_{\varepsilon})_{S\times\Omega^{f}_{\varepsilon}}$ $=-(u\theta^{f},\nabla_{\tilde{x}}\varphi_{0})_{S\times\Sigma\times Z}-(u\theta^{f},\nabla_{y}\varphi_{1})_{S\times\Sigma\times Z}+(u_{BC}(\tilde{x},y_{d})\cdot n_{\Sigma}\theta^{f},\varphi_{0})_{S\times\partial\Sigma\times Y^{d}}.$
-Integration by parts in the second term on the right-hand side, that $\operatorname{div}_{y}(u\theta^{f})=0$, the periodicity and boundary values of $u$ yield $(u\theta^{f},\nabla_{y}\varphi_{1})_{S\times\Sigma\times Z}=0$.
-Define $\bar{u}(t,\tilde{x})=\int_{Z}u(t,\tilde{x},y)\,\mathrm{d}y$.
-Then integrating the first term by parts yields
-$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}(u_{\varepsilon}\nabla\theta_{\varepsilon},\varphi_{\varepsilon})_{S\times\Omega^{f}_{\varepsilon}}$ $=(\bar{u}\nabla_{\tilde{x}}\theta^{f},\varphi_{0})_{S\times\Sigma}+\left(\bar{u}\cdot n_{\Sigma}-\int_{Y^{d}}u_{BC}(\tilde{x},y_{d})\cdot n_{\Sigma}\,\mathrm{d}y,\theta^{f}\varphi_{0}\right)_{S\times\partial\Sigma}.$
-To proceed we now investigate the outer boundary conditions of $\bar{u}$.
-By the two-scale convergence of the Dirichlet condition $u_{\varepsilon,BC}$, see Eq.
-4.3, testing with a $\varphi\in C^{\infty}(\overline{\Sigma})$, independent of $y$ and $t$ (which is reasonable since $u_{\varepsilon}$ two-scale converges a.e.
-in $S$), yields
-00 0 $=\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{\Omega^{f}_{\varepsilon}}\operatorname{div}{(u_{\varepsilon}(x))}\varphi(\tilde{x})\,\mathrm{d}x=-\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{\Omega^{f}_{\varepsilon}}u_{\varepsilon}(x)\nabla\varphi(\tilde{x})\,\mathrm{d}x+\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{\partial\Omega^{f}_{\varepsilon}}u_{\varepsilon}\cdot n_{\varepsilon}\varphi(\tilde{x})\,\mathrm{d}\sigma_{x}$
+Via an integration by parts, we have
+$(u_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi_{\varepsilon})_{S\times\Omega^{f}_{\varepsilon}}$ $=-(u_{\varepsilon}\theta_{\varepsilon},\nabla_{\tilde{x}}\varphi_{0})_{S\times\Omega^{f}_{\varepsilon}}-\varepsilon(u_{\varepsilon}\theta_{\varepsilon},\nabla_{\tilde{x}}\varphi_{1})_{S\times\Omega^{f}_{\varepsilon}}-(u_{\varepsilon}\theta_{\varepsilon},\nabla_{y}\varphi_{1})_{S\times\Omega^{f}_{\varepsilon}}$
+$\hskip 42.67912pt+(u_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi_{0})_{S\times\partial\Omega^{f}_{\varepsilon}}+\varepsilon(u_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi_{1})_{S\times\partial\Omega^{f}_{\varepsilon}}$
+At the boundary $\partial\Omega^{f}_{\varepsilon}$, we have $u_{\varepsilon}\cdot n=0$ outside of $\Sigma_{\varepsilon}^{in}\cup\Sigma_{\varepsilon}^{out}$ and $u_{\varepsilon}=u_{\varepsilon,BC}$.
+At $\Sigma_{\varepsilon}^{in}$, we have $\theta_{\varepsilon}^{f}=0$, therefore
+$(u_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi_{0})_{S\times\partial\Omega^{f}_{\varepsilon}}+\varepsilon(u_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi_{1})_{S\times\partial\Omega^{f}_{\varepsilon}}=(u_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi_{0})_{S\times\Sigma_{\varepsilon}^{out}}+\varepsilon(u_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi_{1})_{S\times\Sigma_{\varepsilon}^{out}}.$
+Moreover, as $\theta_{\varepsilon}\in L^{\infty}(S\times\Omega^{f}_{\varepsilon})\cap L^{2}(S;H^{1}(\Omega^{f}_{\varepsilon}))$ uniformly in $\varepsilon$, its traces are also uniformly bounded in $L^{\infty}(S\times\Sigma_{\varepsilon}^{out})$ which is also the case for $u_{\varepsilon,BC}$.
+As a consequence, with $|\Sigma_{\varepsilon}^{out}|\to 0$ in mind:
+$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\varepsilon(u_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi_{1})_{S\times\Sigma_{\varepsilon}^{out}}=\lim_{\varepsilon\to 0}(u_{\varepsilon,BC}\cdot n\theta_{\varepsilon},\varphi_{1})_{S\times\Sigma_{\varepsilon}^{out}}\to 0.$
+Looking at the remaining boundary term, we have
+$(u_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi_{0})_{S\times\Sigma_{\varepsilon}^{out}}=(\chi_{\Sigma_{\varepsilon}^{out}}u_{\varepsilon,BC}\cdot n\theta_{\varepsilon},\varphi_{0})_{S\times\partial\Sigma\times(0,\varepsilon)}.$
+Now, transforming to an $\varepsilon$-independent domain:
+$(u_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi_{0})_{S\times\Sigma_{\varepsilon}^{out}}=\varepsilon\left(\chi_{\Sigma^{out}}u_{BC}\cdot n\theta_{\varepsilon}(\tilde{x},\varepsilon x_{d}),\varphi_{0}\right)_{S\times\partial\Sigma\times(0,1)}$
+Now, with the uniform bounds on $\theta_{\varepsilon}^{f}$, $\frac{1}{\varepsilon}(u_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi_{0})_{S\times\Sigma_{\varepsilon}^{out}}$ must converge and since $\theta_{\varepsilon}^{f}$ is the only $\varepsilon$-dependent function on the right-hand side, it holds
+$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}(u_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi_{0})_{S\times\Sigma_{\varepsilon}^{out}}=(u_{BC}\cdot n\theta^{f},\varphi_{0})_{S\times\Sigma^{out}\times(0,1)}$
+Utilizing $\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightarrow}}\chi_{Z}\theta^{f}$, $u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}u$, and the above computation, we pass to the limit:
+$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}(u_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi_{\varepsilon})_{S\times\Omega^{f}_{\varepsilon}}=$ $-(u\theta^{f},\nabla_{\tilde{x}}\varphi_{0})_{S\times\Sigma\times Z}-(u\theta^{f},\nabla_{y}\varphi_{1})_{S\times\Sigma\times Z}$
+$+(u_{BC}(\tilde{x},y_{d})\cdot n_{\Sigma}\theta^{f},\varphi_{0})_{S\times\partial\Sigma\times(0,1)}.$
+Since $\theta^{f}$ is a scalar function independent of $y\in Z$, $\operatorname{div}_{y}(u\theta^{f})=0$ follows from $\operatorname{div}_{y}u=0$.
+Also, with $u=0$ almost everywhere on $S\times\Sigma\times\Gamma$ and the periodicity of $u$ with respect to $y\in Z$, it holds
+$(u\theta^{f},\nabla_{y}\varphi_{1})_{S\times\Sigma\times Z}=(\operatorname{div}_{y}(u\theta^{f}),\varphi_{1})_{S\times\Sigma\times Z}=0.$
+We introduce $\bar{u}(t,\tilde{x})=\int_{Z}u(t,\tilde{x},y)\,\mathrm{d}y$ and note that $\operatorname{div}_{\tilde{x}}\bar{u}=0$ (Proposition 3.(iii)(iii)). This yields
+$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}(u_{\varepsilon}\nabla\theta_{\varepsilon},\varphi_{\varepsilon})_{S\times\Omega^{f}_{\varepsilon}}$ $=(\bar{u}\nabla_{\tilde{x}}\theta^{f},\varphi_{0})_{S\times\Sigma}+\left(\bar{u}\cdot n_{\Sigma}-\int_{0}^{1}u_{BC}(\tilde{x},s)\cdot n_{\Sigma}\,\mathrm{d}s,\theta^{f}\varphi_{0}\right)_{S\times\partial\Sigma}.$
+Consequently, the limit for the heat flow inside $\Omega^{f}_{\varepsilon}$ takes the form
+$|Z|\langle\partial_{t}\theta^{f},\varphi_{0}\rangle_{L^{2}(S;H^{1}_{\Sigma^{\text{in}}}(\Sigma)^{\prime})}+(\tilde{\kappa}\nabla_{\tilde{x}}\theta^{f},\nabla_{\tilde{x}}\varphi_{0})_{S\times\Sigma}+(\bar{u}\cdot\nabla_{\tilde{x}}\theta^{f},\varphi_{0})_{S\times\Sigma}\\ -\alpha|\Gamma|(\theta^{s}-\theta^{f},\varphi_{0})_{S\times\Sigma}+\left(\theta^{f}\left[\bar{u}\cdot n_{\Sigma}-\int_{0}^{1}u_{BC}(\tilde{x},s)\cdot n_{\Sigma}\,\mathrm{d}s\right],\varphi_{0}\right)_{S\times\partial\Sigma}=(\bar{f}^{f},\varphi_{0})_{S\times\Sigma}$
+which, by density, holds for all $\varphi_{0}\in L^{2}(S;H^{1}_{\Sigma_{in}}(\Sigma))$.
+In this formulation, however, the additional term over $\partial\Sigma$ is a bit puzzling.
+As a next step, we show that it is vanishing.
+To that end, we investigate the outer boundary behavior of $\bar{u}$.
+With the two-scale convergence of $u_{\varepsilon}$ and the choice of the Dirichlet condition $u_{\varepsilon,BC}$, testing with any $\varphi\in C^{\infty}(\overline{\Sigma})$ yields
+$0$ $=\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{\Omega^{f}_{\varepsilon}}\operatorname{div}{u_{\varepsilon}(x)}\varphi(\tilde{x})\,\mathrm{d}x=\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\left[-\int_{\Omega^{f}_{\varepsilon}}u_{\varepsilon}(x)\cdot\begin{pmatrix}\nabla_{\tilde{x}}\varphi(\tilde{x})\\ 0\end{pmatrix}\,\mathrm{d}x+\int_{\partial\Omega^{f}_{\varepsilon}}u_{\varepsilon}\cdot n\varphi(\tilde{x})\,\mathrm{d}\sigma_{x}\right]$
 $=-\int_{\Sigma}\int_{Z}u(\tilde{x},y)\nabla_{\tilde{x}}\varphi(\tilde{x})\,\mathrm{d}y\,\mathrm{d}\tilde{x}+\int_{\partial\Sigma}\left(\int_{0}^{1}u_{BC}(\tilde{x},s)\,\mathrm{d}s\right)\cdot n_{\Sigma}\varphi(\tilde{x})\,\mathrm{d}\sigma_{\tilde{x}}$.
 Integration by parts in the last line, lets us conclude that
-$\bar{u}(t,\tilde{x})\cdot n_{\Sigma}=\int_{0}^{1}u_{BC}(\tilde{x},s)\,\mathrm{d}s\cdot n_{\Sigma}$ (4.5)
-on $\partial\Sigma$ and for almost all $t\in S$.
-We can conclude that the convection term converges to a convection of the effective quantities.
-After calculating all limits $\varepsilon\to 0$ in Eq.
-3.1, one arrives at the weak form
-$\begin{split}\langle\partial_{t}\theta^{s},\varphi\rangle_{L^{2}(S;H^{1}(\Omega)^{\prime})}+|Z|\langle\partial_{t}\theta^{f},\psi\rangle_{L^{2}(S;H^{1}_{\Sigma^{\text{in}}}(\Sigma)^{\prime})}+\kappa^{s}(\nabla\theta^{s},\nabla\varphi)_{S\times\Omega}&+(\tilde{\kappa}\nabla_{\tilde{x}}\theta^{f},\nabla\psi)_{S\times\Sigma}+(\bar{u}\cdot\nabla_{\tilde{x}}\theta^{f},\psi)_{S\times\Sigma}\\ &+\alpha|\Gamma|(\theta^{s}-\theta^{f},\phi-\psi)_{S\times\Sigma}=(\bar{f},\theta^{s})_{S\times\Sigma}\end{split}$ (4.6)
-which, by a density arguments, holds for all $(\varphi,\psi)\in L^{2}(S;H^{1}(\Omega)\times H^{1}_{\Sigma^{\text{in}}}(\Sigma))$. Here, the entries of $\tilde{\kappa}$ and $\bar{f}$ are given by
-$\tilde{\kappa}_{ij}=\int_{Z}\kappa^{f}(\delta_{ij}+\nabla_{y}\omega_{i}\cdot e_{j})\,\mathrm{d}y,\qquad\bar{f}(t,\tilde{x})=\int_{\Gamma}f(t,\tilde{x},y)\,\mathrm{d}\sigma_{y}.$
-Limit procedure for the Stokes system.
+$\bar{u}(t,\tilde{x})\cdot n_{\Sigma}=\int_{0}^{1}u_{BC}(\tilde{x},s)\,\mathrm{d}s\cdot n_{\Sigma}$ (29)
+on $\partial\Sigma$ and for almost all $t\in S$. As a consequence,
+$|Z|\langle\partial_{t}\theta^{f},\varphi_{0}\rangle_{L^{2}(S;H^{1}_{\Sigma^{\text{in}}}(\Sigma)^{\prime})}+(\tilde{\kappa}\nabla_{\tilde{x}}\theta^{f},\nabla_{\tilde{x}}\varphi_{0})_{S\times\Sigma}+(\bar{u}\cdot\nabla_{\tilde{x}}\theta^{f},\varphi_{0})_{S\times\Sigma}\\ -\alpha|\Gamma|(\theta^{s}-\theta^{f},\varphi_{0})_{S\times\Sigma}=(\bar{f}^{f},\varphi_{0})_{S\times\Sigma}$ (30)
+for all $\varphi_{0}\in L^{2}(S;H^{1}_{\Sigma_{in}}(\Sigma))$.
+Finally, we show that $\theta^{f}$ satisfies a homogeneous Dirichlet condition on $\Sigma_{in}$.
+To that end, take $\varphi\in C^{\infty}(\overline{S\times\Sigma})$ with $\varphi=0$ on $\Sigma_{out}$ and $\psi\in C_{\#}^{\infty}(Z)^{d}$ with $\operatorname{div}_{y}{\psi}=0$ and $\psi=0$ on $\Gamma_{d}$.
+With $\phi_{\varepsilon}\in C^{\infty}(\overline{S\times\Omega})$ given by $\phi_{\varepsilon}(t,x)=\varphi(t,\tilde{x})\psi(\nicefrac{{x}}{{\varepsilon}})$, it holds
+$\frac{1}{\varepsilon}\int_{S}\int_{\Omega^{f}_{\varepsilon}}\theta^{f}_{\varepsilon}(t,x)\operatorname{div}_{x}\phi_{\varepsilon}(t,x)\,\mathrm{d}x\,\mathrm{d}t=-\frac{1}{\varepsilon}\int_{S}\int_{\Omega^{f}_{\varepsilon}}\nabla_{x}\theta^{f}_{\varepsilon}(t,x)\cdot\varphi(t,\tilde{x})\psi(\nicefrac{{x}}{{\varepsilon}})\,\mathrm{d}x\,\mathrm{d}t.$
+On the left-hand side, we simplify via $\operatorname{div}_{x}\phi_{\varepsilon}(t,x)=\nabla_{\tilde{x}}\varphi(t,\tilde{x})\cdot\psi(\nicefrac{{x}}{{\varepsilon}})$ (due to $\operatorname{div}_{y}\psi=0$).
+Passing to the limit $\varepsilon\to 0$ on both sides, we then obtain
+$\int_{S}\int_{\Sigma}\int_{Z}\theta^{f}(t,\tilde{x})\nabla_{\tilde{x}}\varphi(t,\tilde{x})\cdot\psi(y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t\\ =-\int_{S}\int_{\Sigma}\int_{Z}\left(\nabla_{\tilde{x}}\theta^{f}(t,\tilde{x})+\nabla_{y}\theta^{f}_{1}(t,\tilde{x},y)\right)\cdot\varphi(t,\tilde{x})\psi(y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t$ (31)
+Via an integration by parts, we can write the right-hand side as
+$\int_{S}\int_{\Sigma}\int_{Z}\left(\nabla_{\tilde{x}}\theta^{f}(t,\tilde{x})+\nabla_{y}\theta^{f}_{1}(t,\tilde{x},y)\right)\cdot\varphi(t,\tilde{x})\psi(y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t\\ =\int_{S}\int_{\Sigma}\int_{Z}\theta^{f}(t,\tilde{x})\nabla_{\tilde{x}}\varphi(t,\tilde{x})\cdot\psi(y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t-\int_{S}\int_{\Sigma^{\text{in}}}\theta^{f}(t,\tilde{x})\varphi(t,\tilde{x})\int_{Z}\psi(y)\,\mathrm{d}y\cdot n_{\Sigma}\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t\\ +\int_{S}\int_{\Sigma}\int_{Z}\theta^{f}_{1}(t,\tilde{x},y)\varphi(t,\tilde{x})\operatorname{div}_{y}{\psi(y)}\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t-\int_{S}\int_{\Sigma}\int_{\partial Z}\theta^{f}_{1}(t,\tilde{x},y)\varphi(t,\tilde{x})\psi(y)\cdot n_{Z}\,\mathrm{d}y\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t.$ (32)
+Given the properties of $\psi$ and the periodicity of $\theta^{f}_{1}$, the last two integrals vanish.
+In addition, seeing that the first term of the right-hand side of Eq.
+32 cancels with the left-hand side of Eq.
+31, we obtain
+$\int_{S}\int_{\Sigma^{\text{in}}}\theta^{f}(t,\tilde{x})\varphi(t,\tilde{x})\int_{Z}\psi(y)\,\mathrm{d}y\cdot n_{\Sigma}\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t=0$
+for all $\psi\in C_{\#}^{\infty}(Z)^{d}$ with $\operatorname{div}_{y}{\psi}=0$ in $Z$ and $\psi=0$ on $\Gamma_{d}$.
+Choosing any such $\psi$ with $\int_{Z}\psi(y)\,\mathrm{d}y\cdot n_{\Sigma}\neq 0$, we conclude
+$\int_{S}\int_{\Sigma^{\text{in}}}\theta^{f}(t,\tilde{x})\varphi(t,\tilde{x})\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t=0.$
+By density, this holds for all $\varphi\in L^{2}(S;H^{1}_{\Sigma^{\text{out}}}(\Sigma))$, which implies $\theta^{f}=0$ almost everywhere on $S\times\Sigma^{\text{in}}$ (note that $\varphi$ is not restricted on $\Sigma^{\text{in}}$).
+Combining Eqs. 31 and 32, we arrive at the following weak form as the limit of Eq. 25:
+$\begin{split}&\langle\partial_{t}\theta^{s},\varphi\rangle_{L^{2}(S;H^{1}(\Omega)^{\prime})}+|Z|\langle\partial_{t}\theta^{f},\psi\rangle_{L^{2}(S;H^{1}_{\Sigma^{\text{in}}}(\Sigma)^{\prime})}+\kappa^{s}(\nabla\theta^{s},\nabla\varphi)_{S\times\Omega}+(\tilde{\kappa}\nabla_{\tilde{x}}\theta^{f},\nabla_{\tilde{x}}\psi)_{S\times\Sigma}\\ &+(\bar{u}\cdot\nabla_{\tilde{x}}\theta^{f},\psi)_{S\times\Sigma}+\alpha|\Gamma|(\theta^{s}-\theta^{f},\varphi-\psi)_{S\times\Sigma}=(\bar{f}^{s},\varphi)_{S\times\Sigma}+(\bar{f}^{f},\psi)_{S\times\Sigma}\end{split}$ (33)
+which holds for all $(\varphi,\psi)\in L^{2}(S;H^{1}(\Omega)\times H^{1}_{\Sigma^{\text{in}}}(\Sigma))$. Here, the entries of $\tilde{\kappa}$ and $\bar{f}^{r}$ ($r=f,s$) are given by
+$\tilde{\kappa}_{ij}=\int_{Z}\kappa^{f}(\delta_{ij}+\nabla_{y}\omega_{i}\cdot e_{j})\,\mathrm{d}y,\qquad\bar{f}^{r}(t,\tilde{x})=\int_{\Gamma}f^{r}(t,\tilde{x},y)\,\mathrm{d}\sigma_{y}.$
+4.2.2 Limit procedure for the Stokes system.
 Shifting to the Stokes system, we observe that our setup is somewhat close to fluid flows in porous media.
 The following steps are therefore similar to the classical derivation of the Darcy equation, see [40, 51].
-There is also quite a lot of literature considering thin porous layers, which is even closer related to the present setup, for example, [22, 37, 26].
-Nevertheless, to better understand the final homogenized model we briefly present the procedure and point out any differences compared to deriving a Darcy equation.
-Let $\varphi\in C_{c}^{\infty}(\Sigma;C_{\#}^{\infty}(\overline{Z}))^{d}$ be a smooth function with compact support in $\Sigma$ and $\varphi(\tilde{x},y)=0$ if $y\in\Gamma\cup\{y_{d}=0\}$.
+There is also extensive literature considering thin porous layers, which is even closer to the present setup, for example, [35, 52, 49].
+Nevertheless, to better understand the final homogenized model, we briefly present the procedure and point out any differences compared to deriving a Darcy equation.
+We start by recalling that the $\varepsilon$-solutions $(u_{\varepsilon},p_{\varepsilon})\in V_{\varepsilon}\times Q_{\varepsilon}$ satisfy
+$(\mu(\theta_{\varepsilon}^{f})\nabla u_{\varepsilon},\nabla\varphi)_{\Omega^{f}_{\varepsilon}}-(p_{\varepsilon},\operatorname{div}{\varphi})_{\Omega^{f}_{\varepsilon}}=0$ (34)
+for all $\varphi\in H_{0}^{1}(\Omega^{f}_{\varepsilon})$ and almost all $t\in S$.
+Let $\varphi\in C_{c}^{\infty}(\Sigma;C_{\#}^{\infty}(Z))^{d}$ with $\varphi=0$ on $\Sigma\times\Gamma_{d}$.
 We test the weak form Eq.
-3.2 with $\varphi_{\varepsilon}\in H_{0}^{1}(\Omega^{f}_{\varepsilon})^{d}$ given by $\varphi_{\varepsilon}(x)=\varepsilon^{k}\varphi(\tilde{x},\nicefrac{{x}}{{\varepsilon}})$ ($k=1,2$) to get (for almost all $(t\in S)$)
+34 with $\varphi_{\varepsilon}\in H_{0}^{1}(\Omega^{f}_{\varepsilon})^{d}$ given by $\varphi_{\varepsilon}(x)=\varepsilon^{k}\varphi(\tilde{x},\nicefrac{{x}}{{\varepsilon}})$ and get (for almost all $(t\in S)$)
 $\left(\mu(\theta_{\varepsilon}^{f})\nabla u_{\varepsilon},\left[\varepsilon^{k}\nabla_{\tilde{x}}\varphi+\varepsilon^{k-1}\nabla_{y}\varphi\right](\tilde{x},\nicefrac{{x}}{{\varepsilon}})\right)_{\Omega^{f}_{\varepsilon}}-\left(p_{\varepsilon},\left[\varepsilon^{k}\operatorname{div}_{\tilde{x}}{\varphi}+\varepsilon^{k-1}\operatorname{div}_{y}{\varphi}\right](\tilde{x},\nicefrac{{x}}{{\varepsilon}})\right)_{\Omega^{f}_{\varepsilon}}=0.$
-If we set $k=2$ and pass to the limit $\varepsilon\to 0$, we see that
-$(p(t,\tilde{x},y),\operatorname{div}_{y}{\varphi}(\tilde{x},y))_{\Sigma\times Z}=0.$
-It follows that $p$ is constant in $y$.
-On the other hand, if we take only functions with $\operatorname{div}_{y}{\varphi}=0$ and $\operatorname{div}_{\tilde{x}}{\left(\int_{Z}\varphi\,\mathrm{d}y\right)}=0$ and set $k=1$, the limit yields
-$(\mu(\theta^{f}(t,\tilde{x}))\nabla_{y}u(t,\tilde{x},y),\nabla_{y}\varphi(\tilde{x},y))_{\Sigma\times Z}=0.$ (4.7)
-Here, we used Corollary 2. Next, let us introduce
-$V=\left\{\varphi\in L^{2}(\Sigma;H^{1}_{\#}(Z))^{d}:\,\begin{array}[]{c}\operatorname{div}_{y}{\varphi}=0,\,\operatorname{div}_{\tilde{x}}{\left(\int_{Z}\varphi\,\mathrm{d}y\right)}=0,\\ \varphi=0\text{ on }\Gamma\cup\{y_{d}=0\}\text{ and }\int_{Z}\varphi\,\mathrm{d}y\cdot n_{\Sigma}=0\text{ on }\partial\Sigma\end{array}\right\}.$
-By a density argument Eq.
-4.7 also holds for all $\varphi\in V$.
-By the choice of test functions the pressure $p$ does not appear in (4.7).
-To recover the pressure, we use the properties of the orthogonal of $V$.
-With the results from [22, Lemma 5.3 and Remark 5.5.$(i)$], one obtains that
-$V^{\perp}=\left\{q=\nabla_{\tilde{x}}q_{0}+\nabla_{y}q_{1}:\,q_{0}\in H^{1}(\Sigma)/\mathbb{R},\,q_{1}\in L^{2}_{0}(\Sigma\times Z)\right\}\subset L^{2}(\Sigma;\mathcal{H}(Z)^{\prime}),$
-where $\mathcal{H}(Z)\coloneqq\left\{\varphi\in H^{1}_{\#}(Z)^{d}:\varphi(y)=0\text{ if }y\in\Gamma\cup\{y_{d}=0\}\right\}$.
-The decomposition into $(q_{0},q_{1})$ is also unique for every $q\in V^{\perp}$.
-Therefore, there exists a $q=(q_{0},q_{1})\in L^{2}(S;V^{\perp})$ such that $\mu(\theta^{f})\Delta_{y}u=\nabla_{\tilde{x}}q_{0}+\nabla_{y}q_{1}$ (for almost all $t\in S$) understood as an element in $L^{2}(\Sigma;\mathcal{H}(Z)^{\prime})$.
-Collecting the boundary conditions of $u$ derived in Eq.
-4.5 and Proposition 1 as well as the above characterization, we see that
-$(\bar{u},u,q_{0},q_{1})\in L^{2}(S;H(\operatorname{div},\Sigma)\times L^{2}(\Sigma;H^{1}(Z))\times H^{1}(\Sigma)/\mathbb{R}\times L^{2}_{0}(\Sigma\times Z))$
+We set $k=2$ and pass to the limit $\varepsilon\to 0$ with $\varepsilon^{2}p_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}p$ (see Proposition 3.(iv)(iv) and note that an additional $\varepsilon^{-1}$ is consumed via the two-scale limit, see Definition 1) to get
+$(p(t,\tilde{x},y),\operatorname{div}_{y}{\varphi}(\tilde{x},y))_{\Sigma\times Z}=0$
+which, as this holds for all $\varphi\in C_{c}^{\infty}(\Sigma;C_{\#}^{\infty}(Z))^{d}$ with $\varphi=0$ on $\Sigma\times\Gamma_{d}$, implies that $p$ is constant in $y$.
+Next, we set $k=1$ and restrict testing to functions for which, additionally, $\operatorname{div}_{y}{\varphi}=0$ holds. The corresponding limit yields
+$(\mu(\theta^{f})\nabla_{y}u,\nabla_{y}\varphi)_{\Sigma\times Z}-(p,\operatorname{div}_{\tilde{x}}\varphi)_{\Sigma\times Z}=0$ (35)
+where we used $\mu(\theta_{\varepsilon}^{f})\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\mu(\theta_{\varepsilon}^{f})$ (Corollary 2), $\varepsilon\nabla u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\nabla_{y}u$ (Proposition 3.(iii)(iii)), and $\varepsilon^{2}p_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}p$ (Proposition 3.(iv)(iv)).
+Equation 35 must, by density, also hold for all $\varphi\in L^{2}(\Sigma;H^{1}_{\#,\Gamma_{d}}(Z))$ with $\operatorname{div}_{\tilde{x}}{\left(\int_{Z}\varphi\,\mathrm{d}y\right)}\in L^{2}(\Sigma)$.
+Next, let us introduce
+$V=\left\{\varphi\in L^{2}(\Sigma;H^{1}_{\#}(Z))^{d}:\,\begin{array}[]{c}\operatorname{div}_{y}{\varphi}=0,\,\operatorname{div}_{\tilde{x}}{\left(\int_{Z}\varphi\,\mathrm{d}y\right)}=0,\\ \varphi=0\text{ on }\Gamma_{d}\text{ and }\int_{Z}\varphi\,\mathrm{d}y\cdot n_{\Sigma}=0\text{ on }\partial\Sigma\end{array}\right\}.$
+Naturally, Eq. 35 also holds for every $\varphi\in V$. With the results from [35, Lemma 5.3 and Remark 5.5.(i)(i)], the orthogonal of $V$ is characterized by
+$V^{\perp}=\left\{q=\nabla_{\tilde{x}}q_{0}+\nabla_{y}q_{1}:\,q_{0}\in H^{1}(\Sigma)/\mathbb{R},\,q_{1}\in L^{2}_{0}(\Sigma\times Z)\right\}\subset L^{2}(\Sigma;H^{1}_{\#,\Gamma_{d}}(Z)^{\prime})^{d}.$
+This decomposition into $(q_{0},q_{1})$ is also unique for every $q\in V^{\perp}$.
+Therefore, there exists a $q=(q_{0},q_{1})\in L^{2}(S;V^{\perp})$ such that $\mu(\theta^{f})\Delta_{y}u=\nabla_{\tilde{x}}q_{0}+\nabla_{y}q_{1}$ (for almost all $t\in S$) understood as an element in $L^{2}(\Sigma;H^{1}_{\#,\Gamma_{d}}(Z)^{\prime})^{d}$.
+With Eq.
+35 it follows that $q_{0}=p$ almost everywhere up to a constant.
+Collecting the boundary conditions of $u$ derived in Eq. 29 and Proposition 3 as well as the above characterization, we see that
+$(\bar{u},u,p,q_{1})\in L^{2}(S;H^{1}(\Sigma)\times L^{2}(\Sigma;H^{1}(Z))\times H^{1}(\Sigma)/\mathbb{R}\times L^{2}_{0}(\Sigma\times Z))$
 are weak solutions to the following two-pressure Stokes model:
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+$-\mu(\theta^{f})\Delta_{y}u+\nabla_{y}q_{1}$ $=-\nabla_{\tilde{x}}p\quad$ $\text{ in }S\times\Sigma\times Z,$ (36a)
+$\operatorname{div}_{y}{u}$ $=0$ $\text{ in }S\times\Sigma\times Z,$ (36b)
+$u$ $=0$ $\text{ on }S\times\Sigma\times\Gamma,$ (36c)
+$u$ $=u_{\text{motion}}$ $\text{ on }S\times\Sigma\times\{y_{d}=0\},$ (36d)
+$\bar{u}$ $=\int_{Z}u\,\mathrm{d}y$ $\text{ on }S\times\Sigma,$ (36e)
+$\operatorname{div}_{\tilde{x}}\bar{u}$ $=0$ $\text{ in }S\times\Sigma,$ (36f)
+$\bar{u}(t,\tilde{x})\cdot n_{\Sigma}$ $=\int_{0}^{1}u_{BC}(\tilde{x},s)\,\mathrm{d}s\cdot n_{\Sigma}\quad$ $\text{ on }S\times\partial\Sigma.$ (36g)
 Note that both $u_{\text{motion}}\cdot e_{d}=0$ and Assumption (A5) (which guarantees $\int_{\partial\Sigma}n_{\Sigma}\cdot\int_{0}^{1}u_{BC}=0$) are essential to ensure that the boundary conditions are compatible with the incompressibility requirements.
-Furthermore, we introduced the new pressure-like functions $q_{0},q_{1}$ that still have no clear connection to the pressure limit $p$ of the microscale model.
-We first show $p=q_{0}$.
-To this end, we test the original momentum equation of the micro problem with $\varepsilon\varphi$ for $\varphi\in C_{c}^{\infty}(\Sigma;C_{\#}^{\infty}(\overline{Z}))^{d}$ with $\varphi=0$ on $\Gamma\cup\{y_{d}=0\}$ and $\operatorname{div}_{y}{\varphi}=0$:
-$(\mu(\theta^{f})\nabla_{y}u,\nabla_{y}\varphi)_{\Sigma\times Z}-(p,\operatorname{div}_{\tilde{x}}\varphi)_{\Sigma\times Z}=0.$ (4.8)
-With a density argument, this also holds for all $\varphi\in L^{2}(\Sigma;\mathcal{H}(Z))$ with $\operatorname{div}_{\tilde{x}}{\left(\int_{Z}\varphi\,\mathrm{d}y\right)}\in L^{2}(\Sigma)$ and $\operatorname{div}_{y}{\varphi}=0$.
-Since the representation $\mu(\theta^{f})\Delta_{y}u=\nabla_{\tilde{x}}q_{0}+\nabla_{y}q_{1}$ is unique, we deduce from Eq.
-4.8 that it needs to hold $p=q_{0}$.
-We recovered the pressure $p$ for the two-pressure model and therefore replace, in the following, $q_{0}$ by $p$.
-By the linearity of the above two-pressure Stokes model, we can further simplify the problem by constructing suitable cell problems.
-The first set of cell problems are quite standard, see [3] where they are used to characterize the permeability in a porous medium: for $i=1,\dots,d-1$ find $(\xi_{i},\eta_{i})\in H^{1}_{\#}(Z)^{d}\times L^{2}_{0,\#}(Z)$ such that
-$-\Delta_{y}\xi_{i}-\nabla_{y}\eta_{i}$ $=e_{i}\quad$ in $\text{in }Z,$
-$\operatorname{div}_{y}{(\xi_{i})}$ $=0$ in $\text{in }Z,$ (4.9)
-$\xi_{i}$ $=0$ on $\text{on }\Gamma\cup\{y_{d}=0\}.$
+We can further simplify the two-pressure Stokes-problem by constructing suitable cell problems.
+The first set of cell problems are quite standard, see [40] where they are used to characterize the permeability in a porous medium: for $i=1,\dots,d-1$, find $(\xi_{i},\eta_{i})\in H^{1}_{\#}(Z)^{d}\times L^{2}_{0,\#}(Z)$ such that
+$-\Delta_{y}\xi_{i}-\nabla_{y}\eta_{i}$ $=e_{i}\quad$ $\text{in }Z,$
+$\operatorname{div}_{y}{(\xi_{i})}$ $=0$ $\text{in }Z,$ (37)
+$\xi_{i}$ $=0$ $\text{on }\Gamma_{d}.$
 Additionally, a cell problem originating from the nonhomogeneous Dirichlet condition at the bottom boundary appears: find $(\xi_{0},\eta_{0})\in H^{1}_{\#}(Z)^{d}\times L^{2}_{0,\#}(Z)$ such that
-$-\Delta_{y}\xi_{0}-\nabla_{y}\eta_{0}$ $=0\quad$ in $\text{in }Z,$
-$\operatorname{div}_{y}{(\xi_{0})}$ $=0$ in $\text{in }Z,$ (4.10)
-$\xi_{0}$ $=0$ on $\text{on }\Gamma,$
-$\xi_{0}$ $=u_{\text{motion}}\quad$ on $\text{on }\{y_{d}=0\}.$
-Note, the cell problem (4.2) is a well-defined equation and admits a unique solution, given that $u_{\text{motion}}\cdot e_{d}=0$. We have the following property of the cell solutions:
+$-\Delta_{y}\xi_{0}-\nabla_{y}\eta_{0}$ $=0\quad$ $\text{in }Z,$
+$\operatorname{div}_{y}{(\xi_{0})}$ $=0$ $\text{in }Z,$ (38)
+$\xi_{0}$ $=0$ $\text{on }\Gamma,$
+$\xi_{0}$ $=u_{\text{motion}}\quad$ $\text{on }\{y_{d}=0\}.$
+Note, the cell problem (4.2.2) is a well-defined equation and admits a unique solution, given that $u_{\text{motion}}\cdot e_{d}=0$. We have the following property of the cell solutions:
 Proposition 6.
-For the cell solutions $\xi_{i}\in H^{1}_{\#}(Z)^{d}$ ($i=0,\dots,d-1$) of Section 4.2 and Section 4.2, respectively, it holds
+For the cell solutions $\xi_{i}\in H^{1}_{\#}(Z)^{d}$ ($i=0,\dots,d-1$) of Section 4.2.2 and Section 4.2.2, respectively, it holds
 $\int_{Z}\xi_{i}\cdot e_{d}\,\mathrm{d}y=0.$
@@
 By the choice of boundary conditions and periodicity, the above integral over $\partial Z$ vanishes and we obtain the desired result. ∎
-Given the solutions of the cell problems (4.2) and (4.2), we obtain for a.a. $(t,\tilde{x},y)\in S\times\Sigma\times Z$ the representations
-$u(t,\tilde{x},y)=\xi_{0}(y)-\sum_{i=1}^{d-1}\frac{\xi_{i}(y)}{\mu(\theta^{f}(t,\tilde{x}))}\partial_{x_{i}}p(t,\tilde{x}),\quad q_{1}(t,\tilde{x},y)=\eta_{0}(y)-\sum_{i=1}^{d-1}\frac{\eta_{i}(y)}{\mu(\theta^{f}(t,\tilde{x}))}\partial_{x_{i}}p(t,\tilde{x}).$ (4.11)
+Given the solutions of the cell problems (4.2.2) and (4.2.2), we obtain for a.a. $(t,\tilde{x},y)\in S\times\Sigma\times Z$ the representations
+[…]
 This can be checked easily by plugging them into the two-pressure Stokes system.
-Looking at $\bar{u}=\int_{Z}u\,\mathrm{d}y$, we can also check that with $\operatorname{div}_{\tilde{x}}\bar{u}=0$ and by testing with $\psi\in H^{1}(\Sigma)$ while using Eq.
-4.11:
-$\begin{split}0&=\left(\operatorname{div}_{\tilde{x}}{\bar{u}},\psi\right)_{\Sigma}=-\left(\bar{u},\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\left(\bar{u}\cdot n_{\Sigma},\psi\right)_{\partial\Sigma}\\ &=-\left(\int_{Z}\xi_{0}\,\mathrm{d}y,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\sum_{i=1}^{d-1}\left(\frac{\partial_{i}p}{\mu(\theta^{f})}\int_{Z}\xi_{i}\,\mathrm{d}y,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\left(\left(\int_{0}^{1}u_{BC}(\cdot,s)\,\mathrm{d}s\right)\cdot n_{\Sigma},\psi\right)_{\partial\Sigma}\\ &=-(\bar{\xi}_{0},\nabla_{\tilde{x}}\psi)_{\Sigma}+\left(\frac{1}{\mu(\theta^{f})}K\nabla_{\tilde{x}}p,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+(\bar{u}_{\text{BC}}\cdot n_{\Sigma},\psi)_{\partial\Sigma}.\end{split}$ (4.12)
+Looking at $\bar{u}=\int_{Z}u\,\mathrm{d}y$, we can also check that with $\operatorname{div}_{\tilde{x}}\bar{u}=0$ and by testing Eq.
+36a with $\psi\in H^{1}(\Sigma)$ while using Eq.
+39:
+$\begin{split}0&=\left(\operatorname{div}_{\tilde{x}}{\bar{u}},\psi\right)_{\Sigma}=-\left(\bar{u},\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\left(\bar{u}\cdot n_{\Sigma},\psi\right)_{\partial\Sigma}\\ &=-\left(\int_{Z}\xi_{0}\,\mathrm{d}y,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\sum_{i=1}^{d-1}\left(\frac{\partial_{i}p}{\mu(\theta^{f})}\int_{Z}\xi_{i}\,\mathrm{d}y,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\left(\left(\int_{0}^{1}u_{BC}(\cdot,s)\,\mathrm{d}s\right)\cdot n_{\Sigma},\psi\right)_{\partial\Sigma}\\ &=-(\bar{\xi}_{0},\nabla_{\tilde{x}}\psi)_{\Sigma}+\left(\frac{1}{\mu(\theta^{f})}K\nabla_{\tilde{x}}p,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+(\bar{u}_{\text{BC}}\cdot n_{\Sigma},\psi)_{\partial\Sigma}.\end{split}$ (40)
 Where for the last step we defined
-$\bar{\xi}_{0}\coloneqq\int_{Z}\xi_{0}(y)\,\mathrm{d}y,\quad K_{ij}\coloneqq\int_{Z}\xi_{j}(y)\cdot e_{i}\,\mathrm{d}y\quad\text{and}\quad\bar{u}_{\text{BC}}\coloneqq\int_{0}^{1}u_{BC}(\cdot,s)\,\mathrm{d}s.$ (4.13)
-Note, Eq. 4.12 is similar to a Darcy equation, but with an additional term induced by the Dirichlet condition at $y_{d}=0$. The following Theorem 3 collects all equations and shows the complete model.
-Theorem 3 (Homogenized model).
-Let the Assumptions (A1)-(A7) be fulfilled and let $(\theta^{s},\theta^{f},u,p)$ be the limit functions defined in Proposition 1.
-Set $\bar{u}(t,\tilde{x})=\int_{Z}u(t,\tilde{x},y)\,\mathrm{d}y$.
-Then $(\theta^{s},\theta^{f},\bar{u},p)$ are weak solutions of the following problem: inside $\Omega$ it holds
-[…]
-[…]
-[…]
-[…]
-where $\bar{f}(t,\tilde{x})=\int_{\Gamma}f(t,\tilde{x},y)\,\mathrm{d}\sigma_{y}$. The fluid temperature $\theta^{f}$ fulfills
-[…]
-[…]
-[…]
-[…]
-with the effective conductivity $\tilde{\kappa}\in\mathbb{R}^{d\times d}$ given by the cell solutions $\{\omega_{i}\}_{i=1}^{d-1}$ of Eq. 4.4, $\omega_{d}=0$ and
-$\tilde{\kappa}_{ij}=\int_{Z}\kappa^{f}(\delta_{ij}+\nabla_{y}\omega_{i}\cdot e_{j})\,\mathrm{d}y.$
-The fluid movement in the layer is governed by
-[…]
-[…]
-[…]
-[…]
-here the permeability $K\in\mathbb{R}^{d\times d}$ is given by the solutions $\{\xi_{i}\}_{i=1}^{d-1}$ of the cell problems Section 4.2 via
-$K_{ij}=\int_{Z}\xi_{j}\cdot e_{i}\,\mathrm{d}y$
-where we set $\xi_{d}=0$ and $\bar{\xi}_{0}=\int_{Z}\xi_{0}(y)\,\mathrm{d}y$, with $\xi_{0}$ given by Section 4.2.
-Proof.
-That $(\theta^{s},\theta^{f},\bar{u},p)$ fulfill Eq. 4.14 can be derived with the concept of two-scale convergence, as explained and partially carried out above. ∎
-Remark 4.
-Note, that by Proposition 2 and the equality Eq.
-4.14i we obtain that $\bar{u}\cdot e_{d}=0$.
-This is reasonable since we remain with an averaged interface velocity in the effective model and a nonzero flow in direction $d$ would imply a flow into (or out of) the solid domain.
-We recovered, ignoring the temperature-dependent viscosity, the same effective equation as in [12], where the needed assumptions for convergence are reduced, thanks to the two-scale concept.
-Additionally, in 2D the effective problem for the fluid simplifies further.
-By Eq.
-4.14j it holds $\operatorname{div}_{\tilde{x}}(\bar{u})=\partial_{x_{1}}\bar{u}_{1}=0$ and the velocity is a constant solely given by the boundary function $u_{BC}$ in Eq.
-4.14l.
-This linearizes the problem since the temperature equation is independent of the pressure $p$.
-The existence of a solution of model (4.14) is ensured by the existence of the original problem and the two-scale limits.
-Similar to the micro model, for a unique solution, we need additional regularity assumptions.
-Theorem 4 (Uniqueness of the homogenized model).
-If there is a solution of the homogenized problem where $\nabla_{\tilde{x}}\theta\in L^{\infty}(S;L^{2+\delta}(\Sigma))$ and $\nabla_{\tilde{x}}p\in L^{\infty}(S;L^{2+\delta}(\Sigma))$ for some $\delta>0$, then this solution is the only.
-Proof.
-Assume that $(\theta^{f}_{i},\theta^{s}_{i},\bar{u}_{i},p_{i})$, $i=1,2$, are two sets of solutions of the homogenization limit given by System (4.14), i.e., they solve the variational problems given via Eqs.
-4.6 and 4.12.
-In the following, we will denote their differences by $(\bar{\theta}^{f},\bar{\theta}^{s},\bar{\bar{u}},\bar{p})$.
-W.l.o.g.
-we assume that $(\theta_{2}^{f},p_{2})$ is the solution enjoying the higher gradient integrability.
-From Eq. 4.11 we obtain for $\bar{\bar{u}}$, that
-$\|\bar{\bar{u}}\|_{L^{2}(S\times\Sigma)}^{2}\leq C\left(\nabla_{\tilde{x}}\bar{p},\bar{\bar{u}}\right)_{S\times\Sigma}+\left(\left[\frac{1}{\mu(\theta^{f}_{1})}-\frac{1}{\mu(\theta^{f}_{2})}\right]\nabla_{\tilde{x}}p_{2},\bar{\bar{u}}\right)_{S\times\Sigma}$
-which, using the Lipschitz continuity of $\mu$ as well as the higher regularity of $p$, yields
-$\|\bar{\bar{u}}\|_{L^{2}(S\times\Sigma)}\leq C\left(\|\nabla_{\tilde{x}}\bar{p}\|_{L^{2}(S\times\Sigma)}+\|\nabla_{\tilde{x}}p_{2}\|_{L^{\infty}(S;L^{2+\delta}(\Omega))}\|\bar{\theta}^{f}\|_{L^{2}(S;L^{2+\frac{2}{\delta}}(\Sigma))}\right).$
-Testing Eq. 4.12 with $\psi=\bar{p}$, we get:
-$c\|\nabla_{\tilde{x}}\bar{p}\|_{L^{2}(S\times\Sigma)}\leq C\|\nabla_{\tilde{x}}p_{2}\|_{L^{\infty}(S;L^{2+\delta}(\Omega))}\|\bar{\theta}^{f}\|_{L^{2}(S;L^{2+\frac{2}{\delta}}(\Sigma))}$
-As a result, the uniqueness of velocity and pressure follows from the uniqueness of the fluid temperatures via
-$\|\bar{\bar{u}}\|_{L^{2}(S\times\Sigma)}+\|\nabla_{\tilde{x}}\bar{p}\|_{L^{2}(S\times\Sigma)}\leq C\|\bar{\theta}^{f}\|_{L^{2}(S;L^{2+\frac{2}{\delta}}(\Sigma))}.$ (4.15)
-We now look at the weak forms for the heat system (Eq. 4.6) and take their difference:
-$(\partial_{t}\bar{\theta}^{s},\varphi)_{S\times\Omega}+|Z|(\partial_{t}\bar{\theta}^{f},\psi)_{S\times\Sigma}+\kappa^{s}(\nabla\bar{\theta}^{s},\nabla\varphi)_{S\times\Omega}+(\tilde{\kappa}\nabla_{\tilde{x}}\bar{\theta}^{f},\nabla\psi)_{S\times\Sigma}\\ +(\bar{u}_{1}\cdot\nabla_{\tilde{x}}\bar{\theta}^{f},\psi)_{S\times\Sigma}+\alpha|\Gamma|(\bar{\theta}^{s}-\bar{\theta}^{f},\phi-\psi)_{S\times\Sigma}=-(\bar{\bar{u}}\cdot\nabla_{\tilde{x}}\theta^{f}_{2},\psi)_{S\times\Sigma}$
-for all $(\varphi,\psi)\in L^{2}(S;H^{1}(\Omega)\times H^{1}(\Sigma))$.
-For the advection term on the left-hand side, we test with $(\bar{\theta}^{s},\bar{\theta}^{f})\in L^{2}(S;H^{1}(\Omega)\times H^{1}(\Sigma))$ and exploit that $\bar{u}_{1}$ is divergence-free (cf.
-Eq.
-3.5):
-$(\bar{u}_{1}\cdot\nabla_{\tilde{x}}\bar{\theta}^{f},\bar{\theta}^{f})_{S\times\Sigma}=\frac{1}{2}(u_{BC}\cdot n,|\bar{\theta}^{f}|^{2})_{S\times\Sigma}\geq 0.$
-The advective term on the right-hand side can be handled via the additional regularity of $\nabla_{\tilde{x}}\theta^{f}_{2}$.
-With the same line of argument as in Theorem 2 and using the Sobolev embedding $H^{1}\subset L^{2+\frac{2}{\delta}}$ for 2D domains, uniqueness now follows.
-∎
+$\bar{\xi}_{0}\coloneqq\int_{Z}\xi_{0}(y)\,\mathrm{d}y,\quad K_{ij}\coloneqq\int_{Z}\xi_{j}(y)\cdot e_{i}\,\mathrm{d}y\quad\text{and}\quad\bar{u}_{\text{BC}}\coloneqq\int_{0}^{1}u_{BC}(\cdot,s)\,\mathrm{d}s.$ (41)
+Note, Eq. 40 is similar to a Darcy equation, but with an additional term induced by the Dirichlet condition at $y_{d}=0$. This is precisely the weak form of the Darcy-limit given in Theorem 4.
 5 Simulations
 In addition to the previous analysis, we investigate the properties of the effective model with the help of numerical simulations and examine what aspects can be captured in the homogenized problem.
-In particular, we are looking at the effects of the roughness patterns, viscosity function and inflow conditions.
-We also study the accuracy of the homogenization limits by direct comparison with the $\varepsilon$-model for small values of $\varepsilon$.
+In particular, we are looking at the effects of the roughness patterns, viscosity function, and inflow conditions.
+We also study the accuracy of the homogenization limit by direct comparison with the $\varepsilon$-model for small values of $\varepsilon$.
 Note that we focus on the two-dimensional case for this comparison as the computational cost for the $\varepsilon$-problem is prohibitively high in the case $d=3$.
@@
 We use continuous, piecewise linear finite elements to represent temperature fields $(\theta^{f},\theta^{s})$ and pressure $p$, while the velocity $u$ is modeled with continuous piecewise quadratic elements.
-The time stepping is performed with a semi-implicit Euler scheme, where we linearized the convection term by using the velocity field from the previous time step.
+The time stepping is performed with a semi-implicit Euler scheme, where we linearize the convection term by using the velocity field from the previous time step.
 Our implementation is freely available online [39].
-We start by collecting all the parameters that are fixed in the following simulations: We consider the time interval $S=(0,5)$ and the macroscopicdomain is the unit square $\Omega=(0,1)^{2}$.
+We start by collecting all the parameters that are fixed in the following simulations: We consider the time interval $S=(0,5)$, and the macroscopic domain is the unit square $\Omega=(0,1)^{2}$.
 The conductivities are set to $\kappa^{s}=0.5$ and $\kappa^{f}=0.1$, and the heat exchange parameter to $\alpha=1.0$.
-For the initial conditions, everything is set to zero.
-At $\Sigma$, the Dirichlet condition for the velocity is $u_{\text{motion}}=(1,0)^{T}$.
-The heat source is given by
-$f_{\varepsilon}(x)=\frac{1-\nicefrac{{x_{d}}}{{\varepsilon}}}{1-\gamma_{0}},$
-so that there is a higher production at the tip of the roughness, which is motivated by the underlying application of grinding.
+The initial conditions are set to zero.
+At $\Sigma$, the Dirichlet condition of the velocity is set to $u_{\text{motion}}=(1,0)$.
+The heat sources are given by $f^{f}_{\varepsilon}=0$ and
+$f_{\varepsilon}^{s}(x)=\frac{1-\nicefrac{{x_{d}}}{{\varepsilon}}}{1-\gamma_{0}},$
+so that there is a higher production at the tip of the roughness, which is motivated by the underlying application of grinding and the fluid only heats up through exchange with the Robin condition.
 The temperature-dependent viscosity is modeled by a Vogel-Fulcher-Tammann equation:
@@
 A convergence study was carried out with respect to $\Delta t$ and $h$ and the expected linear or quadratic convergence was observed.
-To study the influence of different roughness functions, we investigated two different cases:
+To study the influence of different roughness patterns, we model the solid-fluid interface by the graph of a periodic function $\gamma:Y^{d-1}\to[\gamma_{0},1]$, meaning $\Gamma=\{(\tilde{x},\gamma(\tilde{x})):\tilde{x}\in Y^{d-1}\}$.
+We investigate two different setups:
   • A sine function
@@
   • A pattern given by smoothed rectangles
-    $\gamma_{\text{rect}}(x_{1};\gamma_{0})=\begin{cases}1\quad&\text{, if }x<0.1\text{ or }x>0.9,\\ 1-(1-\gamma_{0})\left(-2\left(\frac{x_{1}-0.1}{0.1}\right)^{3}+3\left(\frac{x_{1}-0.1}{0.1}\right)^{2}\right)\quad&\text{, if }x\in[0.1,0.2),\\ \gamma_{0}\quad&\text{, if }x\in[0.2,0.8],\\ \gamma_{0}+(1-\gamma_{0})\left(-2\left(\frac{x_{1}-0.8}{0.1}\right)^{3}+3\left(\frac{x_{1}-0.8}{0.1}\right)^{2}\right)\quad&\text{, if }x\in(0.8,0.9].\end{cases}$
+[…]
 The reference cell $Z$ as well as the effective parameters for both roughness patterns, and three different $\gamma_{0}$, are stated in Table 1.
+Table 1: Roughness patterns used in the numerical simulations, and the corresponding effective parameters. Since the setup is two-dimensional, all effective parameters reduce to scalar values.
 $Z$ $\gamma_{0}$ 0.1 0.5 0.9
@@
 [Uncaptioned image] $\begin{array}[]{c}\tilde{\kappa}=0.120\,\kappa^{f}\\ K=0.00011\\ \bar{\xi}_{0}=0.0526\end{array}$ $\begin{array}[]{c}\tilde{\kappa}=0.523\,\kappa^{f}\\ K=0.01124\\ \bar{\xi}_{0}=0.2556\end{array}$ $\begin{array}[]{c}\tilde{\kappa}=0.917\,\kappa^{f}\\ K=0.06292\\ \bar{\xi}_{0}=0.4552\end{array}$
-Table 1: Roughness patterns used in the numerical simulations, and the corresponding effective parameters. Since the setup is two-dimensional, all parameters are reduced to scalar values.
-Three different inflow patterns are compared, a linear inflow, a quadratic, and a linear inflow that is not zero above $\gamma_{0}$ given by
-$(i)\ \ u_{\text{BC, lin}}=\begin{cases}\left(1-\frac{x_{2}}{\gamma_{0}}\right)u_{\text{motion}}&\text{, if }x_{2}\leq\gamma_{0},\\ 0&\text{, if }x_{2}>\gamma_{0},\end{cases}\quad$ $,\quad(ii)\ \ u_{\text{BC, quad}}=\begin{cases}\left(1-\left(\frac{x_{2}}{\gamma_{0}}\right)^{2}\right)u_{\text{motion}}&\text{, if }x_{2}\leq\gamma_{0},\\ 0&\text{, if }x_{2}>\gamma_{0},\end{cases}$
-and ($\text{and }(iii)\ \ u_{\text{BC, lin, 2}}$ $=(1-x_{2})u_{\text{motion}}.$
+Three different inflow patterns are compared: a linear inflow, a quadratic inflow, and a linear inflow that is not zero above $\gamma_{0}$ given by
+$(i)\ \ u_{\text{BC, lin}}=\begin{cases}\left(1-\frac{x_{2}}{\gamma_{0}}\right)u_{\text{motion}}&\text{, if }x_{2}\leq\gamma_{0},\\ 0&\text{, if }x_{2}>\gamma_{0},\end{cases}$
+$(ii)\ \ u_{\text{BC, quad}}=\begin{cases}\left(1-\left(\frac{x_{2}}{\gamma_{0}}\right)^{2}\right)u_{\text{motion}}&\text{, if }x_{2}\leq\gamma_{0},\\ 0&\text{, if }x_{2}>\gamma_{0},\end{cases}$
+$(iii)\ \ u_{\text{BC, lin, 2}}=(1-x_{2})u_{\text{motion}}.$
 The effective inflow then equals $\bar{u}_{\text{BC, lin}}=\frac{1}{2}\gamma_{0}$, $\bar{u}_{\text{BC, quad}}=\frac{2}{3}\gamma_{0}$ or $\bar{u}_{\text{BC, lin, 2}}=\frac{1}{2}$.
-With $u_{\text{BC, lin, 2}}$ we investigate how the model behaves if Assumption Item (A5) is not fulfilled.
-In this case, we can not build an extension of the boundary condition to the whole domain by a simple scaling argument.
-Instead, the extension has to be carried out in the thin rough layer directly.
+With $u_{\text{BC, lin, 2}}$ we construct an inflow that does not fulfill (A5) (the inflow is nonzero also for $x_{2}>\gamma_{0}$).
+Hence, we cannot use a simple scaling argument to build an extension of the boundary condition to the whole domain.
+However, we observe that the effective model also works well for this larger class of functions.
 If not specified otherwise, we use $u_{\text{BC, lin}}$ in the following simulation studies.
 Remark 3.
-Please note that the overall simulation scenarios are directly motivated by the grinding application in the way we chose the boundary conditions (inflow/outflow), the roughness patterns, and the viscosity functions.
-However, these simulation experiments are only qualitative, as we are not working with realistic and sensible parameter values for the conductivities, the heat source, etc.
-We refer to [57] where quantitative simulations using realistic material parameters were performed for a single grain scenario.
+Please note that the overall simulation scenarios are motivated by the grinding application.
+However, these simulation experiments are only qualitative, as we are not working with realistic and sensible parameter values for the conductivities, the heat source, etc., and instead only highlight the behavior of the effective model.
+We refer to [8] where quantitative simulations using realistic material parameters were performed for a single-grain scenario.
 Comparison between $\varepsilon$-model and effective model.
-We start by comparing the solutions to the effective model with solutions of the resolved $\varepsilon$-model (where we generally compare with values $\varepsilon=0.1,0.05.,0.01$) for the sinusoidal roughness pattern and $\gamma_{0}=0.5$.
-To give a general overview of the solution, we present them in Fig.
+We start by comparing the solutions to the effective model with solutions of the resolved $\varepsilon$-model (where we compare with $\varepsilon=0.2,0.1,0.05,0.01$) for the sinusoidal roughness pattern and $\gamma_{0}=0.5$.
+We present a general overview of the solution in Fig.
 2 and Fig.
-3(d).
-In the following, we focus mainly on comparing the solutions inside the thin layer as this is the critical area where the interesting features (rough boundary and thin domain) are present.
-As the solutions are defined on different domains (the effective solutions are defined on a line and the $\varepsilon$-solutions on a thin domain), we take an average of the microscopic solution with respect to the vertical direction and compare this with the corresponding effective quantities, as done in similar numerical studies [40, Section 5].
-To this end, we define the averaged functions
-$\begin{split}\bar{\theta}^{f}_{\varepsilon}(t,\tilde{x})\coloneqq\frac{1}{\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}&\int_{0}^{\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}\theta^{f}_{\varepsilon}(t,\tilde{x},s)\,\mathrm{d}s,\quad\bar{p}_{\varepsilon}(t,\tilde{x})\coloneqq\frac{\varepsilon}{\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}\int_{0}^{\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}p_{\varepsilon}(t,\tilde{x},s)\,\mathrm{d}s\\ &\bar{u}_{\varepsilon}(t,\tilde{x})\coloneqq\frac{1}{\varepsilon}\int_{0}^{\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}u_{\varepsilon}(t,\tilde{x},s)\,\mathrm{d}s.\end{split}$ (5.1)
-Note that $\bar{p}_{\varepsilon}$ is scaled with $\varepsilon^{2}$ ($\nicefrac{{\varepsilon^{2}}}{{\varepsilon\gamma}}=\nicefrac{{\varepsilon}}{{\gamma}}$) since the effective pressure $p$ is the limit of $\varepsilon^{2}p_{\varepsilon}$.
-We observe that the effective model captures the resolved $\varepsilon$-model fairly well with regard to both temperature and pressure, as observable in Fig.
-2 to 4.
+3.
+Note that we multiply the microscale pressure $p_{\varepsilon}$ with $\varepsilon^{2}$, since convergence only holds for $\varepsilon^{2}p_{\varepsilon}$.
+In the following, we focus mainly on comparing the solutions inside the thin layer, as this is the critical area where the interesting features (rough boundary and thin domain) are present.
+We observe that the effective model captures the temperature and pressure of the resolved $\varepsilon$-model fairly well (Fig.
+2 and 3).
+To better observe the microscopic fluctuations of temperature and pressure as well as the effective profile of the effective solution, the micro solution is visualized along the line $(0,1)\times\{\nicefrac{{\varepsilon}}{{2}}\}$ in Figure 4 and compared with the effective solution.
 We obtain negative pressure at the inflow ($x_{1}=0$) since the amount of fluid pushed into the domain by $u_{BC,lin}$ is less than the fluid movement induced by the bottom boundary.
-This becomes also apparent, by comparing $\bar{u}_{\text{BC, lin}}$ with the flow induced by the cell problem (4.2), which gives the value $\bar{\xi}_{0}=0.29$ (see Table 1).
-Note that for velocity, the average of the microscopic solution is already close to the effective value of $\bar{u}_{\text{BC, lin}}=\frac{1}{2}\gamma_{0}=0.25$ and only oscillates slightly around this value, see Fig.
-4.
-These oscillations originate mainly from numerical inaccuracies, induced both by the FE-scheme and the numeric integration to compute Eq.
-5.1.
+This becomes also apparent, by comparing $\bar{u}_{\text{BC, lin}}$ with the flow induced by the cell problem (4.2.2), which gives the value $\bar{\xi}_{0}=0.29$ (see Table 1).
 (a) $\varepsilon=0.1$
@@
 Figure 2: Comparison of the temperature profiles, at $t=5$, for the micro and effective model, with sinusoidal roughness.
-The effective fluid temperature $\theta^{f}$ is only defined on the line $(0,1)\times\{0\}$ but for visualization purposes the line is being shown with a thickness
+The effective fluid temperature $\theta^{f}$ is only defined on the line $(0,1)\times\{0\}$, but for visualization purposes, the line is being shown with a thickness.
 (a) $\varepsilon=0.1$
@@
 Figure 3: Comparison of the fluid (top) and pressure (bottom) profiles, at $t=5$, for the micro and effective model, with sinusoidal roughness.
-The line in the effective case is again visualized with a thickness
-Figure 4: Comparison inside the thin layer of the effective solutions with the averaged micro solutions defined in Eq.
-5.1.
-All solutions are shown at the final time step $t=5$ and plotted over $\Sigma$.
-The velocity field for $\varepsilon=0.01$ is left out for a clearer visualization.
-We also want to study numerically the convergence behavior for $\varepsilon\to 0$.
-To this end, we compute inside the thin layer the difference of the effective and averaged solution in the $L^{2}(S\times\Sigma)$ norm, and in the solid bulk domain, we compare in the $L^{2}(S\times\Omega^{s}_{\varepsilon})$ norm.
+The line in the effective case is again visualized with a thickness.
+Figure 4: Comparison of the effective solutions with the microscopic temperature and pressure inside the layer.
+The micro solution is plotted along the line $(0,1)\times\{\nicefrac{{\varepsilon}}{{2}}\}$.
+All solutions are shown at the final time step $t=5$.
+We also want to study numerically the convergence behavior for $\varepsilon\to 0$ explicitly and not only compare visually.
+The solutions, however, are defined on different domains (the effective solutions are defined on a line and the $\varepsilon$-solutions on a thin domain), but given the results in Section 4, the functions in the thin layer can be approximated by the effective solutions via
+[…]
+The functions $\theta^{f}_{1}$ and $p_{1}$ are defined in Equation 27 and (39) and can be computed with the cell solutions given by Equation 28, (4.2.2), (4.2.2) and the homogenized solution.
+With the reconstruction (42), we compute the difference of the micro solution with the effective solution in the $L^{2}(S\times\Omega^{f}_{\varepsilon})$ norm.
 The results are shown in Fig.
 5.
-For temperature and pressure, we observe a convergence of order $\mathcal{O}(\varepsilon)$.
-On the other hand, there is no observable convergence for the velocity, as mentioned before, we assume that this is mainly due to numerical errors.
-The difference between effective and averaged velocity is also already small compared to temperature and pressure error.
+For temperature and pressure, we observe a convergence of order $\mathcal{O}(\varepsilon^{2})$.
+On the other hand, the velocity converges only with order $\mathcal{O}(\varepsilon)$, which is to be expected given that the temperature and pressure reconstruction also include higher order terms compared to the velocity.
+Additionally, the temperature in the solid $\theta^{s}_{\varepsilon}$ converges also with order $\mathcal{O}(\varepsilon)$ in $L^{2}(S\times\Omega^{s}_{\varepsilon})$, but is not depicted for clearer visualization.
 Figure 5: Convergence study for the limit $\varepsilon\to 0$.
 Both axes are on a logarithmic scale.
-Inside the thin layer, meaning for $\theta^{f},p,u$, we present the error in the $L^{2}(S\times\Sigma)$ norm.
-The error for $\theta^{s}$ is computed in $L^{2}(S\times\Omega^{s}_{\varepsilon})$.
-To estimate the order of convergence, the black line equals the identity.
+The error between microscopic resolved simulations and the effective quantities reconstructed with Equation 42 in the $L^{2}(S\times\Omega^{f}_{\varepsilon})$ norm.
+To better estimate the order of convergence, the black lines show different powers of $\varepsilon$.
 Varying the inflow function.
@@
 6.
+Here we keep the visualization scheme from Figure 4 for the temperature and pressure to highlight the effective solution profile, and not explicitly compare the error in any norm.
+For the fluid velocity, we compare along the vertical line $\{0.5\}\times(0,\varepsilon)$ and use Equation 42 to compute the effective profile.
 We first observe that, naturally, a higher average fluid flow leads to a cooler system since the heat energy is removed faster from the system.
 Furthermore, inflows $u_{\text{BC, quad}}$ and $u_{\text{BC, lin, 2}}$ result in a positive pressure at the inflow interface, since the amount of fluid pressed into the domain is larger than the amount moved by the bottom boundary condition.
-The homogenized model can capture the influence of the different inflow conditions fairly well with regard to temperature and pressure.
-This holds true even for $u_{\text{BC, lin, 2}}$, which does not satisfy the assumptions made in Section 2.3.
+The homogenized model can capture the influence of the different inflow conditions fairly well with regard to all quantities.
+This holds true even for $u_{\text{BC, lin, 2}}$, which does not directly satisfy the assumptions made in (A5).
 Figure 6: Comparison of the fluid temperature, pressure, and horizontal velocity component for different inflow functions, at $t=5$.
-In black the micro solution, and in orange the effective results.
-The velocity is visualized along $\{0.5\}\times(0,\varepsilon)$.
-For the other functions, we again compare the average (5.1).
-Varying the grain height $\gamma_{0}$.
+In black, the micro solution, and in orange, the effective results.
+The velocity is compared along $\{0.5\}\times(0,\varepsilon)$.
+Varying the grain height $\mathbf{\gamma_{0}}$.
 We investigate the influence of the minimal height of the interface $\gamma_{0}$.
@@
 We use the inflow function $u_{\text{BC, lin, 2}}$ as it is independent of $\gamma_{0}$; this allows us to focus on the influence of the interface height.
-Similarly, we modify the heat source to $f_{\varepsilon}=\tfrac{1}{|\Gamma|}$ ensuring that the total amount of heat produced is one and is independent of $\gamma_{0}$.
+Similarly, we modify the heat source to $f_{\varepsilon}=\tfrac{1}{|\Gamma|}$, ensuring that the total amount of heat produced is one and is independent of $\gamma_{0}$.
 The simulated solutions, again at $t=5$, are presented in Fig.
@@
 As in the results before, the homogenized model shows good agreement with the exact solutions for $\varepsilon=0.05$.
+Given that we plot along the line at $x_{2}=\nicefrac{{\varepsilon}}{{2}}$, there appear jumps for the case $\gamma_{0}=0.1$ since the line is not always inside the domain.
 The grain height has a strong influence on the pressure profile; since we have a fixed amount of inflow for all cases, a smaller gap necessarily leads to higher pressures inside the layer.
@@
 In this simulation scenario for the two-dimensional case, we demonstrate the influence of different roughness patterns.
-We utilize the same setup as in the previous case for varying roughness height but this time we fix $\gamma_{0}=0.5$ and compare the differences between $\gamma_{\text{sine}}$ and $\gamma_{\text{rect}}$.
+We utilize the same setup as in the previous case for varying roughness height, but this time we fix $\gamma_{0}=0.5$ and compare the differences between $\gamma_{\text{sine}}$ and $\gamma_{\text{rect}}$.
 As before, we present the results at the last time step, in Fig.
@@
 The homogenized model accurately reflects the influence of the roughness pattern on all solution fields.
-For 2D we can conclude, that the derived effective model yields a good approximation of the resolved micro model and the temperature-dependent viscosity has a, rather large, impact on the pressure profile.
+For the 2D case, we can conclude that the derived effective model yields a good approximation of the resolved micro model, and the temperature-dependent viscosity only has an impact on the pressure profile.
 Figure 8: Comparison of the fluid temperature, pressure, and horizontal velocity component for different roughness types, at $t=5$. The solution is visualized as in Fig. 6.
@@
 Additionally, we consider a local heat source given by
-$f_{\varepsilon}(x)=\begin{cases}1\quad&\text{, if }x_{2}>0.5\text{ and }|x_{1}-0.5|<0.1,\\ 0&\text{ else}.\end{cases}$
+$f_{\varepsilon}^{s}(x)=\begin{cases}1\quad&\text{, if }x_{2}>0.5\text{ and }|x_{1}-0.5|<0.1,\\ 0&\text{ else}.\end{cases}$
 This right-hand side is chosen in such a way that local variations in the heat production occur, leading to local differences in the viscosity.
-The inflow is defined by $u_{\text{BC, lin}}$ with $\gamma=0.5$ and $u_{\text{motion}}=(1,0,0)^{T}$.
-For the roughness we define
+The inflow is defined by $u_{\text{BC, lin}}$ with $\gamma_{0}=0.5$ and $u_{\text{motion}}=(1,0,0)^{T}$.
+For the roughness, we again model the reference cell as a graph and define
 $\gamma=1-0.9\left(\cos{(2\pi(x_{2}-0.5)^{2})}\cos(2\pi(x_{1}-0.25))\right)^{2},$
 so that the effective quantities behave differently in directions $e_{1}$ and $e_{2}$. The corresponding effective parameters are collected in Table 2.
+Table 2: Effective parameters for the three-dimensional simulation. The effective values corresponding to dimension $d$ are not listed, since they are zero.
 $Z$ $\tilde{\kappa}$ $K$ $\bar{\xi}_{0}$
 [Uncaptioned image] $\left(\begin{array}[]{cc}0.493&0.0\\ 0.0&0.635\end{array}\right)$ $\left(\begin{array}[]{cc}0.004566&0.0\\ 0.0&0.004947\end{array}\right)$ $\left(\begin{array}[]{c}0.134\\ 0.0\end{array}\right)$
-Table 2: Effective parameters for the three-dimensional simulation. The effective values corresponding to dimension $d$ are not listed, since they are zero.
 The nonlinearity is resolved by splitting up the problem and doing the time stepping for temperature and Stokes problem sequentially.
@@
 With the concept of two-scale convergence, an effective model was derived where the effective fluid temperature is given by an interface function.
-We point out that, for the fluid velocity and pressure, we recovered the results of [12] with an additional coupling to the temperature field through viscosity.
 Multiple simulation results validate the derived model and show that temperature-dependent viscosity also plays a non-negligible role in the limit $\varepsilon\to 0$.
+The main contribution of this article is the $L^{\infty}$ estimate for the temperature equation in the thin rough layer and the homogenization of a highly non-linear coupled problem.
 In reality, other parameters (such as conductivity or heat exchange parameter) may also be temperature dependent.
@@
 Another aspect that could be investigated further is the current $\varepsilon$-scaling of $\kappa$ in the thin layer.
-Removing this scaling complicates the analysis and may lead to an effective equation without diffusion on the interface $\Sigma$.
-Regarding the motivating application of a grinding process, we could derive a justification for the models used in the literature; see [33].
-Currently, we are looking at the idealized scenario of periodic roughness.
+Removing this scaling complicates the analysis and may lead to an effective equation without diffusion on the interface $\Sigma$, as in [33, 4].
+Regarding the motivating application of a grinding process, we could derive a justification for the models used in the literature [4], as explained already in the introduction.
+Currently, we are looking at a strongly idealized scenario of periodic roughness and non-touching roughness with the bottom boundary.
 In future research, we plan to extend this model to real grinding wheels and workpiece surfaces.
-There, we also plan to investigate the effect of including microstructures in the whole process.
+Another interesting direction is studying the impact of direct contact between the workpiece and the wheel on the fluid behavior in the contact zone.
+Additionally, a multiscale approach, also including microscopic material removal, is currently being investigated, studying the effect of microstructures on the whole process.
 Acknowledgements
+We thank the anonymous reviewers for their constructive and detailed feedback, which helped to improve the article.
 TF acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project nr.